### v1

# secReferences in the Online Appendix

Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ objectives and how their strategies evolve over time. Using data from 2,500 campaigns, we find that NGOs are eight times more likely to launch campaigns on Annual General Meeting (AGM) dates. These campaigns increase media exposure and stakeholder scrutiny, raising the likelihood of consumer boycotts and related shareholder proposals at the following AGM of the targeted firm but have no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM outcomes, revealing the trade-offs they face to drive corporate change.
1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels. Through tactics of “private politics”—such as boycotts and shaming campaigns—NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017). Many studies examine how campaigns by both large and small NGOs impact and delegitimate targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009). For instance, just weeks before McDonald’s 2006 Annual General Meeting (AGM), Greenpeace launched a campaign branding the company ‘McAmazon‘ for alleged links to Amazon deforestation, prompting McDonald’s to adopt a new sustainable sourcing policy within days (Goldberg and Droste Yagan, 2007). Similar campaigns have targeted companies across sectors, including Nike, Shell, Monsanto, and BNP Paribas. However, we know relatively little about NGOs’ objectives and trade-offs and how they vary over their life cycle, shaping their campaign strategies toward misbehaving firms and, ultimately, their capacity to drive societal change.
To illustrate how objectives vary over time and in the life of an NGO, our analysis spans a wide range of organizations, from smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club.‘ However, we face several challenges, particularly due to the scarcity of data on NGOs. Donation data for non-political nonprofits is difficult to access because of donor privacy concerns, [^1] Data exist for donations to political parties and charities thanks to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023). This paper differs from this literature in that it does not focus on political donations or the resulting campaigns, but rather on the incentives that NGOs face when deciding to campaign against a firm. and NGOs have much more lenient transparency requirements than publicly listed corporations, complicating the tracking of their fund usage.
The absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot be easily reduced to profit maximization (Maxwell, 2012; Fioretti, 2022). The literature emphasizing extreme behaviors of non-profits, with some contribute to the greater good (e.g., Daubanes and Rochet, 2019) whereas others are captured by corporate interests (e.g., Bertrand et al., 2021). Therefore, this paper conducts a forensic investigation that infers NGOs’ objectives from their reactions to external incentives, focusing on the primary—and often only—observable data on NGOs: their campaigns, which must necessarily be public and observable to influence firms and their stakeholders.
The starting point of our analysis is that NGO campaigns are strategically timed. Two key factors influence an NGO’s ability to capture stakeholders’ attention. First, media coverage amplifies NGO campaigns, making shaming and boycotting efforts more effective. Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from workers, consumers, shareholders, suppliers, and the public. We demonstrate that these two factors create a trade-off: NGOs must choose between seeking influence or enhancing their reputation.
We find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM. The AGM attracts the highest media coverage, as media outlets engage with managers and key shareholders to discuss the firm’s current achievements and future outlook (Dyck et al., 2008). This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, which might proxy for the NGO’s reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017).
However, a late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020). Nevertheless, we find that the heightened visibility of campaigns timed with the AGM increases the likelihood of (i) a shareholder proposal on a related topic the following year and (ii) a boycott campaign in the subsequent months, compared to campaigns launched at other times. In contrast, campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
We argue that NGOs navigate this tradeoff differently depending on their accummulated reputation. Employing over ten years of campaign data against the largest US corporations, we first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly. Our analysis rules out learning this aspect of proxy voting as an explanation for our finding, as we observe the same trend for NGOs whose first campaign occurred just after an AGM, meaning they already knew they could not influence the vote by campaigning that late. We also find no evidence that NGO selects different target firms as they gain more campaigning experience.
Second, we demonstrate that the success of a shareholder proposal at a current AGM is linked to how many times the NGO backing the proposal has campaigned on AGM dates in the past. We interpret this as a reduced-form proxy for the reputation the NGO has built over time. However, there are diminishing returns to targeting AGM dates: once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes by campaigning earlier. Conversely, we find that accumulating experience at campaigning outside AGMs does not help NGOs at swaying shareholder proposal votes.
Empirically, our approach leverages the exogenous timing of AGM meetings, which in the U.S. remains relatively stable within firms over time due to Securities and Exchange Commission (SEC) regulations designed to give shareholders adequate time to review proposals before voting them, while controlling for time-varying confounding factors at the firm, industry, or campaign topic levels. Therefore, while current events or news breaks may influence an NGO’s decision to launch a campaign, they are exogenous to the AGM calendar. We employ these external factors shifting NGO’s objectives to identify the trade-off between influence and visibility and to study how their objectives shift as they accumulate reputation or brand capital (Hasan et al., 2022).
Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
We interpret our results through the lens of firm life cycle models, where consumers initially lack information about a firm’s product quality, spurring firms to invest in reputation, especially when firms are young (e.g., Wernerfelt, 1988; Cabral, 2000, 2009). In this context, waiting for the target firm’s AGM may be seen as a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny and that is salient for shareholders.
Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, allowing them to build their reputation. Once established, they can leverage this reputation to directly impact firms’ decisions, engage in negotiations with firms, or spend their image to directly support shareholder proposals.
Taken together, our results describe a novel trade-off between exposure and (immediate) impact as a key driver of the societal consequences of NGO activism. Unlike our paper, these activists are often assumed to pursue social objectives, which may stem from prosocial behavior as described by Bénabou and Tirole (2006) or from moral preferences, according to Baron (2010). Their level of commitment can also depend on how closely they align with the cause, as noted by Besley and Ghatak (2005). Baron (2010) further illustrates that NGOs can emerge when socially motivated individuals coordinate efforts, drawing on frameworks outlined by Dixit (2004) and Tabellini (2008). When firms become targets of activist campaigns, firms may engage in “self-regulation” to preempt opposition and reduce potential harm. As firms’ capture of regulators has become an unavoidable feature of our economies, NGO activism surges to rebalance public and private interests (Glaeser and Shleifer, 2003; Egorov and Harstad, 2017; Daubanes and Rochet, 2019). [^2] A related literature highlights several instances in which NGOs can become subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021). For example, firms may strategically exploit their relationships with NGOs—through donations or partnerships—to further political agendas (Bertrand et al., 2020) or to regain consumer and investor trust after reputational damages (Akey et al., 2021).
This paper also contributes to the growing body of literature on firm objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (Hawn and Ioannou, 2016) . [^3] A recent literature also suggests that government initiatives might even hinder private ones (Akey et al., 2024). The literature emphasizes that shareholders can influence firms through either voicing concerns or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024)–especially through voting (Cuñat et al., 2016; Gine et al., 2017)–while consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010). However, other agents may play a pivotal role. Among them, NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Fioretti et al., 2023; Chuah et al., 2024; Koenig and Poncet, 2019). This paper sheds new light on NGO as a coordination mechanism across stakeholders by providing new insights into the functioning of NGOs and how stakeholders respond to their activism.
Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the reputations of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024). In contrast, our study focuses on how NGOs can enhance their own reputation by adequately timing their campaigns against misbehaving corporations.
The paper is organized as follows. Section 2 outlines the data construction and provides summary statistics. The empirical approach is described in Section 3. Sections 4 and 5 examine media exposure and influence at the current AGM as potential drivers of AGM actions. Section 6 explores the role of NGO reputation in supporting related shareholder proposals. Section 7 discusses the results and Section 8 concludes.
2 Data and Summary Statistics
Our dataset integrates three sources. First, we use data from Sigwatch, a European consultancy tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^4] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, MacDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common topics are Environment/Pollution, Human & Labor Rights, Animal Rights, and Health.
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the largest (by market capitalization) 3,000 US-listed firms’ Annual General Meetings (AGMs). For those firms, ISS data include detailed information on shareholder proposals for these firms, such as proposals withdrawn before the AGM.
We select environmental and social proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
Lastly, we source firm-level news data from Ravenpack. RavenPack is a data analytics platform that transforms unstructured news content into structured data, offering sentiment analysis and event detection. The data allows us to count the news related to every firm headquartered in the US. [^5] We only retain news with a Relevance Score of at least 75%, which ensures the relevance of the news for the firm. We include only US-listed firms from 2010 to 2020. The sample includes 5,212 US firms.
We link these three datasets using standard identifiers, including year, firms’ CUSIP, and ISIN numbers. Appendix Table A1 presents summary statistics for NGO-related variables (Panel a). On average, NGOs launch 3.40 campaigns during the sample period, but the distribution is skewed, with over 50% of NGOs initiating only one campaign. Most campaigns focus on social issues rather than environmental ones. Panel (b) displays the variation in key variables for the targeted firms, including market capitalization, assets, and news coverage.
The following three sections examine the payoffs accruing to an NGO in connection with the timing of the campaigns it launched, as well as the reactions to the campaigns of central stakeholders such as consumers and shareholders. We present the key takeaways from each analysis at the end of each section. The empirical approach is discussed in detail next.
3 Empirical Strategy
Our empirical approach investigates the time gap between a firm’s AGM and the actions taken by NGOs or other stakeholders, while controlling for time-varying confounding factors at the firm, industry, and campaign topic levels. Importantly, our identification strategy relies on the independence of the corporate calendar from events that may motivate NGOs to launch campaigns. The timing of AGMs is largely exogenous, as SEC regulations mandate a stable schedule across years. Specifically, Rule 14a-8 ensures shareholders have ample time to review proposals before voting, requiring that shareholder proposals be filed at least 120 days before the proxy statement is released, based on the previous year’s AGM timing. This stabilizes AGM dates across consecutive years (Glac, 2014), as shown in Appendix Figure B3, which plots AGM months across adjacent years. Most variation outside the diagonal line—indicating significant fluctuations in AGM months—is attributed to the pandemic, as the SEC allowed firms to adjust AGM dates in 2020 to prevent COVID-19 spread and facilitate shareholder attendance (SEC, 2020). Our results remain robust even when excluding data from 2020 and 2021.
While current events may influence an NGO’s decision to launch a campaign against a firm, they are unlikely to affect the timing of AGMs. These external factors may shift an NGO’s objectives by altering incentives to target a firm. Our analysis leverages these shifts to identify a key trade-off for NGOs: whether to influence a firm’s strategies by launching campaigns early or to gain greater visibility by timing campaigns around the firm’s AGM.
Our results highlight NGO experience as a crucial factor in this trade-off. Before interpreting our findings, we present data in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past. Panel (c) reveals that both small and large NGOs—in terms of their number of campaigns in the sample—target firms with similar market capitalization and total assets. They also have similar numbers of yearly news mentions and environmental and social scores, with comparable counts of environmental and social campaigns. Panel (d) examines target selection by regressing each one of these firm-level variables against the number of past campaigns launched by the campaigning NGO, excluding (Column 1) or including (Column 2) fixed effects at the firm-by-year, NGO-by-year, and topic-by-year levels. This correlation analysis shows that NGOs’ campaign targets do not change as they gain experience, although they are more likely to focus on environmental issues rather than social ones in later years as climate change grew in importance.
4 Timing Campaigns for Media Exposure
We start our analysis by investigating when NGOs are most likely to launch their campaigns relative to a firm’s corporate calendar. Figure 1(a) shows that NGOs are significantly more likely to launch campaigns on a firm’s AGM date. Campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100 campaigns. [^6] Appendix Figure B2 confirms the result showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$), following Cattaneo et al. (2020).
Figure 1: Timing of NGO campaigns and media attention around AGMs
(a) Timing of NGO campaigns
(b) Media coverage in the AGM week (gray)
Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,489 NGO campaigns. Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms. The NGO data come from Sigwatch; the news data are from Ravenpack. Calendar weeks are re-centered around the firm’s AGM.
This rise could be related to the heightened media coverage during AGMs (see also Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars). To test if media exposure of NGO campaigns is highest around AGMs, we aggregate news by firm ($n$) and month-by-year ($m,y$) and run the following event study:
$\text{News}_{n,m,y}$ $=\sum^{\tau=6}_{\begin{subarray}{c}\tau=-6\\ \tau\neq-3\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{m=\tau\}}\cdot\text{NGO Campaign}_{n,m,y}\right)+\mathbf{X}_{n,m,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{m,y,s(n)}+u_{n,m,y}.$ (1)
First, note that we adjusted the calendar year to align with the corporate year so that the year starts six months before the AGM month (month 0) and ends six months later for each firm. Importantly, since firms are unlikely to change their AGM month over time due to SEC regulations as outlined in Section 3, moving from the calendar year to the corporate year does not affect the distribution of months by firms (i.e., months in following years are not double counted). Thus, the recalendarization is a mere relabeling of the time variation within a firm. [^7] In each year, we define months -6 and +6 as half months, so the calendar year has twelve full months.
Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise. We use $\tau=-3$ as the reference period because shareholders typically receive voting materials within 90 days of the AGM, marking a date when the AGM’s visibility in the news could increase (ISS, 2022).
The vector $\mathbf{X}_{n,m,y}$ captures the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable. Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-month-by-year fixed effects ($\rho_{m,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs). Standard errors are clustered at the firm level.
Figure 2 reports the estimated coefficients, revealing that media coverage peaks during the month of the AGM. To quantify its magnitude, an NGO campaign launched in the AGM month increases media coverage by 80% compared to a campaign started 6 months earlier, on average. After the AGM, coverage returns to average levels, indicating that NGOs gain more media attention by campaigning close to the AGM date.
Result 1.
Media coverage of NGO campaigns is substantially greater in the month of the AGM.
Figure 2: Media exposure of NGO campaigns around an AGM
Note: We regress the number of news about firm $f$ in month $t$ on a dummy indicating the publication of a negative NGO report about firm $f$ in month $t$, a dummy equal to one when month $t$ is $k$ months away from the nearest AGM, where we take $k$ to vary between -6 and +6 months, the interaction between these two dummy variables, and firm and month by industry fixed effects. Month -6 and +6 are “half” months, so the year still consists of 12 full months. The figure also reports the 95% confidence interval when the standard errors are clustered at the level of the firm. The sample period goes from January 2010 to December 2019 and covers 4,676 firms. NGO campaign data from Sigwatch. AGM data from ISS. News data from Ravenpack. Industry data from Compustat. The regression also includes firm and year-by-industry fixed effects. Standard errors are clustered at the firm level.
4.1 Stakeholders’ Reactions
In this section, we explore how campaign timing affects critical stakeholders of the targeted firms, such as shareholders and consumers. Therefore, timing a campaign on an AGM date has significant implications with consequences that go beyond media coverage. We report the exhibits related to the analysis in this section in the Online Appendix.
4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same topic of the campaign in the previous year. Appendix C describes the matching procedure.
For each campaign in our sample, we look at the probability that the next AGM has a related shareholder proposal, depending on the timing of the campaign:
[…]
Figure B4 shows that only NGO campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same topic at the AGM. This substantial increase is as large as 12 percentage points, or 44%, above the average probability of observing a proposal. Thus, NGO campaigns from the previous AGMs likely lead to a shareholder proposal at the current AGM, potentially due to heightened media attention.
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately affect a firm’s behavior, without having to wait a year to advocate for change at the next AGM.
4.1.2 Consumers
Drawing from quarterly data from Compustat, we investigate the impact of NGO campaigns on consumers as the stakeholder consuming a firm’s values (Conway and Boxell, 2024) by studying how the Sales to Asset ratio changed around the time of an NGO campaign to the AGM date through
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value rebases a firm’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms). That is, we investigate potential consumer boycotts in the months following an NGO campaign. Standard errors are clustered at the firm level.
Appendix Figure B5 reports the estimated $\beta_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles). We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so two quarters after a campaign targeted on the AGM day, implying a loss of sales to total assets of 0.7pp, or about 5%. Similar sale losses as a result of negative ESG corporate events have been documented in the literature (Conway and Boxell, 2024; Saint-Jean, 2024).
Result 2.
NGO campaigns on an AGM date impact both shareholders, who are more likely to make related proposals in future years, and consumers, who are more likely to boycott the firm in the following months.
5 Timing Campaigns to Influence AGMs
Early campaigns could benefit an NGO aiming to influence current AGM outcomes. Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes. For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting (Microsoft, 2023), suggesting that a campaign close to the AGM date may have less impact on the voting results.
We define a proposal as successful if either it is withdrawn by the shareholder before the AGM or it secures at least $50\%+1$ vote at the AGM. We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+u_{i,n,m,y},$ (4)
where $\text{Success}_{i,n,m,y}$ is a dummy indicating the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+l,t}$ is a dummy indicating the occurrence of an NGO campaign on the same topic as proposal $i$, $l$ months before the AGM. The reference period covers the months that occur more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year and topic-by-year fixed effects, thereby exploiting only variation across campaigns and topics within firms. The standard errors are clustered by firm.
Figure 3(a) plots the estimated $\hat{\beta}_{l}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more than seven months before.
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (e.g., Shu, 2024). Those recommendations have a critical influence on the vote decisions. We thus include management and ISS’s recommendations as covariates in (4).
Figure 3(b) shows that only NGO campaigns occurring a month or three months before the AGM increases support for related shareholder proposals. This increase is by nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this translates to a 8% increase in shareholder support when an NGO campaign addresses the same topic, on average. In contrast to Figure 3(a), there is no significant increase in support for campaigns initiated as early as six months before the AGM. This suggests that the difference between the two figures is due to shareholder proposals being withdrawn after reaching an agreement with the firm.
Figure 3: NGOs’ influence on the vote of E&S shareholder proposals
(a) Proposals accepted or withdrawn
(b) Shareholder support for proposal
Note: E&S denote environmental and social proposals. The figure plots the estimated coefficients of regressions of a dummy indicating that a shareholder proposal is withdrawn or passes if it goes to a vote (left panel, based on 911 proposals), or the percentage of votes in favor of the proposal when it goes to the vote (right panel, based on 2,078 proposals), on dummies indicating the occurrence of an NGO campaign on the same topic $k$ months before the AGM, for $k$ varying from -6 to 0. The regressions also include firm and year-by-industry fixed effects. Standard errors are clustered at the firm level.
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). For this reason, we investigate all E&S proposals in Figure 3(a). Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms (Levit and Malenko, 2011) as large shareholder support to a cause may be hard to conceal for managers.
Result 3.
NGO campaigns are more effective at influencing a firm’s decisions when launched shortly before the AGM.
6 Exposure vs. Influence
Therefore, Sections 4 and 5 outline a tradeoff: late campaigns impact society by boosting an NGO’s visibility among the media and key stakeholders, while early ones are more effective at influencing a firm’s decisions. How do NGOs navigate this tradeoff?
In this section, we first explore whether NGOs, after gaining visibility by targeting AGMs in the past, switch to campaigning before the AGM. We then examine whether this “visibility capital” from previous campaigns increases the likelihood of success at the AGM for shareholder proposals related to the NGO campaign.
6.1 Campaign Timing and Accumulated NGO Visibility
Section 4 suggests that targeting a firm’s AGM increases an NGO’s visibility in the media and among stakeholders. This visibility could enhance an NGO’s reputation for commitment to a cause. In this section, we explore whether NGOs, after establishing such a reputation, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
We investigate this mechanism by:
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise. Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation, we include NGO-by-year fixed effects. The regression also includes firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. The standard errors are clustered at the NGO level. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample.
To interpret the estimated $\beta_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^8] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year.
Alternative mechanism—learning. This result is unlikely due to NGOs learning that campaigning on the AGM date does not influence voting decisions, as most shareholders have already voted by proxy. Appendix Figure B6 shows the same downward trend observed in Figure 4 when subsetting to NGOs whose first campaign took place within 2 weeks after the AGM, when the NGO surely knew that it could not affect votes.
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time, thereby excluding selection on observables. For instance, Appendix Figures B7 and B8 show that NGO experience does not influence the choice of targets in terms of size (market capitalization), and Environmental and Social scores for Environmental and Social campaigns, respectively.
Result 4.
NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of the votes.
Figure 4: NGO experience and the timing of NGO campaigns
Note: This figure plots the estimated coefficients from the regression of the distance between NGO campaigns and the targeted firm’s next AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. Regressions are based on a sample of 1,136 NGO campaigns. Standard errors are clustered at the NGO level.
6.2 NGO Success and Reputation
We investigate whether NGO reputation, as proxied by past campaigns on AGM dates, makes NGOs more effective at producing positive corporate change. We assess this channel with the following regression:
$\text{Success}_{i,n,y}$ $=\beta_{1}\text{Campaign on Same Topic}_{c(i,n,g),y}+\beta_{2}\text{Experience}_{g,y}$ (6)
$+\beta_{3}\text{Campaign on Same Topic}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+u_{i,n,y},$
where $\text{Success}_{i,n,y}$ is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, $\text{Campaign on Same Topic}_{i,n,y}$ is one if there has been a campaign by NGO $g$ against firm $n$ on topic $i$ in the ten months preceding the AGM and zero otherwise, and $\text{Experience}_{g,y}$ indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the firm-by-year, ngo-by-year, and topic-by-year through fixed effects. Standard errors are clustered by firm.
Table 1: NGO experience at targeting AGMs and success at supporting current shareholder proposals
Dependent Variable: Proposal Passes (0/1)
Model: (1) (2) (3) (4)
Main Variables
Campaign on Same Topic 0.018 0.083∗∗ 0.046 0.076
(0.068) (0.040) (0.059) (0.042)
Experience 0.018
Campaign on Same Topic $\times$ Experience 0.024
AGM Experience (0/1) -0.036 -0.035
(0.072) (0.071)
Campaign on Same Topic $\times$ AGM Experience (0/1) 0.141∗∗ 0.126
(0.070) (0.070)
Non-AGM Experience (0/1) 0.005
Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.054
AGM Experience -0.019
Squared AGM Experience 0.002∗∗
Campaign on Same Topic $\times$ AGM Experience 0.068∗∗
Campaign on Same Topic $\times$ Squared AGM Experience -0.005∗∗∗
Fixed-effects
Firm x Year
Topic x Year
NGO x Year
Fit statistics
Observations 38,613 38,613 38,613 38,613
R2 0.31881 0.31820 0.31844 0.32041
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
  • Note: This table reports the estimated coefficients from the regression of a dummy indicating whether an E&S shareholder proposal is withdrawn before voting or passes after voting (Success) on a dummy indicating the occurrence of an NGO campaign on the same topic as the proposal in the 10 months preceding the AGM (Campaign on Same Topic), a dummy indicating whether the NGO has launched a campaign on a firm’s AGM day in the past (AGM Experience (0/1)), and their interaction. In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic. In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic. All regressions also include Firm x Year, NGO x Year, and Topic × Year fixed effects. Clustered (firm) standard errors are reported between parentheses.
Column (1) of Table 1 reports the estimates for $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same topic as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a topic being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns. We replace $\text{Experience}_{g,y}$ for the variable $\text{AGM Experience}_{g,y}$, which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise. The interaction coefficient is significant and economically important. While a campaign on the same proposal topic increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
[…]
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces $\text{Experience}_{g,y}$ for the cumulative count of the campaigns when NGO $g$ targeted an AGM date ($\text{AGM Experience}_{g,y}$) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and Campaign on Same Topic imply a concave relationship between AGM experience and AGM success. The results are consistent with Figure 4, which shows that only the very first campaigns happen close to an AGM date.
Result 5.
Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but display decreasing returns.
7 Discussion
Our results demonstrate that NGOs strategically use the timing of their campaigns to influence firms’ decisions. In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaign against a target firm’s practice or strategy (Result 1). By doing so, NGOs prompt reactions from key stakeholders, such as shareholders, who are more likely to file a related proposal at the following AGM, and consumers, as the likelihood of boycotts increases if NGOs campaign on the AGM date (Result 2). The goal of these campaigns is to alter a firm’s strategy by damaging its reputation (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
However, campaigning before the AGM offers direct advantages. Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3). Since shareholder proposals generally receive low support, early NGO campaigns can play a role in disciplining managers (Levit and Malenko, 2011). Importantly, these analyses control for NGO-by-year fixed effects and topic-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., WWF) do not solely drive these results. Also, we provide evidence suggesting that NGOs campaigns on AGMs do not target different firms compared to other campaign times in Appendix Table A2.
Therefore, NGOs face a trade-off, which they navigate by adapting strategies over time. NGOs with limited activism history tend to campaign on the AGM date. These NGOs may favor early campaigns to build a reputation, as without it, they may be unable to influence shareholder votes. In contrast, more established NGOs prefer targeting firms just before their AGM to directly affect shareholder votes (Result 4), consistent with reputation as a signaling device.
The public may also expect mature NGOs not only to generate negative publicity but also to produce immediate results. Thus, there are diminishing returns to reputation from consistently targeting AGMs. We show that this practice enables NGOs with a history of targeting AGMs to help related shareholder proposals pass. However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns, decreases support for shareholder proposals (Result 5).
In conclusion, NGOs play a crucial role in increasing information and aligning corporate interests with societal needs, and their growth in recent years is well-documented (e.g., Egorov and Harstad, 2017; Daubanes and Rochet, 2019). However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited because of limited information on NGOs. This paper offers a detailed analysis of NGO campaigns to uncover these trade-offs and incentives by leveraging the exogeneity of the corporate calendar to NGO’s goals. Similar approaches have been successfully used to identify the preferences of other stakeholders and may help reframe the role of AGMs not just as voting venues but also as forums where firms engage with and negotiate alongside outside stakeholders in ways that are not often observed by researchers but can be inferred from the underlying system of incentives (Fioretti et al., 2023). We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to that employed in this paper to quantitatively study the incentives for collaboration.
By focusing on the timing of NGO campaigns, we demonstrate how this seemingly innocuous decision has significant implications for both targeted firms and the success of future initiatives. We show that, similar to for-profit organizations, NGOs are influenced by their reputation, which affects the timing of the information presented to stakeholders and, consequently, their responses. Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize their externalities.
8 Conclusion
This paper examines the timing and motives behind NGO campaigns, revealing that these campaigns are not randomly timed but often strategically target firms’ Annual General Meetings. While this timing does not affect votes on environmental and social proposals, which are cast beforehand, it enhances the NGO’s visibility, leading to greater consumer reactions and increasing the likelihood of related shareholder proposals in the following year. Mature NGOs utilize the reputation similarly built over the years to influence current firm decisions by supporting relevant shareholder proposals. Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its lifecycle, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
Appendix A Omitted Tables
Table A1: Summary Statistics
Panel a: Variables varying by NGO
Mean 25% 50% 75% 95%
Number of campaigns by NGO 3.20 1 1 3 13
      – E campaigns 1.51 0 1 1 5
      – S campaigns 1.70 0 1 1 6
      – On AGM date 0.24 0 0 0 1
Campaign sentiment -1.72 -2 -2 -1 -1
Panel b: Variables varying by target firm
Mean 25% 50% 75% SD
Market Cap (USD bn) 175.66 21.71 75.05 209.47 263.66
Total Assets (USD bn) 98.81 21.80 59.27 138.68 124.36
Yearly Sales (USD bn) 37.26 6.85 17.35 42.41 150.58
Yearly Number of News 1,183 328 616 1,046 1,695.25
E score 5.66 4 5.58 7.3 2.22
S score 4.44 3.30 4.70 5.60 1.58
E campaign 0.47 0 0 1 0.50
S campaign 0.53 0 1 1 0.50
Panel c: Variables varying by target firm and NGO experience
High Experience Low Experience Test difference
Mean S.E. Mean S.E. p-value
Market Cap (decile) 9.23 0.10 9.07 0.04 0.13
Total Assets (decile) 9.26 0.08 9.23 0.04 0.72
Sales (decile) 9.31 0.09 9.14 0.04 0.07
Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
E score 5.64 0.12 5.72 0.06 0.58
S score 4.44 0.09 4.45 0.04 0.89
E campaign 0.46 0.03 0.5 0.01 0.15
S campaign 0.54 0.03 0.49 0.01 0.12
Panel d: OLS Regression of $y_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
Dependent Variable ($y_{ny}$): (1) (2) (3) (4)
Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
Total Assets Decile 0.003 (0.004) -0.002 (0.012) 1,338
Sales Decile 0.001 (0.004) -0.002 (0.03) 1,337
Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
E score 0.004 (0.009) 0.041 (0.041) 1,338
S score -0.0006 (0.006) 0.019 (0.032) 1,338
E campaign 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
S campaign -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
  • Note: This table reports summary statistics on NGO campaigns (Panel a, NGO campaign data from Sigwatch and AGM dates from ISS), firms targeted by the campaigns (Panel b, accounting data from Compustat, ESG scores from MSCI, and news count based on RavenPack data), as well as an analysis of the relationship between NGO experience and targeted firms’ characteristics (Panels c and d). In Panel c, the Market Cap, Total Assets, and Sales deciles are based on the universe of US firms in Compustat on the quarter of the campaign. We use deciles instead of dollar values because highly ranked firms saw a large expansion in market cap and assets over the last years. Ranking measures like deciles avoid such a mechanical correlation with NGO Experience, which, being defined as the number of campaigns launched by NGO g in previous years, also grows over time. In Panel d, the fixed-effects are: NGO×$\times$×Year, Firm×$\times$×Year and Topic×$\times$×Year, except for the last two rows where Topic×$\times$×Year fixed effects are not included because of perfect collinearity with the dependent variables. Standard errors are clustered by firm.
Table A2: Campaign timing and target selection
Market Cap Assets Sales Yearly Environmental Social Campaign
Dependent Variables: (decile) (decile) (decile) News Score Score Sentiment
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Campaign on AGM day (0/1) 0.469 0.495 0.374 83.2 0.469 -0.522 0.022
(0.628) (0.695) (0.550) (185.7) (0.919) (0.715) (0.383)
Fixed-effects
Industry-Year
NGO-Year
Topic-Year
Fit statistics
Observations 1,338 1,338 1,337 1,299 1,338 1,338 1,338
R2 0.93364 0.92669 0.90388 0.99232 0.95420 0.92167 0.84525
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
  • Note: This table reports the estimated coefficients from regressions of different target firm characteristics on a dummy indicating whether the campaign took place on the target firm’s AGM day. All regressions also include Industry-by-Year, NGO-by-Year, and Topic-by-Year fixed effects. The Market Cap, Assets, and Sales deciles are based on the entire universe of US firms in Compustat in the quarter of the campaign. Yearly News refers to the four-quarter moving average of news mentioning the target firm. Accounting data from Compustat. NGO campaign (including Sentiment) data from Sigwatch. News data from RavenPack. Environmental and Social scores are from MSCI. Standard errors are clustered by firm.
Appendix B Omitted Figures
Figure B1: E&S shareholder proposals, and their success
Note: This figure plots the number of E&S proposals at Russell 3,000 firms from 2006 to 2021 (solid line) and the number of successful E&S proposals (blue area) in the same period. A successful proposal is defined as a proposal that is either withdrawn before the AGM, or that goes to vote and passes. Data from ISS. Based on 5,937 shareholder proposals filed at 979 firms.
Figure B2: Testing NGO campaign timing around the AGM date
Note: This figure plots the density test for the occurrence of NGO campaigns around firms’ AGM, following Cattaneo et al. (2020). The test is based on 521 NGO campaigns launched within 50 days of the firm’s closest AGM. The NGO data come from Sigwatch; the AGM dates are from ISS.
Figure B3: Persistence of the AGM month over time
Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis). A larger and darker circle indicates a greater proportion of meetings. Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
Figure B4: NGO campaigns and shareholder proposals
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same topic, $k$ months ahead. The regression also includes month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
Figure B5: NGO campaigns and consumer reactions
Note: This figure plots the estimated coefficients of the regression of sales over total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$. The regressions also include firm and quarter-by-industry fixed effects. The number of observations is 351,351 and 10,188, respectively. Standard errors are clustered at the firm level.
Figure B6: Campaign timing for NGOs whose first campaign is after the AGM
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
Figure B7: NGO experience and target firm’s size
Note: This figure plots the estimated coefficients from the regressions of on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. Standard errors are clustered at the firm level. The number of observations is 1,338.
Figure B8: NGO experience and target firm’s E&S characteristics
(a) Target firm’s E score for Environmental campaigns
(b) Target firm’s S score for Social campaigns
Note: This figure plots the estimated coefficients from the regressions of on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.
Appendix C Matching Proposals And NGO Campaigns
This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign topic.
Sigwatch assigns a“Generic Topic” to each campaign in their dataset. To streamline the analysis, we combine certain topics when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category. Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” topic, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
Using these consolidated “Generic Topic Names”, we then categorize each shareholder proposal based on its short description. When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for topic classification. We start by assigning proposals to topics based on keywords in their descriptions and then manually classify any remaining proposals.
Table C1: Proposal Topics
NGO Campaign Category Example of Shareholder Proposals
Topic Classification (E or S)
Climate Change E Encourage and Support Carbon Tax to Address Climate Change
Environment E Stop financing mountain top removal coal mining
Pollution E Report on Efforts to Reduce Pollution From Products and Operations
Rainforest E Report on impact on rainforest
Sustainability E Establish board committee on sustainability
Waste Generation E Do not increase radioactive waste production
Water E Report on supply chain water impacts
Animal Rights & Welfare S Review suppliers’ animal slaughter methods
CSR S Disclose Charitable Contributions
Food Safety S Label Genetically Modified Ingredients
Health S Reduce Tobacco Harm to Health
Human & Labor Rights S End Sri Lanka Trade Partnerships Until Human Rights Improve
Indigenous People S Adopt Global Policy Regarding the Rights of Indigenous People
Intellectual Property Rights S Implement a Process to Fully Vet Whether a
Supervisor is a Bona Fide Inventor Before a Patent is Filed
Politics S Affirm Political Nonpartisanship
Product Safety S Report on Nanomaterial Product Safety
Worker Safety S Report on Reducing Risks of Accidents
  • Note: This Table reports the 17 topics assigned to each shareholder proposal in the sample. Shareholder proposals data from ISS. The topics are from Sigwatch’s categorization of campaigns.

### v2

@@
 We analyze the timing of NGO campaigns to shed light on NGOs’ objectives and how their strategies evolve over time.
-Using data from 2,500 campaigns, we find that NGOs are eight times more likely to launch campaigns on Annual General Meeting (AGM) dates.
-These campaigns increase media exposure and stakeholder scrutiny, raising the likelihood of consumer boycotts and related shareholder proposals at the following AGM of the targeted firm but have no impact on current AGM votes.
-As NGOs build reputational capital, they adjust their timing to influence AGM outcomes, revealing the trade-offs they face to drive corporate change.
+Using data from 2,500 campaigns, we find that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date.
+Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer boycotts and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes.
+As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
 1 Introduction
 Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels.
-Through tactics of “private politics”—such as boycotts and shaming campaigns—NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017).
-Many studies examine how campaigns by both large and small NGOs impact and delegitimate targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009).
+Through tactics of “private politics,” such as boycotts and shaming campaigns, NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017).
+Many studies examine how campaigns by both large and small NGOs impact targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009).
 For instance, just weeks before McDonald’s 2006 Annual General Meeting (AGM), Greenpeace launched a campaign branding the company ‘McAmazon‘ for alleged links to Amazon deforestation, prompting McDonald’s to adopt a new sustainable sourcing policy within days (Goldberg and Droste Yagan, 2007).
-Similar campaigns have targeted companies across sectors, including Nike, Shell, Monsanto, and BNP Paribas.
-However, we know relatively little about NGOs’ objectives and trade-offs and how they vary over their life cycle, shaping their campaign strategies toward misbehaving firms and, ultimately, their capacity to drive societal change.
-To illustrate how objectives vary over time and in the life of an NGO, our analysis spans a wide range of organizations, from smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club.‘ However, we face several challenges, particularly due to the scarcity of data on NGOs.
-Donation data for non-political nonprofits is difficult to access because of donor privacy concerns, [^1] Data exist for donations to political parties and charities thanks to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023).
-This paper differs from this literature in that it does not focus on political donations or the resulting campaigns, but rather on the incentives that NGOs face when deciding to campaign against a firm.
-and NGOs have much more lenient transparency requirements than publicly listed corporations, complicating the tracking of their fund usage.
-The absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot be easily reduced to profit maximization (Maxwell, 2012; Fioretti, 2022).
-The literature emphasizing extreme behaviors of non-profits, with some contribute to the greater good (e.g., Daubanes and Rochet, 2019) whereas others are captured by corporate interests (e.g., Bertrand et al., 2021).
-Therefore, this paper conducts a forensic investigation that infers NGOs’ objectives from their reactions to external incentives, focusing on the primary—and often only—observable data on NGOs: their campaigns, which must necessarily be public and observable to influence firms and their stakeholders.
+Similar campaigns have targeted companies across various sectors, including Nike, Shell, Monsanto, and BNP Paribas.
+Despite the significance of these campaigns in shaping NGO strategies and influencing corporate behavior, we know relatively little about NGOs’ objectives, trade-offs, and how these evolve over the course of an NGO’s life.
+Researchers addressing these questions face several challenges, particularly due to the scarcity of data on NGOs.
+For instance, donation data for non-political nonprofits is difficult to access because of donor privacy concerns.
+[^1] Data on donations to political parties and charities are more readily available due to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023).
+This paper differs from this literature in that it does not focus on the connection between NGO and politics, but rather on the incentives that NGOs face when deciding to campaign against a firm.
+NGOs are also subject to much less stringent transparency requirements than other firms such as publicly listed corporations, which complicates the tracking of their fund usage.
+Moreover, the absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot easily be reduced to profit maximization (Maxwell, 2012; Fioretti, 2022).
+To obviate to this problem, the literature emphasizes extreme behaviors of NGOs, with some contributing to the greater good (e.g., Daubanes and Rochet, 2019), while others are captured by corporate interests (e.g., Bertrand et al., 2021, 2023).
 The starting point of our analysis is that NGO campaigns are strategically timed.
-Two key factors influence an NGO’s ability to capture stakeholders’ attention.
-First, media coverage amplifies NGO campaigns, making shaming and boycotting efforts more effective.
-Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from workers, consumers, shareholders, suppliers, and the public.
-We demonstrate that these two factors create a trade-off: NGOs must choose between seeking influence or enhancing their reputation.
-We find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM.
+Drawing on campaign data from a wide range of NGOs—spanning smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club‘—we identify two key factors that influence an NGO’s ability to capture stakeholders’ attention.
+First, media coverage amplifies NGO campaigns, making efforts like shaming and boycotting more effective.
+[^2] Consistent with this, Hawn (2021) finds that media coverage of a firm’s irresponsible behavior can significantly amplify negative public perceptions (Shea and Hawn, 2019), potentially resulting in the failure of planned merger.
+Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from stakeholders such as workers, consumers, shareholders, suppliers, and the public.
+We demonstrate that these two factors create a trade-off regarding the strategic decision of campaign timing.
+NGOs must choose between maximizing influence on firms and enhancing their reputation.
+Employing over ten years of campaign data against the largest US corporations, we find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM.
 The AGM attracts the highest media coverage, as media outlets engage with managers and key shareholders to discuss the firm’s current achievements and future outlook (Dyck et al., 2008).
-This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, which might proxy for the NGO’s reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017).
-However, a late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020).
+This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, thereby accumulating reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017).
+Since our estimation routine compares the timing of the campaigns of NGOs campaigning against multiple firms in a year, we can account for any confounding factors varying within an NGO (and target firms) across years through fixed effects, such as their size.
+A late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020).
 Nevertheless, we find that the heightened visibility of campaigns timed with the AGM increases the likelihood of (i) a shareholder proposal on a related topic the following year and (ii) a boycott campaign in the subsequent months, compared to campaigns launched at other times.
-In contrast, campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
-We argue that NGOs navigate this tradeoff differently depending on their accummulated reputation.
-Employing over ten years of campaign data against the largest US corporations, we first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly.
+Campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
+We argue that NGOs navigate this tradeoff differently depending on their reputation.
+We first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly.
 Our analysis rules out learning this aspect of proxy voting as an explanation for our finding, as we observe the same trend for NGOs whose first campaign occurred just after an AGM, meaning they already knew they could not influence the vote by campaigning that late.
-We also find no evidence that NGO selects different target firms as they gain more campaigning experience.
+We also find no evidence that NGOs select different target firms as they gain more campaigning experience.
 Second, we demonstrate that the success of a shareholder proposal at a current AGM is linked to how many times the NGO backing the proposal has campaigned on AGM dates in the past.
-We interpret this as a reduced-form proxy for the reputation the NGO has built over time.
-However, there are diminishing returns to targeting AGM dates: once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes by campaigning earlier.
-Conversely, we find that accumulating experience at campaigning outside AGMs does not help NGOs at swaying shareholder proposal votes.
-Empirically, our approach leverages the exogenous timing of AGM meetings, which in the U.S.
-remains relatively stable within firms over time due to Securities and Exchange Commission (SEC) regulations designed to give shareholders adequate time to review proposals before voting them, while controlling for time-varying confounding factors at the firm, industry, or campaign topic levels.
-Therefore, while current events or news breaks may influence an NGO’s decision to launch a campaign, they are exogenous to the AGM calendar.
-We employ these external factors shifting NGO’s objectives to identify the trade-off between influence and visibility and to study how their objectives shift as they accumulate reputation or brand capital (Hasan et al., 2022).
-Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
+We interpret this as a reduced-form proxy for the reputation built by the NGO for fighting over issues.
+As we document diminishing returns to targeting AGM dates, once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes through early campaigns.
+Conversely, we find that accumulating experience at campaigning outside AGM dates does not help NGOs at swaying shareholder proposal votes.
+Empirically, our approach leverages exogenous variation in the timing of AGM meetings, which in the U.S.
+we show to be relatively stable within firms over time due to regulations by the Securities and Exchange Commission (SEC) designed to give shareholders adequate time to review proposals before voting them.
+Across analyses, our approach also allows us to exclude time-varying confounding factors at the firm or industry, and campaign topic levels.
+In particular, while current events and news may influence an NGO’s decision to launch a campaign, they cannot influence the AGM calendar.
+As a result, objectives are identified by the tradeoff between targeting a firm’s AGM date to increase their exposure and impacting the firm’s decisions by campaigning early in support of the news or event.
+Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type (reputation)—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
 We interpret our results through the lens of firm life cycle models, where consumers initially lack information about a firm’s product quality, spurring firms to invest in reputation, especially when firms are young (e.g., Wernerfelt, 1988; Cabral, 2000, 2009).
-In this context, waiting for the target firm’s AGM may be seen as a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny and that is salient for shareholders.
-Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, allowing them to build their reputation.
+In this context, waiting for the target firm’s AGM is a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny at a time that is salient for shareholders and decision makers.
+Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, reflecting a better reputation.
 Once established, they can leverage this reputation to directly impact firms’ decisions, engage in negotiations with firms, or spend their image to directly support shareholder proposals.
@@
 For example, firms may strategically exploit their relationships with NGOs—through donations or partnerships—to further political agendas (Bertrand et al., 2020) or to regain consumer and investor trust after reputational damages (Akey et al., 2021).
-This paper also contributes to the growing body of literature on firm objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024).
+This paper also contributes to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; George et al., 2023; Fangwa et al., 2024).
 Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities.
-However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (Hawn and Ioannou, 2016) .
+However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (e.g., Hawn and Ioannou, 2016; Hartzmark and Sussman, 2019; Hasan et al., 2022).
 [^4] A recent literature also suggests that government initiatives might even hinder private ones (Akey et al., 2024).
-The literature emphasizes that shareholders can influence firms through either voicing concerns or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024)–especially through voting (Cuñat et al., 2016; Gine et al., 2017)–while consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010).
+The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016; Gine et al., 2017)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010).
 However, other agents may play a pivotal role.
 Among them, NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Fioretti et al., 2023; Chuah et al., 2024; Koenig and Poncet, 2019).
-This paper sheds new light on NGO as a coordination mechanism across stakeholders by providing new insights into the functioning of NGOs and how stakeholders respond to their activism.
-Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the reputations of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024).
-In contrast, our study focuses on how NGOs can enhance their own reputation by adequately timing their campaigns against misbehaving corporations.
+Therefore, this short paper sheds new light on NGOs as coordinating various stakeholders such as shareholders and consumers, empirically highlighting a set of tradeoffs that arise from stakeholder involvement (Porter and Kramer, 2011; Kaplan, 2020).
+We provide new insights into the incentives NGOs face and how they can help align a firm’s action with the interests of its stakeholders (McGahan and Pongeluppe, 2023).
+In this direction, the results also suggest that other stakeholders can gain privately (i.e., reputation) by engaging with a firm, bringing empirical evidence to theories suggesting that stakeholders would not contribute resources to an organization without the potential to claim at least some of the value created (cf., Barney, 2018; McGahan, 2023).
+Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the performances and prosocial outcomes of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024).
+Our study turns the table by focusing on the incentives that non-profit firms face when campaigning against for-profit firms.
 The paper is organized as follows.
@@
 Our dataset integrates three sources.
-First, we use data from Sigwatch, a European consultancy tracking global NGO campaigns against corporations (Hatte and Koenig, 2020).
+First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020).
 The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative.
@@
 The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace.
-The most targeted companies are Monsanto, MacDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil.
+The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil.
 The most common topics are Environment/Pollution, Human & Labor Rights, Animal Rights, and Health.
 Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the largest (by market capitalization) 3,000 US-listed firms’ Annual General Meetings (AGMs).
-For those firms, ISS data include detailed information on shareholder proposals for these firms, such as proposals withdrawn before the AGM.
+For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
 We select environmental and social proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total).
@@
 We present the key takeaways from each analysis at the end of each section.
-The empirical approach is discussed in detail next.
+Next, we discuss the empirical approach in detail.
 3 Empirical Strategy
@@
 Our results highlight NGO experience as a crucial factor in this trade-off.
-Before interpreting our findings, we present data in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past.
+Before interpreting our findings, we present results in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past.
 Panel (c) reveals that both small and large NGOs—in terms of their number of campaigns in the sample—target firms with similar market capitalization and total assets.
-They also have similar numbers of yearly news mentions and environmental and social scores, with comparable counts of environmental and social campaigns.
+Targets also have similar numbers of yearly news mentions, environmental and social scores, and suffered a similar number of environmental and social campaigns.
 Panel (d) examines target selection by regressing each one of these firm-level variables against the number of past campaigns launched by the campaigning NGO, excluding (Column 1) or including (Column 2) fixed effects at the firm-by-year, NGO-by-year, and topic-by-year levels.
-This correlation analysis shows that NGOs’ campaign targets do not change as they gain experience, although they are more likely to focus on environmental issues rather than social ones in later years as climate change grew in importance.
+This correlation analysis shows that NGOs’ campaign targets do not change as NGOs gain experience, although they are more likely to focus on environmental issues rather than social ones in later years due to the increasing prominence of environmental challenges over social ones.
 4 Timing Campaigns for Media Exposure
@@
 [^8] In each year, we define months -6 and +6 as half months, so the calendar year has twelve full months.
-Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise.
+Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with the variable $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise.
 We use $\tau=-3$ as the reference period because shareholders typically receive voting materials within 90 days of the AGM, marking a date when the AGM’s visibility in the news could increase (ISS, 2022).
-The vector $\mathbf{X}_{n,m,y}$ captures the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable.
+The vector $\mathbf{X}_{n,m,y}$ contains the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable.
 Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-month-by-year fixed effects ($\rho_{m,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs).
@@
 Result 1.
-Media coverage of NGO campaigns is substantially greater in the month of the AGM.
+Media coverage of NGO campaigns is substantially greater during the target’s AGM.
 Figure 2: Media exposure of NGO campaigns around an AGM
@@
 Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome.
-If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately affect a firm’s behavior, without having to wait a year to advocate for change at the next AGM.
+If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately gain broad support on the shareholder’s proposal, without having to wait a year to advocate for change at the next AGM.
 4.1.2 Consumers
@@
 where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively.
-The dependent value rebases a firm’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms).
+The dependent value rebases $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms).
 That is, we investigate potential consumer boycotts in the months following an NGO campaign.
@@
 Result 2.
-NGO campaigns on an AGM date impact both shareholders, who are more likely to make related proposals in future years, and consumers, who are more likely to boycott the firm in the following months.
+NGO campaigns on an AGM date impact stakeholders.
+(i) Shareholders are more likely to make related proposals in future years.
+(ii) Consumers are more likely to boycott the firm in the following months.
 5 Timing Campaigns to Influence AGMs
@@
 We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
-$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+u_{i,n,m,y},$ (4)
+$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
 where $\text{Success}_{i,n,m,y}$ is a dummy indicating the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+l,t}$ is a dummy indicating the occurrence of an NGO campaign on the same topic as proposal $i$, $l$ months before the AGM.
 The reference period covers the months that occur more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$.
-We account for firm-by-year and topic-by-year fixed effects, thereby exploiting only variation across campaigns and topics within firms.
+We account for firm-by-year, and topic-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and topics within firm-year pairs.
 The standard errors are clustered by firm.
@@
 Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”).
-Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (e.g., Shu, 2024).
+Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (Shu, 2024).
 Those recommendations have a critical influence on the vote decisions.
@@
 The figure plots the estimated coefficients of regressions of a dummy indicating that a shareholder proposal is withdrawn or passes if it goes to a vote (left panel, based on 911 proposals), or the percentage of votes in favor of the proposal when it goes to the vote (right panel, based on 2,078 proposals), on dummies indicating the occurrence of an NGO campaign on the same topic $k$ months before the AGM, for $k$ varying from -6 to 0.
-The regressions also include firm and year-by-industry fixed effects.
+The regressions also include month, firm-by-year, and topic-by-year fixed effects.
 Standard errors are clustered at the firm level.
@@
 We investigate this mechanism by:
-$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
-which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
+$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
+which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{c,g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
 Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010.
-To control for this truncation, we include NGO-by-year fixed effects.
-The regression also includes firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign.
+To control for this truncation as well as time varying confounders—e.g., the size of the NGO social media network across years (Tufekci, 2017)—we include NGO-by-year fixed effects.
+We also include firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign.
+Hence, the identification of $\beta_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^9] In Appendix Figure B6, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
+Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample.
 The standard errors are clustered at the NGO level.
-Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample.
 To interpret the estimated $\beta_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM.
@@
 [^10] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year.
+We postpone the interpretation of this result to Section 7, which sums up and discuss the results.
 Alternative mechanism—learning.
@@
 Alternative mechanism—selection.
-NGO do not increasingly launch campaigns against firms that may be more easily influenced over time, thereby excluding selection on observables.
-For instance, Appendix Figures B7 and B8 show that NGO experience does not influence the choice of targets in terms of size (market capitalization), and Environmental and Social scores for Environmental and Social campaigns, respectively.
+NGO do not increasingly launch campaigns against firms that may be more easily influenced over time.
+Appendix Figures B8 and B9 exclude selection on observables.
+For instance, target sizes do not covary (market capitalization) with NGO experience.
+In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
 Result 4.
@@
 $+\beta_{3}\text{Campaign on Same Topic}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
-$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+u_{i,n,y},$
+$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
 where $\text{Success}_{i,n,y}$ is defined as in (4)—namely, either a successful vote or a withdrawn proposal.
 On the right-hand side, $\text{Campaign on Same Topic}_{i,n,y}$ is one if there has been a campaign by NGO $g$ against firm $n$ on topic $i$ in the ten months preceding the AGM and zero otherwise, and $\text{Experience}_{g,y}$ indicates the number of campaigns NGO $g$ had launched in the past.
-We control for all confounding variations at the firm-by-year, ngo-by-year, and topic-by-year through fixed effects.
+We control for all confounding variations at the month, firm-by-year, NGO-by-year, and topic-by-year through fixed effects.
 Standard errors are clustered by firm.
@@
 Main Variables
-Campaign on Same Topic 0.018 0.083∗∗ 0.046 0.076
+Campaign on Same Topic 0.008 0.076 0.032 0.070
 (0.064) (0.039) (0.056) (0.041)
 Experience 0.018
-Campaign on Same Topic $\times$ Experience 0.024
-AGM Experience (0/1) -0.036 -0.035
+Campaign on Same Topic $\times$ Experience 0.026
+AGM Experience (0/1) -0.042 -0.041
 (0.070) (0.069)
-Campaign on Same Topic $\times$ AGM Experience (0/1) 0.141∗∗ 0.126
+Campaign on Same Topic $\times$ AGM Experience (0/1) 0.169∗∗ 0.150∗∗
 (0.074) (0.073)
 Non-AGM Experience (0/1) 0.005
-Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.054
-AGM Experience -0.019
+Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.065
+AGM Experience -0.018
 Squared AGM Experience 0.002∗∗
-Campaign on Same Topic $\times$ AGM Experience 0.068∗∗
+Campaign on Same Topic $\times$ AGM Experience 0.074∗∗
 Campaign on Same Topic $\times$ Squared AGM Experience -0.005∗∗∗
@@
 NGO x Year
+Month
 Fit statistics
 Observations 38,613 38,613 38,613 38,613
-R2 0.31881 0.31820 0.31844 0.32041
+R2 0.32100 0.32065 0.32101 0.32260
 *–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
@@
 Clustered (firm) standard errors are reported between parentheses.
-Column (1) of Table 1 reports the estimates for $\{\beta_{1},\beta_{2},\beta_{3}\}$.
+Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$.
 Campaigning on the same topic as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero.
@@
 Our results demonstrate that NGOs strategically use the timing of their campaigns to influence firms’ decisions.
-In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaign against a target firm’s practice or strategy (Result 1).
+In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaigns against target firms’ practices or strategies (Result 1).
 By doing so, NGOs prompt reactions from key stakeholders, such as shareholders, who are more likely to file a related proposal at the following AGM, and consumers, as the likelihood of boycotts increases if NGOs campaign on the AGM date (Result 2).
-The goal of these campaigns is to alter a firm’s strategy by damaging its reputation (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
+The goal of these campaigns is to alter a firm’s strategy by damaging its reputation and showing stakeholders’ disappointment (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
 However, campaigning before the AGM offers direct advantages.
@@
 Also, we provide evidence suggesting that NGOs campaigns on AGMs do not target different firms compared to other campaign times in Appendix Table A2.
-Therefore, NGOs face a trade-off, which they navigate by adapting strategies over time.
-NGOs with limited activism history tend to campaign on the AGM date.
-These NGOs may favor early campaigns to build a reputation, as without it, they may be unable to influence shareholder votes.
-In contrast, more established NGOs prefer targeting firms just before their AGM to directly affect shareholder votes (Result 4), consistent with reputation as a signaling device.
+Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time.
+NGOs with limited activism histories tend to campaign on the AGM date, while more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
+This pattern is consistent with the idea of using campaigning time as a signaling device.
+One interpretation of this result is that, as NGOs gain more experience with campaigning—and potentially run multiple campaigns in a year—they accumulate exposure through standard news media and may already have a large social media following.
+As a result, there may be less pressure to campaign specifically on the AGM date.
+Another interpretation is that, as NGOs become more complex organizations over time, media exposure becomes less vital for influencing decision-making at target firms.
+Disentangling these interpretations suggest several potential avenues for future research, contributing to the broader conversation on civil society lobbying and activism (Bertrand et al., 2021, 2023).
 The public may also expect mature NGOs not only to generate negative publicity but also to produce immediate results.
-Thus, there are diminishing returns to reputation from consistently targeting AGMs.
+Thus, there may be diminishing returns to reputation from consistently targeting AGMs.
 We show that this practice enables NGOs with a history of targeting AGMs to help related shareholder proposals pass.
-However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns, decreases support for shareholder proposals (Result 5).
+However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns on average, decreases support for shareholder proposals (Result 5).
 In conclusion, NGOs play a crucial role in increasing information and aligning corporate interests with societal needs, and their growth in recent years is well-documented (e.g., Egorov and Harstad, 2017; Daubanes and Rochet, 2019).
-However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited because of limited information on NGOs.
+However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited due to the scant data available on NGOs and models to interpret their actions.
 This paper offers a detailed analysis of NGO campaigns to uncover these trade-offs and incentives by leveraging the exogeneity of the corporate calendar to NGO’s goals.
@@
 We show that, similar to for-profit organizations, NGOs are influenced by their reputation, which affects the timing of the information presented to stakeholders and, consequently, their responses.
-Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize their externalities.
+Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize firms’ externalities.
 8 Conclusion
@@
 Mature NGOs utilize the reputation similarly built over the years to influence current firm decisions by supporting relevant shareholder proposals.
-Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its lifecycle, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
+Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its life, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
 Appendix A Omitted Tables
@@
 Standard errors are clustered by firm.
-Table A2: Campaign timing and target selection
+Table A2: Campaign timing and characteristics of target firms
+Characteristics of Target Firm
 Market Cap Assets Sales Yearly Environmental Social Campaign
@@
 Based on 5,937 shareholder proposals filed at 979 firms.
+Figure B2: Persistence of the AGM month over time
+Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis).
+A larger and darker circle indicates a greater proportion of meetings.
+Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
 Figure B3: Testing NGO campaign timing around the AGM date
@@
 The NGO data come from Sigwatch; the AGM dates are from ISS.
-Figure B3: Persistence of the AGM month over time
-Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis).
-A larger and darker circle indicates a greater proportion of meetings.
-Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
 Figure B4: NGO campaigns and shareholder proposals
@@
 Standard errors are clustered at the firm level.
+Figure B6: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects
+Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
+The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year.
+Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in NGO experience over years.
+Standard errors are clustered at the NGO level.
 Figure B7: Campaign timing for NGOs whose first campaign is after the AGM
-Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
+Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
 The panel focuses on NGOs whose first campaign was within the two weeks following the AGM.
@@
 Figure B8: NGO experience and target firm’s size
-Note: This figure plots the estimated coefficients from the regressions of on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
+Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the experience of the NGO, measured by the number of campaigns it had launched in previous year as in (5).
 Standard errors are clustered at the firm level.
@@
 (b) Target firm’s S score for Social campaigns
-Note: This figure plots the estimated coefficients from the regressions of on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
+Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating the experience of the NGO, measured by either the number of environmental (top panel) or social (bottom panel) campaigns it had launched in previous year as in (5).
 Standard errors are clustered at the firm level.

### v3

@@
 Abstract
-We analyze the timing of NGO campaigns to shed light on NGOs’ objectives and how their strategies evolve over time.
-Using data from 2,500 campaigns, we find that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date.
-Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer boycotts and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes.
+We analyze the timing of NGO campaigns to shed light on NGOs’ strategies and how they evolve over time.
+Data from 2,500 campaigns show that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date.
+Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer backlash and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes.
 As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
+JEL classifications: L21, L31
+Keywords: Non-Governmental Organizations, campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
 1 Introduction
-Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels.
-Through tactics of “private politics,” such as boycotts and shaming campaigns, NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017).
-Many studies examine how campaigns by both large and small NGOs impact targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009).
-For instance, just weeks before McDonald’s 2006 Annual General Meeting (AGM), Greenpeace launched a campaign branding the company ‘McAmazon‘ for alleged links to Amazon deforestation, prompting McDonald’s to adopt a new sustainable sourcing policy within days (Goldberg and Droste Yagan, 2007).
-Similar campaigns have targeted companies across various sectors, including Nike, Shell, Monsanto, and BNP Paribas.
-Despite the significance of these campaigns in shaping NGO strategies and influencing corporate behavior, we know relatively little about NGOs’ objectives, trade-offs, and how these evolve over the course of an NGO’s life.
-Researchers addressing these questions face several challenges, particularly due to the scarcity of data on NGOs.
-For instance, donation data for non-political nonprofits is difficult to access because of donor privacy concerns.
-[^1] Data on donations to political parties and charities are more readily available due to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023).
-This paper differs from this literature in that it does not focus on the connection between NGO and politics, but rather on the incentives that NGOs face when deciding to campaign against a firm.
-NGOs are also subject to much less stringent transparency requirements than other firms such as publicly listed corporations, which complicates the tracking of their fund usage.
-Moreover, the absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot easily be reduced to profit maximization (Maxwell, 2012; Fioretti, 2022).
-To obviate to this problem, the literature emphasizes extreme behaviors of NGOs, with some contributing to the greater good (e.g., Daubanes and Rochet, 2019), while others are captured by corporate interests (e.g., Bertrand et al., 2021, 2023).
-The starting point of our analysis is that NGO campaigns are strategically timed.
-Drawing on campaign data from a wide range of NGOs—spanning smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club‘—we identify two key factors that influence an NGO’s ability to capture stakeholders’ attention.
-First, media coverage amplifies NGO campaigns, making efforts like shaming and boycotting more effective.
-[^2] Consistent with this, Hawn (2021) finds that media coverage of a firm’s irresponsible behavior can significantly amplify negative public perceptions (Shea and Hawn, 2019), potentially resulting in the failure of planned merger.
-Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from stakeholders such as workers, consumers, shareholders, suppliers, and the public.
-We demonstrate that these two factors create a trade-off regarding the strategic decision of campaign timing.
-NGOs must choose between maximizing influence on firms and enhancing their reputation.
-Employing over ten years of campaign data against the largest US corporations, we find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM.
-The AGM attracts the highest media coverage, as media outlets engage with managers and key shareholders to discuss the firm’s current achievements and future outlook (Dyck et al., 2008).
-This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, thereby accumulating reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017).
-Since our estimation routine compares the timing of the campaigns of NGOs campaigning against multiple firms in a year, we can account for any confounding factors varying within an NGO (and target firms) across years through fixed effects, such as their size.
-A late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020).
-Nevertheless, we find that the heightened visibility of campaigns timed with the AGM increases the likelihood of (i) a shareholder proposal on a related topic the following year and (ii) a boycott campaign in the subsequent months, compared to campaigns launched at other times.
-Campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
-We argue that NGOs navigate this tradeoff differently depending on their reputation.
-We first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly.
-Our analysis rules out learning this aspect of proxy voting as an explanation for our finding, as we observe the same trend for NGOs whose first campaign occurred just after an AGM, meaning they already knew they could not influence the vote by campaigning that late.
-We also find no evidence that NGOs select different target firms as they gain more campaigning experience.
-Second, we demonstrate that the success of a shareholder proposal at a current AGM is linked to how many times the NGO backing the proposal has campaigned on AGM dates in the past.
-We interpret this as a reduced-form proxy for the reputation built by the NGO for fighting over issues.
-As we document diminishing returns to targeting AGM dates, once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes through early campaigns.
-Conversely, we find that accumulating experience at campaigning outside AGM dates does not help NGOs at swaying shareholder proposal votes.
-Empirically, our approach leverages exogenous variation in the timing of AGM meetings, which in the U.S.
-we show to be relatively stable within firms over time due to regulations by the Securities and Exchange Commission (SEC) designed to give shareholders adequate time to review proposals before voting them.
-Across analyses, our approach also allows us to exclude time-varying confounding factors at the firm or industry, and campaign topic levels.
-In particular, while current events and news may influence an NGO’s decision to launch a campaign, they cannot influence the AGM calendar.
-As a result, objectives are identified by the tradeoff between targeting a firm’s AGM date to increase their exposure and impacting the firm’s decisions by campaigning early in support of the news or event.
-Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type (reputation)—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
-We interpret our results through the lens of firm life cycle models, where consumers initially lack information about a firm’s product quality, spurring firms to invest in reputation, especially when firms are young (e.g., Wernerfelt, 1988; Cabral, 2000, 2009).
-In this context, waiting for the target firm’s AGM is a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny at a time that is salient for shareholders and decision makers.
-Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, reflecting a better reputation.
-Once established, they can leverage this reputation to directly impact firms’ decisions, engage in negotiations with firms, or spend their image to directly support shareholder proposals.
-Taken together, our results describe a novel trade-off between exposure and (immediate) impact as a key driver of the societal consequences of NGO activism.
-Unlike our paper, these activists are often assumed to pursue social objectives, which may stem from prosocial behavior as described by Bénabou and Tirole (2006) or from moral preferences, according to Baron (2010).
-Their level of commitment can also depend on how closely they align with the cause, as noted by Besley and Ghatak (2005).
-Baron (2010) further illustrates that NGOs can emerge when socially motivated individuals coordinate efforts, drawing on frameworks outlined by Dixit (2004) and Tabellini (2008).
-When firms become targets of activist campaigns, firms may engage in “self-regulation” to preempt opposition and reduce potential harm.
-As firms’ capture of regulators has become an unavoidable feature of our economies, NGO activism surges to rebalance public and private interests (Glaeser and Shleifer, 2003; Egorov and Harstad, 2017; Daubanes and Rochet, 2019).
-[^3] A related literature highlights several instances in which NGOs can become subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021).
-For example, firms may strategically exploit their relationships with NGOs—through donations or partnerships—to further political agendas (Bertrand et al., 2020) or to regain consumer and investor trust after reputational damages (Akey et al., 2021).
-This paper also contributes to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; George et al., 2023; Fangwa et al., 2024).
+Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017).
+However, their actions are often difficult to observe, and evidence on the incentives they face and how these shape their strategic decisions remains limited (Yaziji and Doh, 2009).
+Thus, NGOs are frequently perceived as passive actors, either selecting the most favorable action from a common good perspective (Daubanes and Rochet, 2019) or being dysfunctional and susceptible to corporate capture (Bertrand et al., 2020).
+This paper introduces a novel approach to analyzing how NGOs make decisions with implications for corporations and their stakeholders.
+Consider an NGO with information about corporate wrongdoing that faces a strategic choice: launch an immediate campaign to foster dialogue or wait for the optimal moment when the firm is most sensitive to external pressure.
+While early action encourages engagement and potential change, strategic timing can amplify impact by signaling strong opposition among stakeholders.
+A time when firms are particularly sensitive to external pressure is during their Annual General Meetings (AGMs), recurring shareholder meetings where strategic decisions are revealed and executives are held accountable.
+We show that NGOs are four to six times more likely to campaign on a target’s AGM date than at other times, a pattern driven by the heightened media coverage AGMs attract.
+This visibility helps NGOs reach stakeholders and the broader public while imposing reputational costs on firms.
+However, while AGMs provide a platform for NGOs to maximize exposure, their direct influence is limited—most shareholders vote before the meeting via proxy, reducing the immediate impact on voting outcomes (Bebchuk et al., 2020).
+Empirically, we leverage the exogenous timing of AGMs, which remains stable within firms due to SEC regulations, allowing us to focus on NGOs’ actions while holding firms’ strategic choices constant.
+Our findings reveal the “life cycle” of NGO strategies (Wernerfelt, 1988; Cabral, 2000).
+Early-stage NGOs prioritize visibility by campaigning on AGM dates, despite their limited direct influence.
+We show that AGM-timed campaigns prompt consumer backlash and increase the likelihood of related shareholder proposals the following year.
+However, while these campaigns generate exposure, they do not directly drive change, as, for instance, the target’s environmental and social scores remain unchanged.
+Over time, NGOs targeting AGMs start campaigning before AGMs to maximize direct influence on corporate decisions.
+Early campaigns increase the likelihood of related shareholder proposals succeeding, ultimately driving measurable improvements in firms’ environmental and social performance.
+Our interpretation is that AGM exposure can enhance NGO reputation.
+Supporting this view, we find that the success of a shareholder proposal on a given cause increases when an NGO campaigns on the same issue and year, with past AGM-date campaigns by the NGO amplifying this effect.
+However, the returns to AGM-date campaigns diminish as their frequency rises.
+In contrast, past NGO campaigns on non-AGM dates have no impact.
+We view AGM-date campaigns as a way for NGOs to signal their type to stakeholders and build their “reputation,” as evidenced by the diminishing returns of AGM-targeting over time, which is consistent with reputation models (Bar-Isaac and Tadelis, 2008).
+[^1] This practice also generates additional negative news to a target firm only for an NGO’s first few campaigns.
+Later campaigns inflict no additional damage.
+Thus, the NGO may be willing to incur a cost to build its “reputation-for-toughness,” which it can then leverage in future campaigns (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
+Hence, more mature NGOs shift to campaign before AGMs.
+Furthermore, NGO reputation is likely multidimensional, shaped both by a sustained track record of activism and by a demonstrated commitment to a cause.
+As such, NGOs may transition to early campaigning not only due to diminishing returns on visibility but also to establish credibility through constructive engagement with key stakeholders (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
+By comparing NGO campaign outcomes across firms and issues within years, our analysis accounts for factors that vary across years within NGOs, firms, and issues, such as organizational competence, learning, moral authority, resources, and the salience of specific issues.
+NGOs also engage firms through partnerships and other channels, but since such activities are voluntarily disclosed—and thus subject to selection biases (Burger and Owens, 2010)—we rather focus on campaign timing, which must be public to be effective and, thus, observable to researchers.
+Our findings open new avenues for research, as future work may empirically explore how NGOs choose between campaigning and other activist tactics or whether these different approaches act as complements or substitutes.
+The exogenous timing of AGMs provides a useful identification strategy for such analyses, as it holds the target firm’s strategic choices constant, allowing researchers to isolate NGOs’ choices and their incentives.
+Our results reveal a novel tradeoff between exposure and direct impact as a key driver of the societal consequences of NGO activism (Briscoe and Gupta, 2016).
+Unlike our framework, much of the existing literature assumes that activists primarily pursue social objectives, either stemming from prosocial behavior (Besley and Ghatak, 2005; Bénabou and Tirole, 2006) or moral preferences (Baron, 2010), without discussing their incentives.
+At the same time, several studies highlight instances in which NGOs themselves are subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021).
+Firms may also strategically leverage their relationships with NGOs—through donations or partnerships—to advance political agendas (Bertrand et al., 2020), or to restore consumer and investor trust following reputational damage (Akey et al., 2021).
+Our main contribution in this literature is to show that NGOs are not passive actors.
+Rather, their tactics evolve over time, shaping firms and their stakeholders’ responses in different ways depending on the strategies they employ (e.g., Hendricks and Sorensen, 2009; Cabral and Hortacsu, 2010).
+Thus, understanding NGOs’ incentives is essential to learn why firms internalize their externalities.
+We also contribute to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024).
 Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities.
-However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (e.g., Hawn and Ioannou, 2016; Hartzmark and Sussman, 2019; Hasan et al., 2022).
-[^4] A recent literature also suggests that government initiatives might even hinder private ones (Akey et al., 2024).
-The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016; Gine et al., 2017)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010).
+However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear (e.g., Hawn and Ioannou, 2016).
+The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010).
 However, other agents may play a pivotal role.
-Among them, NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Fioretti et al., 2023; Chuah et al., 2024; Koenig and Poncet, 2019).
-Therefore, this short paper sheds new light on NGOs as coordinating various stakeholders such as shareholders and consumers, empirically highlighting a set of tradeoffs that arise from stakeholder involvement (Porter and Kramer, 2011; Kaplan, 2020).
-We provide new insights into the incentives NGOs face and how they can help align a firm’s action with the interests of its stakeholders (McGahan and Pongeluppe, 2023).
-In this direction, the results also suggest that other stakeholders can gain privately (i.e., reputation) by engaging with a firm, bringing empirical evidence to theories suggesting that stakeholders would not contribute resources to an organization without the potential to claim at least some of the value created (cf., Barney, 2018; McGahan, 2023).
-Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the performances and prosocial outcomes of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024).
-Our study turns the table by focusing on the incentives that non-profit firms face when campaigning against for-profit firms.
+Among them, we highlight that NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Koenig and Poncet, 2019; Fioretti et al., 2023; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023).
 The paper is organized as follows.
@@
 Section 6 explores the role of NGO reputation in supporting related shareholder proposals.
-Section 7 discusses the results and Section 8 concludes.
-2 Data and Summary Statistics
+Section 7 discusses the results and concludes.
+2 Data
 Our dataset integrates three sources.
@@
 The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil.
-The most common topics are Environment/Pollution, Human & Labor Rights, Animal Rights, and Health.
-Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the largest (by market capitalization) 3,000 US-listed firms’ Annual General Meetings (AGMs).
+The most common causes are “environment/pollution,” “human & labor rights,” “animal rights,” and “health.”
+Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms by market capitalization.
 For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
-We select environmental and social proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total).
+We select environmental and social (E&S) proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total).
 Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote.
-Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023).
+[^3] Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023).
 We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021.
@@
 Panel (b) displays the variation in key variables for the targeted firms, including market capitalization, assets, and news coverage.
-The following three sections examine the payoffs accruing to an NGO in connection with the timing of the campaigns it launched, as well as the reactions to the campaigns of central stakeholders such as consumers and shareholders.
+The following three sections examine an NGO’s payoffs in connection with the timing of its campaigns, and the reactions of central stakeholders.
 We present the key takeaways from each analysis at the end of each section.
-Next, we discuss the empirical approach in detail.
-3 Empirical Strategy
-Our empirical approach investigates the time gap between a firm’s AGM and the actions taken by NGOs or other stakeholders, while controlling for time-varying confounding factors at the firm, industry, and campaign topic levels.
-Importantly, our identification strategy relies on the independence of the corporate calendar from events that may motivate NGOs to launch campaigns.
-The timing of AGMs is largely exogenous, as SEC regulations mandate a stable schedule across years.
-Specifically, Rule 14a-8 ensures shareholders have ample time to review proposals before voting, requiring that shareholder proposals be filed at least 120 days before the proxy statement is released, based on the previous year’s AGM timing.
-This stabilizes AGM dates across consecutive years (Glac, 2014), as shown in Appendix Figure B2, which plots AGM months across adjacent years.
-Most variation outside the diagonal line—indicating significant fluctuations in AGM months—is attributed to the pandemic, as the SEC allowed firms to adjust AGM dates in 2020 to prevent COVID-19 spread and facilitate shareholder attendance (SEC, 2020).
-Our results remain robust even when excluding data from 2020 and 2021.
-While current events may influence an NGO’s decision to launch a campaign against a firm, they are unlikely to affect the timing of AGMs.
-These external factors may shift an NGO’s objectives by altering incentives to target a firm.
-Our analysis leverages these shifts to identify a key trade-off for NGOs: whether to influence a firm’s strategies by launching campaigns early or to gain greater visibility by timing campaigns around the firm’s AGM.
-Our results highlight NGO experience as a crucial factor in this trade-off.
-Before interpreting our findings, we present results in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past.
-Panel (c) reveals that both small and large NGOs—in terms of their number of campaigns in the sample—target firms with similar market capitalization and total assets.
-Targets also have similar numbers of yearly news mentions, environmental and social scores, and suffered a similar number of environmental and social campaigns.
-Panel (d) examines target selection by regressing each one of these firm-level variables against the number of past campaigns launched by the campaigning NGO, excluding (Column 1) or including (Column 2) fixed effects at the firm-by-year, NGO-by-year, and topic-by-year levels.
-This correlation analysis shows that NGOs’ campaign targets do not change as NGOs gain experience, although they are more likely to focus on environmental issues rather than social ones in later years due to the increasing prominence of environmental challenges over social ones.
+Before that, we detail our empirical approach.
+3 Exploiting AGM Dates
+Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations.
+Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014).
+Appendix Figure B2 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs.
+[^5] Results remain robust excluding the covid period (2020–2021).
+While external events may influence NGOs’ campaign decisions, they do not affect AGM timing.
+Thus, AGMs allow us to isolate NGOs’ strategic trade-off from firms’ responses.
+Target selection.
+Our analysis focuses on campaign timing, abstracting from NGOs’ selection of a target, which we leave for future research.
+However, Appendix Table A1 shows that NGOs’ targets does not depend on prior campaign experience, a key variable of our analysis.
+Panel (c) finds no differences in market capitalization, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs.
+Panel (d) regresses firm characteristics on NGOs’ past campaigns, showing no correlation between target selection and NGO experience.
+However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.
 4 Timing Campaigns for Media Exposure
-We start our analysis by investigating when NGOs are most likely to launch their campaigns relative to a firm’s corporate calendar.
-Figure 1(a) shows that NGOs are significantly more likely to launch campaigns on a firm’s AGM date.
-Campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100 campaigns.
-[^7] Appendix Figure B3 confirms the result showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$), following Cattaneo et al.
-(2020).
+NGOs are substantially more likely to launch their campaigns around AGM dates, on average.
+According to Figure 1(a), NGO campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100.666Appendix Figure B3 confirms this finding showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$) (Cattaneo et al., 2020).
 Figure 1: Timing of NGO campaigns and media attention around AGMs
-(a) Timing of NGO campaigns
-(b) Media coverage in the AGM week (gray)
+(a) Campaigns peak during AGMs
+(b) Media coverage rises during AGMs
 Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,489 NGO campaigns.
@@
 Calendar weeks are re-centered around the firm’s AGM.
-This rise could be related to the heightened media coverage during AGMs (see also Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars).
-To test if media exposure of NGO campaigns is highest around AGMs, we aggregate news by firm ($n$) and month-by-year ($m,y$) and run the following event study:
-$\text{News}_{n,m,y}$ $=\sum^{\tau=6}_{\begin{subarray}{c}\tau=-6\\ \tau\neq-3\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{m=\tau\}}\cdot\text{NGO Campaign}_{n,m,y}\right)+\mathbf{X}_{n,m,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{m,y,s(n)}+u_{n,m,y}.$ (1)
-First, note that we adjusted the calendar year to align with the corporate year so that the year starts six months before the AGM month (month 0) and ends six months later for each firm.
-Importantly, since firms are unlikely to change their AGM month over time due to SEC regulations as outlined in Section 3, moving from the calendar year to the corporate year does not affect the distribution of months by firms (i.e., months in following years are not double counted).
-Thus, the recalendarization is a mere relabeling of the time variation within a firm.
-[^8] In each year, we define months -6 and +6 as half months, so the calendar year has twelve full months.
-Second, Equation (1) interacts the event-study dummy, $\mathbf{1}_{\{m=\tau\}}$, with the variable $\text{NGO Campaign}_{n,m,y}$, which is 1 if firm $n$ experienced an NGO campaign in month $m$ of year $y$, and 0 otherwise.
-We use $\tau=-3$ as the reference period because shareholders typically receive voting materials within 90 days of the AGM, marking a date when the AGM’s visibility in the news could increase (ISS, 2022).
-The vector $\mathbf{X}_{n,m,y}$ contains the direct effects of $\mathbf{1}_{\{m=\tau\}}$ and $\text{NGO Campaign}_{n,m,y}$ on the dependent variable.
-Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-month-by-year fixed effects ($\rho_{m,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs).
+This finding could be related to the heightened media coverage during AGMs (Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars).
+To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
+$\text{News}_{n,w,y}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}{\{M_{n,w,y}=\tau\}}\,\cdot\,\text{NGO Campaign}_{n,w,y}\right)$ (1)
+$+\mathbf{X}_{n,w,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y}\,\,,$
+where “$\text{News}_{n,w,y}$” is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event.
+The independent variables include a set of dummies “$\mathbf{1}\{M_{n,w,y}=\tau\}$” indicating the distance in months from the closest AGM.
+The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise.777 The analysis aligns the calendar year with the corporate year, starting five months before the AGM month (month 0) and ending five months later, with months -5 and +5 slightly longer.
+Since SEC regulations make AGM months stable over time (Section 3), this relabeling does not alter the distribution of months by firms or double-count adjacent years.
+We interact these event-study indicators, with the variable “$\text{NGO Campaign}_{n,m,y}$,” an indicator variable equal to 1 if firm $n$ experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise.
+We set the reference month to four months before the AGM, following SEC rules that require the AGM date to be set at least four months in advance.
+Figure 2: Firms’ negative media coverage increases if NGOs campaign during AGMs
+Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Standard errors are clustered at the firm level.
+The vector “$\mathbf{X}_{n,w,y}$” includes “$\mathbf{1}{\{M_{n,w,y}=\tau\}}$” and “$\text{NGO Campaign}_{n,m,y}$” separately, and controls for (i) the occurrence of negative news and (ii) all E&S news over the previous four weeks to account for a firm’s recent media coverage.
+[^8] Adding controls for the current count of E&S news to account for NGOs targeting AGMs of high-coverage firms does not qualitatively change the results.
+Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-week-by-year fixed effects ($\rho_{w,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs).
 Standard errors are clustered at the firm level.
-Figure 2 reports the estimated coefficients, revealing that media coverage peaks during the month of the AGM.
-To quantify its magnitude, an NGO campaign launched in the AGM month increases media coverage by 80% compared to a campaign started 6 months earlier, on average.
-After the AGM, coverage returns to average levels, indicating that NGOs gain more media attention by campaigning close to the AGM date.
+Figure 2 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients.
+Negative news follow a bell-shaped pattern around the AGM, peaking significantly above zero in the AGM month, highlighting the benefits of targeting AGMs.
+[^9] In section 6.2, we show that the spike on AGM date is particularly strong for an NGO’s first few campaigns but it fades away for later ones (Appendix Figure B4).
+To quantify the magnitude of the AGM peak, an NGO campaign launched in the AGM week doubles the probability of a negative E&S news compared to a campaign initiated four months earlier, on average.
+After the AGM, coverage returns to baseline levels, indicating that NGOs attract more media attention when campaigning closer to the AGM.
 Result 1.
 Media coverage of NGO campaigns is substantially greater during the target’s AGM.
-Figure 2: Media exposure of NGO campaigns around an AGM
-Note: We regress the number of news about firm $f$ in month $t$ on a dummy indicating the publication of a negative NGO report about firm $f$ in month $t$, a dummy equal to one when month $t$ is $k$ months away from the nearest AGM, where we take $k$ to vary between -6 and +6 months, the interaction between these two dummy variables, and firm and month by industry fixed effects.
-Month -6 and +6 are “half” months, so the year still consists of 12 full months.
-The figure also reports the 95% confidence interval when the standard errors are clustered at the level of the firm.
-The sample period goes from January 2010 to December 2019 and covers 4,676 firms.
-NGO campaign data from Sigwatch.
-AGM data from ISS.
-News data from Ravenpack.
-Industry data from Compustat.
-The regression also includes firm and year-by-industry fixed effects.
-Standard errors are clustered at the firm level.
+Importantly, we rule out the hypothesis that NGOs select more negative causes when targeting AGMs.
+Without an adequate ranking of causes by negativity, we estimate (1) with firm $n$’s average E&S news sentiment in week $w$ as the dependent variable.
+[^10] RavenPack scores the tone of a firm’s news coverage between -1 (adverse) and 1 (favorable).
+Examples of positive and negative news can be found in Appendix Figure B5.
+The estimated coefficients show that campaigns launched during AGM months are no different than other campaigns (Appendix Figure B6).
 4.1 Stakeholders’ Reactions
-In this section, we explore how campaign timing affects critical stakeholders of the targeted firms, such as shareholders and consumers.
-Therefore, timing a campaign on an AGM date has significant implications with consequences that go beyond media coverage.
-We report the exhibits related to the analysis in this section in the Online Appendix.
+This section explores how campaign timing affects critical stakeholders of a targeted firm, such as shareholders and consumers, while Section 4.2 examines its ramifications for the target’s operations.
+The relevant exhibits are in the Online Appendix.
 4.1.1 Shareholders
@@
 Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest.
-We define as “related” a shareholder proposal on the same topic of the campaign in the previous year.
+We define as “related” a shareholder proposal on the same cause of the campaign in the previous year.
 Appendix C describes the matching procedure.
@@
 […]
+$+\alpha_{i,s(n),g,y}+u_{i,n,y}$
 […]
-Figure B4 shows that only NGO campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same topic at the AGM.
+Figure B7 shows that only campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same cause.
 This substantial increase is as large as 12 percentage points, or 44%, above the average probability of observing a proposal.
@@
 4.1.2 Consumers
-Drawing from quarterly data from Compustat, we investigate the impact of NGO campaigns on consumers as the stakeholder consuming a firm’s values (Conway and Boxell, 2024) by studying how the Sales to Asset ratio changed around the time of an NGO campaign to the AGM date through
-$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
+Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
+$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
 where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively.
-The dependent value rebases $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms).
-That is, we investigate potential consumer boycotts in the months following an NGO campaign.
+The dependent value normalizes $n$’s sales by its asset to account for variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms).
 Standard errors are clustered at the firm level.
-Appendix Figure B5 reports the estimated $\beta_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles).
-We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so two quarters after a campaign targeted on the AGM day, implying a loss of sales to total assets of 0.7pp, or about 5%.
-Similar sale losses as a result of negative ESG corporate events have been documented in the literature (Conway and Boxell, 2024; Saint-Jean, 2024).
+Appendix Figure B8 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles).
+We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so to campaigns targeted on the AGM date, implying a loss of sales to total assets of 0.7pp, or about 5%—a magnitude consistent with other papers (Conway and Boxell, 2024; Saint-Jean, 2024).
+4.2 Does AGM-Targeting Create Positive Change?
+If public shaming were an effective strategy for influencing a firm’s operations, targeting the AGM month should lead to the greatest improvements in a firm’s E&S score.
+However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before an AGM result in a subsequent increase in firms’ E&S scores, with a 2 to 3% rise over the following four quarters.
+In contrast, campaigns launched at other times, including during the AGM month, have an economically and statistically insignificant effect on firms’ scores.
+Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms.
+Therefore,
 Result 2.
-NGO campaigns on an AGM date impact stakeholders.
-(i) Shareholders are more likely to make related proposals in future years.
-(ii) Consumers are more likely to boycott the firm in the following months.
-5 Timing Campaigns to Influence AGMs
+NGO campaigns on an AGM date impact stakeholders but do not lead to improvements in E&S scores.
+(i) Shareholders are more likely to make related proposals the following year.
+(ii) The target’s sales drop in the months after a campaign.
+5 Timing Campaigns to Influence AGM Votes
 Early campaigns could benefit an NGO aiming to influence current AGM outcomes.
 Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes.
-For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting (Microsoft, 2023), suggesting that a campaign close to the AGM date may have less impact on the voting results.
+For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting, indicating that campaigning around AGM dates may have little impact on shareholders’ votes (Microsoft, 2023).
 We define a proposal as successful if either it is withdrawn by the shareholder before the AGM or it secures at least $50\%+1$ vote at the AGM.
@@
 $\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
-where $\text{Success}_{i,n,m,y}$ is a dummy indicating the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+l,t}$ is a dummy indicating the occurrence of an NGO campaign on the same topic as proposal $i$, $l$ months before the AGM.
-The reference period covers the months that occur more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$.
-We account for firm-by-year, and topic-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and topics within firm-year pairs.
+where “$\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and “$\text{Campaign}_{i,n,m+l,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $l$ months before the AGM.
+The reference period covers the months that are more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$.
+We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs.
 The standard errors are clustered by firm.
-Figure 3(a) plots the estimated $\hat{\beta}_{l}$.
+Figure 3: NGOs’ campaigns before the AGM help E&S shareholder proposals succeed
+(a) Proposals accepted or withdrawn
+(b) Vote share in favor of proposal
+Note: E&S denote environmental and social proposals, respectively.
+The figure plots the estimated coefficients from (4).
+The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b).
+Standard errors are clustered at the firm level.
+Figure 3(a) plots the estimated $\{\hat{\beta}_{l}\}_{l}$.
 We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals.
-They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more than seven months before.
+They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more six months earlier.
 To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable.
 Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”).
-Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (Shu, 2024).
-Those recommendations have a critical influence on the vote decisions.
-We thus include management and ISS’s recommendations as covariates in (4).
-Figure 3(b) shows that only NGO campaigns occurring a month or three months before the AGM increases support for related shareholder proposals.
-This increase is by nearly 2 percentage points.
-Given the generally low baseline support for these proposals (24%), this translates to a 8% increase in shareholder support when an NGO campaign addresses the same topic, on average.
-In contrast to Figure 3(a), there is no significant increase in support for campaigns initiated as early as six months before the AGM.
-This suggests that the difference between the two figures is due to shareholder proposals being withdrawn after reaching an agreement with the firm.
-Figure 3: NGOs’ influence on the vote of E&S shareholder proposals
-(a) Proposals accepted or withdrawn
-(b) Shareholder support for proposal
-Note: E&S denote environmental and social proposals.
-The figure plots the estimated coefficients of regressions of a dummy indicating that a shareholder proposal is withdrawn or passes if it goes to a vote (left panel, based on 911 proposals), or the percentage of votes in favor of the proposal when it goes to the vote (right panel, based on 2,078 proposals), on dummies indicating the occurrence of an NGO campaign on the same topic $k$ months before the AGM, for $k$ varying from -6 to 0.
-The regressions also include month, firm-by-year, and topic-by-year fixed effects.
-Standard errors are clustered at the firm level.
+Institutional investors often follow proxy advisor recommendations, particularly from ISS (Shu, 2024).
+We include management and ISS recommendations as covariates in (4).
+Figure 3(b) shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points.
+Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
+In contrast to Figure 3(a), Figure 3(b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM.
+This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm.
+As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
+Moreover, the fact that campaigns launched one to three months before an AGM help secure more votes aligns with the timing of proxy voting.
+Shareholders typically receive voting materials within 90 days of the AGM, a period when AGM visibility in the news may increase as well as the salience of the campaign for shareholders (ISS, 2022).
 Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015).
-For this reason, we investigate all E&S proposals in Figure 3(a).
+For this reason, we investigate all E&S proposals in Figure 3(a), which also include withdrawn proposals.
 Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote.
-However, those votes can still be helpful in disciplining firms (Levit and Malenko, 2011) as large shareholder support to a cause may be hard to conceal for managers.
+However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011).
 Result 3.
-NGO campaigns are more effective at influencing a firm’s decisions when launched shortly before the AGM.
+NGO campaigns are more effective at influencing a firm’s decisions when launched before the AGM.
 6 Exposure vs. Influence
-Therefore, Sections 4 and 5 outline a tradeoff: late campaigns impact society by boosting an NGO’s visibility among the media and key stakeholders, while early ones are more effective at influencing a firm’s decisions.
+Sections 4 and 5 outline a tradeoff: late campaigns expose a firm’s wrongdoing to its stakeholders and boost the NGO’s visibility, while early ones are more effective at directly influencing a firm’s decisions.
 How do NGOs navigate this tradeoff?
@@
 We then examine whether this “visibility capital” from previous campaigns increases the likelihood of success at the AGM for shareholder proposals related to the NGO campaign.
-6.1 Campaign Timing and Accumulated NGO Visibility
-Section 4 suggests that targeting a firm’s AGM increases an NGO’s visibility in the media and among stakeholders.
-This visibility could enhance an NGO’s reputation for commitment to a cause.
-In this section, we explore whether NGOs, after establishing such a reputation, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
+6.1 Accumulating Visibility
+Targeting AGM-dates could enahnce an NGO’s visibility and its reputation for commitment to a cause.
+In this section, we explore whether NGOs, after accumulating some visibility, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
 We investigate this mechanism by:
-$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
-which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{c,g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
+$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}$ (5)
+$+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$
+which examines “$\text{Distance to AGM}_{c(i,n,g),y}$,” the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using an indicator, “$\text{Campaign \#}_{c,g,y}$,” that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
 Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010.
-To control for this truncation as well as time varying confounders—e.g., the size of the NGO social media network across years (Tufekci, 2017)—we include NGO-by-year fixed effects.
-We also include firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign.
-Hence, the identification of $\beta_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^9] In Appendix Figure B6, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
+To control for this truncation as well as time varying confounders such as NGO size, we include NGO-by-year fixed effects.
+We also include firm-by-year and cause-by-year fixed effects to account for firm and cause-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign.
+Hence, the identification of $\{\beta_{\tau}\}_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^11] In Appendix Figure B10, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
 Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample.
 The standard errors are clustered at the NGO level.
-To interpret the estimated $\beta_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM.
+To interpret the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM.
 Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date.
 [^12] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year.
-We postpone the interpretation of this result to Section 7, which sums up and discuss the results.
+We postpone the interpretation of this result to Section 7, which sums up and discusses the results.
 Alternative mechanism—learning.
@@
 NGO do not increasingly launch campaigns against firms that may be more easily influenced over time.
-Appendix Figures B8 and B9 exclude selection on observables.
+Appendix Figures B12 and B13 exclude selection on observables.
 For instance, target sizes do not covary (market capitalization) with NGO experience.
@@
 NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of the votes.
-Figure 4: NGO experience and the timing of NGO campaigns
-Note: This figure plots the estimated coefficients from the regression of the distance between NGO campaigns and the targeted firm’s next AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
-Regressions are based on a sample of 1,136 NGO campaigns.
+Figure 4: More experienced NGOs campaign earlier
+Note: This figure plots the estimated coefficients from (5).
+The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date.
+Only firms with at least two campaigns are considered (1,136 observations).
 Standard errors are clustered at the NGO level.
@@
 We assess this channel with the following regression:
-$\text{Success}_{i,n,y}$ $=\beta_{1}\text{Campaign on Same Topic}_{c(i,n,g),y}+\beta_{2}\text{Experience}_{g,y}$ (6)
-$+\beta_{3}\text{Campaign on Same Topic}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
+$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g),y}+\beta_{2}\,\text{Experience}_{g,y}$ (6)
+$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
 $+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
-where $\text{Success}_{i,n,y}$ is defined as in (4)—namely, either a successful vote or a withdrawn proposal.
-On the right-hand side, $\text{Campaign on Same Topic}_{i,n,y}$ is one if there has been a campaign by NGO $g$ against firm $n$ on topic $i$ in the ten months preceding the AGM and zero otherwise, and $\text{Experience}_{g,y}$ indicates the number of campaigns NGO $g$ had launched in the past.
-We control for all confounding variations at the month, firm-by-year, NGO-by-year, and topic-by-year through fixed effects.
+where “$\text{Success}_{i,n,y}$” is defined as in (4)—namely, either a successful vote or a withdrawn proposal.
+On the right-hand side, “$\text{Campaign on Same Cause}_{i,n,y}$” is one if there has been a campaign by NGO $g$ against firm $n$ on cause $i$ in the ten months preceding the AGM and zero otherwise, and “$\text{Experience}_{g,y}$” indicates the number of campaigns NGO $g$ had launched in the past.
+We control for all confounding variations at the month, firm-by-year, NGO-by-year, and cause-by-year through fixed effects.
 Standard errors are clustered by firm.
@@
 Observations 38,613 38,613 38,613 38,613
-R2 0.32100 0.32065 0.32101 0.32260
+R2 0.321 0.321 0.321 0.323
 *–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
-  • Note: This table reports the estimated coefficients from the regression of a dummy indicating whether an E&S shareholder proposal is withdrawn before voting or passes after voting (Success) on a dummy indicating the occurrence of an NGO campaign on the same topic as the proposal in the 10 months preceding the AGM (Campaign on Same Topic), a dummy indicating whether the NGO has launched a campaign on a firm’s AGM day in the past (AGM Experience (0/1)), and their interaction.
+  • Note: This table reports the estimated coefficients from (6) in Column 1.
+In Column 1, Experience is capped at 5 to avoid outliers–as 85% of NGOs never reach that level of experience.
 In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic.
 In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic.
-All regressions also include Firm x Year, NGO x Year, and Topic × Year fixed effects.
-Clustered (firm) standard errors are reported between parentheses.
+All regressions also include firm-by-year, NGO-by-year, topic-by-year, and month fixed effects.
+Clustered (firm) standard errors are reported in parentheses.
 Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$.
-Campaigning on the same topic as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero.
-Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a topic being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
+Campaigning on the same cause as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero.
+Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a cause being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
 Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns.
-We replace $\text{Experience}_{g,y}$ for the variable $\text{AGM Experience}_{g,y}$, which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise.
+We replace “$\text{Experience}_{g,y}$” for the variable “$\text{AGM Experience}_{g,y}$,” which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise.
 The interaction coefficient is significant and economically important.
-While a campaign on the same proposal topic increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
+While a campaign on the same cause as the shareholder proposal increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
 […]
 Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates.
-Column (4) replaces $\text{Experience}_{g,y}$ for the cumulative count of the campaigns when NGO $g$ targeted an AGM date ($\text{AGM Experience}_{g,y}$) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and Campaign on Same Topic imply a concave relationship between AGM experience and AGM success.
-The results are consistent with Figure 4, which shows that only the very first campaigns happen close to an AGM date.
+Column (4) replaces “$\text{Experience}_{g,y}$” for the cumulative count of the campaigns when NGO $g$ targeted an AGM date (“$\text{AGM Experience}_{g,y}$”) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and “$\text{Campaign on Same Cause}_{i,n,g,y}$” imply a concave relationship between past AGM-targeting and proposals’ success.
+Therefore, the support of NGO’s to a shareholder proposal display diminishing returns.
+This finding supports the results in Appendix Figure B4, discussed in Section 4, which demonstrate that AGM-targeting increases the probability that target receives negative media coverage only if the NGO has little campaign experience (i.e., having run less than four campaigns).
+Later campaigns do not generate additional awareness on AGM dates compared to other dates.
+In fact, “mature NGOs” seem to be better off campaigning one or two months earlier, indicating diminishing returns from campaigning on AGM dates, even in terms of visibility.
+We conclude that,
 Result 5.
-Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but display decreasing returns.
-7 Discussion
-Our results demonstrate that NGOs strategically use the timing of their campaigns to influence firms’ decisions.
-In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaigns against target firms’ practices or strategies (Result 1).
-By doing so, NGOs prompt reactions from key stakeholders, such as shareholders, who are more likely to file a related proposal at the following AGM, and consumers, as the likelihood of boycotts increases if NGOs campaign on the AGM date (Result 2).
-The goal of these campaigns is to alter a firm’s strategy by damaging its reputation and showing stakeholders’ disappointment (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
+Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but displays decreasing returns.
+7 Discussion and Conclusion
+By analyzing the timing of NGO campaigns, we demonstrate that this seemingly minor decision has substantial consequences for both targeted firms and the effectiveness of future initiatives.
+Furthermore, we demonstrate that NGOs’ tactics evolve over time as they seek to either indirectly pressure firms (“exposure”) or directly shape their operations (“influence”).
+Thus, systematically analyzing NGOs’ behavior can shed new light on how activism helps firms internalize externalities.
+We first show that NGOs tend to launch their campaigns on the AGM date, as media coverage is higher during AGMs (Result 1).
+By doing so, NGOs elicit reactions from key stakeholders, such as shareholders and consumers (Result 2).
+The immediate goal of these campaigns is to influence a firm’s strategy by damaging its reputation and voicing stakeholder dissatisfaction.
 However, campaigning before the AGM offers direct advantages.
 Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3).
-Since shareholder proposals generally receive low support, early NGO campaigns can play a role in disciplining managers (Levit and Malenko, 2011).
-Importantly, these analyses control for NGO-by-year fixed effects and topic-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., WWF) do not solely drive these results.
-Also, we provide evidence suggesting that NGOs campaigns on AGMs do not target different firms compared to other campaign times in Appendix Table A2.
+Since shareholder proposals generally receive low support, early NGO campaigns can discipline managers (Levit and Malenko, 2011).
+Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results.
+Also, we provide evidence suggesting that NGO campaigns on AGMs do not target different firms compared to campaigns at other times, which helps mitigate selection issues (Appendix Table A2).
 Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time.
-NGOs with limited activism histories tend to campaign on the AGM date, while more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
-This pattern is consistent with the idea of using campaigning time as a signaling device.
-One interpretation of this result is that, as NGOs gain more experience with campaigning—and potentially run multiple campaigns in a year—they accumulate exposure through standard news media and may already have a large social media following.
-As a result, there may be less pressure to campaign specifically on the AGM date.
-Another interpretation is that, as NGOs become more complex organizations over time, media exposure becomes less vital for influencing decision-making at target firms.
-Disentangling these interpretations suggest several potential avenues for future research, contributing to the broader conversation on civil society lobbying and activism (Bertrand et al., 2021, 2023).
-The public may also expect mature NGOs not only to generate negative publicity but also to produce immediate results.
-Thus, there may be diminishing returns to reputation from consistently targeting AGMs.
-We show that this practice enables NGOs with a history of targeting AGMs to help related shareholder proposals pass.
-However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns on average, decreases support for shareholder proposals (Result 5).
-In conclusion, NGOs play a crucial role in increasing information and aligning corporate interests with societal needs, and their growth in recent years is well-documented (e.g., Egorov and Harstad, 2017; Daubanes and Rochet, 2019).
-However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited due to the scant data available on NGOs and models to interpret their actions.
-This paper offers a detailed analysis of NGO campaigns to uncover these trade-offs and incentives by leveraging the exogeneity of the corporate calendar to NGO’s goals.
-Similar approaches have been successfully used to identify the preferences of other stakeholders and may help reframe the role of AGMs not just as voting venues but also as forums where firms engage with and negotiate alongside outside stakeholders in ways that are not often observed by researchers but can be inferred from the underlying system of incentives (Fioretti et al., 2023).
+NGOs with limited activism histories tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
+One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
+Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5).
+First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device.
+This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following ”reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future.
+[^13] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
+Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics.
+Indeed, we find that mature NGOs—those with multiple prior campaigns in the sample—no longer inflict additional reputation losses by campaigning on the AGM date, whereas they did so in their first campaigns (compare the blue dots and red triangles in Appendix Figure B4).
+In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
+Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause.
+For instance, organizational aging (e.g., older NGOs becoming larger and more complex, with multiple ways to influence firms) is accounted for through fixed effects.
+These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
+In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied.
+This paper analyzes NGO campaigns, leveraging the exogeneity of the corporate calendar to uncover their trade-offs.
+Similar approaches have identified stakeholder preferences, showing that AGMs are more than voting venues—they also serve as forums for firm-stakeholder negotiations, often unseen by researchers but inferred from underlying incentives (Fioretti et al., 2023).
 We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010).
-Future research could capitalize on similar exogenous variation to that employed in this paper to quantitatively study the incentives for collaboration.
-By focusing on the timing of NGO campaigns, we demonstrate how this seemingly innocuous decision has significant implications for both targeted firms and the success of future initiatives.
-We show that, similar to for-profit organizations, NGOs are influenced by their reputation, which affects the timing of the information presented to stakeholders and, consequently, their responses.
-Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize firms’ externalities.
-8 Conclusion
-This paper examines the timing and motives behind NGO campaigns, revealing that these campaigns are not randomly timed but often strategically target firms’ Annual General Meetings.
-While this timing does not affect votes on environmental and social proposals, which are cast beforehand, it enhances the NGO’s visibility, leading to greater consumer reactions and increasing the likelihood of related shareholder proposals in the following year.
-Mature NGOs utilize the reputation similarly built over the years to influence current firm decisions by supporting relevant shareholder proposals.
-Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its life, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
+Future research could capitalize on similar exogenous variation to quantitatively study the incentives for collaboration.
 Appendix A Omitted Tables
@@
 Appendix B Omitted Figures
-Figure B1: E&S shareholder proposals, and their success
+Figure B1: Count of E&S shareholder proposals and their success
 Note: This figure plots the number of E&S proposals at Russell 3,000 firms from 2006 to 2021 (solid line) and the number of successful E&S proposals (blue area) in the same period.
@@
 The NGO data come from Sigwatch; the AGM dates are from ISS.
+Figure B4: Media exposure of NGO campaigns around an AGM
+(a) Main specification
+(b) Controlling also for current E&S news
+Note: The figure presents estimates from (1).
+The dependent variable is an indicator that a firm receives at least one negative E&S news in a week.
+Panel (b) also controls for the current count of E&S news to account for firm-week specific events that increase news coverage (not necessarily negative coverage).
+Standard errors are clustered at the firm level.
+Figure B5: Examples of positive and negative E&S news
+(a) Example of a positive E or S news
+(b) Example of a negative E or S news
+Notes: This figure displays examples of a positive (top) and negative (bottom) E&S news. News data from RavenPack.
+Figure B6: Media sentiment of NGO campaigns around an AGM
+Note:The figure presents estimates from (1). The dependent variable is the average sentiment of E&S news received by firm in a week. Standard errors are clustered at the firm level.
 Figure B7: NGO campaigns and shareholder proposals
-Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same topic, $k$ months ahead.
+Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead.
 The regression also includes month-by-year-by-industry-by-topic fixed effects.
@@
 The number of observations is 351,351 and 10,188, respectively.
+Standard errors are clustered at the firm level.
+Figure B9: Campaign Timing and E&S Score Improvement
+Note: This figure plots the estimated coefficients of the following regression:
+$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
+where $\Delta_{4}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter $q$ and quarter $q+4$, and $\sum_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM.
+The regression also controls for firm size (logged total assets), current average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$).
+ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
+The regression has 47,082 observations.
 Standard errors are clustered at the firm level.
@@
 Appendix C Matching Proposals And NGO Campaigns
-This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign topic.
+This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign cause.
 Sigwatch assigns a“Generic Topic” to each campaign in their dataset.
-To streamline the analysis, we combine certain topics when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category.
-Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” topic, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
-Using these consolidated “Generic Topic Names”, we then categorize each shareholder proposal based on its short description.
-When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for topic classification.
-We start by assigning proposals to topics based on keywords in their descriptions and then manually classify any remaining proposals.
-Table C1: Proposal Topics
+To streamline the analysis, we combine certain causes when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category.
+Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” cause, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
+Using these consolidated “Generic Cause Names”, we then categorize each shareholder proposal based on its short description.
+When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for cause classification.
+We start by assigning proposals to causes based on keywords in their descriptions and then manually classify any remaining proposals.
+Table C1: Proposal Causes
 NGO Campaign Category Example of Shareholder Proposals

### v4

@@
 1 Introduction
-Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017).
-However, their actions are often difficult to observe, and evidence on the incentives they face and how these shape their strategic decisions remains limited (Yaziji and Doh, 2009).
-Thus, NGOs are frequently perceived as passive actors, either selecting the most favorable action from a common good perspective (Daubanes and Rochet, 2019) or being dysfunctional and susceptible to corporate capture (Bertrand et al., 2020).
-This paper introduces a novel approach to analyzing how NGOs make decisions with implications for corporations and their stakeholders.
-Consider an NGO with information about corporate wrongdoing that faces a strategic choice: launch an immediate campaign to foster dialogue or wait for the optimal moment when the firm is most sensitive to external pressure.
-While early action encourages engagement and potential change, strategic timing can amplify impact by signaling strong opposition among stakeholders.
-A time when firms are particularly sensitive to external pressure is during their Annual General Meetings (AGMs), recurring shareholder meetings where strategic decisions are revealed and executives are held accountable.
-We show that NGOs are four to six times more likely to campaign on a target’s AGM date than at other times, a pattern driven by the heightened media coverage AGMs attract.
-This visibility helps NGOs reach stakeholders and the broader public while imposing reputational costs on firms.
-However, while AGMs provide a platform for NGOs to maximize exposure, their direct influence is limited—most shareholders vote before the meeting via proxy, reducing the immediate impact on voting outcomes (Bebchuk et al., 2020).
-Empirically, we leverage the exogenous timing of AGMs, which remains stable within firms due to SEC regulations, allowing us to focus on NGOs’ actions while holding firms’ strategic choices constant.
-Our findings reveal the “life cycle” of NGO strategies (Wernerfelt, 1988; Cabral, 2000).
-Early-stage NGOs prioritize visibility by campaigning on AGM dates, despite their limited direct influence.
-We show that AGM-timed campaigns prompt consumer backlash and increase the likelihood of related shareholder proposals the following year.
-However, while these campaigns generate exposure, they do not directly drive change, as, for instance, the target’s environmental and social scores remain unchanged.
-Over time, NGOs targeting AGMs start campaigning before AGMs to maximize direct influence on corporate decisions.
-Early campaigns increase the likelihood of related shareholder proposals succeeding, ultimately driving measurable improvements in firms’ environmental and social performance.
-Our interpretation is that AGM exposure can enhance NGO reputation.
-Supporting this view, we find that the success of a shareholder proposal on a given cause increases when an NGO campaigns on the same issue and year, with past AGM-date campaigns by the NGO amplifying this effect.
-However, the returns to AGM-date campaigns diminish as their frequency rises.
-In contrast, past NGO campaigns on non-AGM dates have no impact.
-We view AGM-date campaigns as a way for NGOs to signal their type to stakeholders and build their “reputation,” as evidenced by the diminishing returns of AGM-targeting over time, which is consistent with reputation models (Bar-Isaac and Tadelis, 2008).
-[^1] This practice also generates additional negative news to a target firm only for an NGO’s first few campaigns.
-Later campaigns inflict no additional damage.
-Thus, the NGO may be willing to incur a cost to build its “reputation-for-toughness,” which it can then leverage in future campaigns (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
-Hence, more mature NGOs shift to campaign before AGMs.
-Furthermore, NGO reputation is likely multidimensional, shaped both by a sustained track record of activism and by a demonstrated commitment to a cause.
-As such, NGOs may transition to early campaigning not only due to diminishing returns on visibility but also to establish credibility through constructive engagement with key stakeholders (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
+Over recent decades, non-governmental organizations (NGOs) have emerged as influential actors shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017).
+However, their actions are difficult to observe, and evidence on their strategic incentives remains limited (Yaziji and Doh, 2009).
+As a result, NGOs are often viewed either as principled actors promoting social good (Daubanes and Rochet, 2019) or as dysfunctional entities vulnerable to corporate capture (Bertrand et al., 2020).
+This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value.
+We focus on campaign timing—one of the few observable aspects of NGO activity, which remains far more opaque than that of for-profit firms.
+[^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base.
+Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored.
+While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior.
+We identify the main determinants of how NGOs resolve this trade-off and show that their choices shape which stakeholders are mobilized and how firms ultimately respond.
+Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000).
+Early-stage NGOs prioritize visibility by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals.
+However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly.
+Over time, NGOs shift to campaigning before a target’s AGM date to exert direct influence, increasing the success rate of current shareholder proposals and leading to measurable improvements in corporate behavior.
+Consistent with reputation-building models, shareholder proposals are more likely to succeed when accompanied by NGO campaigns on the same issue—especially when the NGO has a history of targeting AGM dates (Bar-Isaac and Tadelis, 2008).
+However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns.
+In contrast, past non-AGM campaigns have no influence on proposal outcomes.
+NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
+As a result, more mature NGOs shift toward earlier, pre-AGM campaigns, aiming to build credibility through constructive stakeholder engagement (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
+Our identification strategy leverages the timing of firms’ AGMs, which is exogenously fixed by SEC regulations and stable within firms over time.
+This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant.
+Our main endogenous variable, therefore, is NGO campaign timing, which we explain through NGO experience: more experienced NGOs campaign earlier, trading visibility for impact.
+While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience.
+This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
 By comparing NGO campaign outcomes across firms and issues within years, our analysis accounts for factors that vary across years within NGOs, firms, and issues, such as organizational competence, learning, moral authority, resources, and the salience of specific issues.
@@
 Before that, we detail our empirical approach.
-3 Exploiting AGM Dates
+3 AGM Dates
 Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations.
@@
 […]
-Figure B7 shows that only campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same cause.
-This substantial increase is as large as 12 percentage points, or 44%, above the average probability of observing a proposal.
-Thus, NGO campaigns from the previous AGMs likely lead to a shareholder proposal at the current AGM, potentially due to heightened media attention.
+Figure B7 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue.
+The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline.
+Since SEC 14a8(e)(2) rule prevents shareholders from submitting proposals within 120 days of the proxy statement’s release, the x-axis is capped at four months before the previous year’s AGM.
+This timing suggests that campaigns shortly after an AGM may spur proposals at the next one, possibly due to heightened media attention.
 Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome.
@@
 Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
-$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
+$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
 where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively.
-The dependent value normalizes $n$’s sales by its asset to account for variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms).
+The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample.
 Standard errors are clustered at the firm level.
-Appendix Figure B8 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles).
-We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so to campaigns targeted on the AGM date, implying a loss of sales to total assets of 0.7pp, or about 5%—a magnitude consistent with other papers (Conway and Boxell, 2024; Saint-Jean, 2024).
+Appendix Figure B8 shows the estimated ${\hat{\beta}\tau}\tau$ coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b).
+We find no consumer response to the latter, but sales drop by 0.7pp (about 5%) for the former—a magnitude consistent with Conway and Boxell (2024); Saint-Jean (2024).
+In Panel (c), the sales gap between AGM and non-AGM campaigns emerges one quarter before the event and continues to widen, possibly due to the greater media coverage of AGM dates.
 4.2 Does AGM-Targeting Create Positive Change?
-If public shaming were an effective strategy for influencing a firm’s operations, targeting the AGM month should lead to the greatest improvements in a firm’s E&S score.
-However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before an AGM result in a subsequent increase in firms’ E&S scores, with a 2 to 3% rise over the following four quarters.
-In contrast, campaigns launched at other times, including during the AGM month, have an economically and statistically insignificant effect on firms’ scores.
+If public shaming were an effective strategy to influence firm behavior, targeting the AGM month should yield the largest improvements in E&S scores.
+However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by 2–3% over the following year.
+By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
 Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms.
@@
 Therefore, the support of NGO’s to a shareholder proposal display diminishing returns.
+[^13] Note that the combined effect of AGM Experience and AGM Experience Squared never yields a negative probability of proposal success—even at the highest observed levels of AGM Experience—suggesting our reputation proxy does not exhibit negative returns.
 This finding supports the results in Appendix Figure B4, discussed in Section 4, which demonstrate that AGM-targeting increases the probability that target receives negative media coverage only if the NGO has little campaign experience (i.e., having run less than four campaigns).
@@
 Standard errors are clustered at the firm level.
-Figure B8: NGO campaigns and consumer reactions
-Note: This figure plots the estimated coefficients of the regression of sales over total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$.
+Figure B8: Examples of positive and negative E&S news
+(a) NGO Campaigns on AGM
+(b) NGO Campaigns on Other Days
+(c) Estimated Effect of Campaigning on AGM
+Notes: Panels (a) and (b) plot the estimated coefficients of the regression of sales over (lagged) total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$ from (3).
+Panel(c) plots the estimated $\hat{\beta}$ from
+$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign on AGM}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)$
+$+\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\tilde{\beta}_{\tau}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
+which interacts the time dummy, the occurrence of a campaign, and a dummy indicating that the campaign happened on AGM day.
 The regressions also include firm and quarter-by-industry fixed effects.
-The number of observations is 351,351 and 10,188, respectively.
+The reference period is -1 quarter, consistent with the other analyses.
+The number of observations is 10,188, 351,351, and 361,539, respectively.
 Standard errors are clustered at the firm level.
@@
 Note: This figure plots the estimated coefficients of the following regression:
-$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
+$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
 where $\Delta_{4}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter $q$ and quarter $q+4$, and $\sum_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM.
@@
 Standard errors are clustered at the firm level.
+Unlike an event study, equation (B1) is a distributed leads-and-lags regression.
+We omit a reference period, so the coefficients are interpreted relative to firms that did not experience a campaign.
+This choice allows us to directly compare the effects of early campaigns with those launched on the AGM date relative to firms without campaigns.
+Results remain similar if we instead set month –4 as the reference period, as in the other analyses.
 Figure B10: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects

### v5

@@
 This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value.
-We focus on campaign timing—one of the few observable aspects of NGO activity, which remains far more opaque than that of for-profit firms.
+We focus on campaign timing—one of the few observable aspects of NGO activity, which is far more opaque than that of for-profit firms.
 [^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base.
-Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored.
+Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored, especially outside developing-countries contexts (e.g., Deserranno et al., 2025).
 While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior.
@@
 Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000).
-Early-stage NGOs prioritize visibility by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals.
+Early-stage NGOs prioritize exposure by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals.
 However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly.
@@
 However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns.
-In contrast, past non-AGM campaigns have no influence on proposal outcomes.
+In contrast, past non-AGM campaigns have no influence on proposal success.
 NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
@@
 This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant.
-Our main endogenous variable, therefore, is NGO campaign timing, which we explain through NGO experience: more experienced NGOs campaign earlier, trading visibility for impact.
-While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience.
+Our main endogenous variable, therefore, is NGO campaign timing, which we explain by reference to each NGO’s prior campaign experience: more experienced NGOs campaign earlier, trading visibility for impact.
+While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience or campaign timing.
 This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
@@
 [^2] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ.
+Those single-firm campaigns represent 93% of the original sample.
 The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms.
@@
 Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote.
-[^3] Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023).
+[^3] Evidence shows that shareholders generally agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023).
 We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021.
@@
 Target selection.
-Our analysis focuses on campaign timing, abstracting from NGOs’ selection of a target, which we leave for future research.
-However, Appendix Table A1 shows that NGOs’ targets does not depend on prior campaign experience, a key variable of our analysis.
-Panel (c) finds no differences in market capitalization, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs.
-Panel (d) regresses firm characteristics on NGOs’ past campaigns, showing no correlation between target selection and NGO experience.
+While our analysis centers on campaign timing—leaving NGOs’ target selection to future research—Appendix Table A2 shows that NGOs campaigning on AGM dates do not select firms that differ in observable characteristics (market capitalization, assets, sales, media coverage, or environmental and social scores) from those targeted at other times, and campaign sentiment is likewise comparable across both groups.
+Moreover, Table A1 demonstrates that target choice is independent of prior campaign experience: Panel (c) reveals no significant differences in market cap, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs, and Panel (d) finds no correlation between firm characteristics and NGOs’ past campaigns.
 However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.
@@
 Result 4.
-NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of the votes.
+NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of shareholder proposals.
 Figure 4: More experienced NGOs campaign earlier
@@
 7 Discussion and Conclusion
-By analyzing the timing of NGO campaigns, we demonstrate that this seemingly minor decision has substantial consequences for both targeted firms and the effectiveness of future initiatives.
-Furthermore, we demonstrate that NGOs’ tactics evolve over time as they seek to either indirectly pressure firms (“exposure”) or directly shape their operations (“influence”).
-Thus, systematically analyzing NGOs’ behavior can shed new light on how activism helps firms internalize externalities.
+Examining NGO campaign timing reveals that not just the decision to intervene, but its timing, can materially affect firm behavior and shape the course of future activism.
+Over time, NGOs shift from visibility-driven exposure tactics to influence current strategies, highlighting how the temporal dimension of their strategy channels pressure that leads firms to internalize social and environmental externalities.
 We first show that NGOs tend to launch their campaigns on the AGM date, as media coverage is higher during AGMs (Result 1).
@@
 However, campaigning before the AGM offers direct advantages.
-Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3).
-Since shareholder proposals generally receive low support, early NGO campaigns can discipline managers (Levit and Malenko, 2011).
+Specifically, when an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of acceptance—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3).
+Given that proposals typically attract low support, early NGO interventions can effectively discipline managers (Levit and Malenko, 2011).
 Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results.
-Also, we provide evidence suggesting that NGO campaigns on AGMs do not target different firms compared to campaigns at other times, which helps mitigate selection issues (Appendix Table A2).
+Furthermore, we show that selection issues are not a central concern: Panel d of Appendix Table A1 shows target‐firm characteristics are invariant to an NGO’s past campaign experience, and Appendix Table A2 confirms NGOs campaigning at AGMs do not systematically target different firms than those campaigning at other times.
 Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time.
@@
 Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics.
-Indeed, we find that mature NGOs—those with multiple prior campaigns in the sample—no longer inflict additional reputation losses by campaigning on the AGM date, whereas they did so in their first campaigns (compare the blue dots and red triangles in Appendix Figure B4).
+Consistent with this, mature NGOs—those with multiple prior campaigns—do not inflict greater reputation harm by campaigning on the AGM date, whereas they do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B4).
 In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
 Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause.
-For instance, organizational aging (e.g., older NGOs becoming larger and more complex, with multiple ways to influence firms) is accounted for through fixed effects.
+For instance, organizational aging (e.g., older NGOs becoming larger and more complex) is accounted for through fixed effects.
 These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
@@
 Standard errors are clustered at the firm level.
-Figure B8: Examples of positive and negative E&S news
+Figure B8: NGO campaigns and consumer backlash
 (a) NGO Campaigns on AGM

### v6

@@
-secReferences in the Online Appendix
+NGO Activism:
 Abstract
-We analyze the timing of NGO campaigns to shed light on NGOs’ strategies and how they evolve over time.
-Data from 2,500 campaigns show that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date.
-Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer backlash and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes.
-As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
+We examine how NGOs time campaigns to influence corporate decisions at Annual General Meetings (AGMs).
+Using data on 2,500 U.S.
+campaigns, we find that NGOs frequently launch campaigns on the targeted firm’s AGM date, which boosts media exposure and triggers future shareholder proposals but does not sway current AGM votes or affect sales.
+As NGOs gain experience and reputational capital, they shift to earlier pre-AGM campaigns that increase shareholder support at the AGM and lead to improvements in corporate behavior.
+These patterns reveal a strategic trade-off between maximizing exposure and achieving direct influence.
 JEL classifications: L21, L31
-Keywords: Non-Governmental Organizations, campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
+Keywords: Non-Governmental Organizations, NGO campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
 1 Introduction
-Over recent decades, non-governmental organizations (NGOs) have emerged as influential actors shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017).
-However, their actions are difficult to observe, and evidence on their strategic incentives remains limited (Yaziji and Doh, 2009).
-As a result, NGOs are often viewed either as principled actors promoting social good (Daubanes and Rochet, 2019) or as dysfunctional entities vulnerable to corporate capture (Bertrand et al., 2020).
-This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value.
-We focus on campaign timing—one of the few observable aspects of NGO activity, which is far more opaque than that of for-profit firms.
-[^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base.
-Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored, especially outside developing-countries contexts (e.g., Deserranno et al., 2025).
-While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior.
-We identify the main determinants of how NGOs resolve this trade-off and show that their choices shape which stakeholders are mobilized and how firms ultimately respond.
-Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000).
-Early-stage NGOs prioritize exposure by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals.
-However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly.
-Over time, NGOs shift to campaigning before a target’s AGM date to exert direct influence, increasing the success rate of current shareholder proposals and leading to measurable improvements in corporate behavior.
-Consistent with reputation-building models, shareholder proposals are more likely to succeed when accompanied by NGO campaigns on the same issue—especially when the NGO has a history of targeting AGM dates (Bar-Isaac and Tadelis, 2008).
-However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns.
-In contrast, past non-AGM campaigns have no influence on proposal success.
-NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
-As a result, more mature NGOs shift toward earlier, pre-AGM campaigns, aiming to build credibility through constructive stakeholder engagement (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
-Our identification strategy leverages the timing of firms’ AGMs, which is exogenously fixed by SEC regulations and stable within firms over time.
-This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant.
-Our main endogenous variable, therefore, is NGO campaign timing, which we explain by reference to each NGO’s prior campaign experience: more experienced NGOs campaign earlier, trading visibility for impact.
-While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience or campaign timing.
-This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
-By comparing NGO campaign outcomes across firms and issues within years, our analysis accounts for factors that vary across years within NGOs, firms, and issues, such as organizational competence, learning, moral authority, resources, and the salience of specific issues.
-NGOs also engage firms through partnerships and other channels, but since such activities are voluntarily disclosed—and thus subject to selection biases (Burger and Owens, 2010)—we rather focus on campaign timing, which must be public to be effective and, thus, observable to researchers.
-Our findings open new avenues for research, as future work may empirically explore how NGOs choose between campaigning and other activist tactics or whether these different approaches act as complements or substitutes.
-The exogenous timing of AGMs provides a useful identification strategy for such analyses, as it holds the target firm’s strategic choices constant, allowing researchers to isolate NGOs’ choices and their incentives.
-Our results reveal a novel tradeoff between exposure and direct impact as a key driver of the societal consequences of NGO activism (Briscoe and Gupta, 2016).
-Unlike our framework, much of the existing literature assumes that activists primarily pursue social objectives, either stemming from prosocial behavior (Besley and Ghatak, 2005; Bénabou and Tirole, 2006) or moral preferences (Baron, 2010), without discussing their incentives.
-At the same time, several studies highlight instances in which NGOs themselves are subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021).
-Firms may also strategically leverage their relationships with NGOs—through donations or partnerships—to advance political agendas (Bertrand et al., 2020), or to restore consumer and investor trust following reputational damage (Akey et al., 2021).
-Our main contribution in this literature is to show that NGOs are not passive actors.
-Rather, their tactics evolve over time, shaping firms and their stakeholders’ responses in different ways depending on the strategies they employ (e.g., Hendricks and Sorensen, 2009; Cabral and Hortacsu, 2010).
-Thus, understanding NGOs’ incentives is essential to learn why firms internalize their externalities.
-We also contribute to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024).
-Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities.
-However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear (e.g., Hawn and Ioannou, 2016).
-The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010).
-However, other agents may play a pivotal role.
-Among them, we highlight that NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Koenig and Poncet, 2019; Fioretti et al., 2023; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023).
+Non-governmental organizations (NGOs) are increasingly prominent actors in shaping firms’ environmental and social policies (Baron, 2001; Egorov and Harstad, 2017; Yaziji and Doh, 2009).
+Yet, their strategic choices remain poorly understood.
+Much of the literature treats NGOs as either principled advocates motivated by social good (Besley and Ghatak, 2005; Bénabou and Tirole, 2006; Daubanes and Rochet, 2019) or as dysfunctional entities prone to corporate capture and strategic manipulation (Grossman and Helpman, 1994; Bertrand et al., 2020; Akey et al., 2021).
+We argue and demonstrate empirically that NGOs adapt their strategies in response to incentives, shaping which stakeholders they engage and ultimately affecting the social and environmental performance of targeted firms.
+To study how NGOs respond to incentives and their implications for firms, we focus on a setting where firms are especially exposed to outside pressure: the Annual General Meeting (AGM).
+AGMs are predictable focal points at which executives present results, shareholders vote on proposals, and media attention peaks (Fioretti et al., 2025).
+Figure 1 shows that NGO campaigns cluster heavily around AGM weeks: incidence is four to six times higher than in non-AGM weeks (Panel a), and media coverage spikes simultaneously (Panel b).
+These striking patterns raise a central question: do NGOs target AGMs mainly to maximize exposure, or do they also use campaigns to exert influence over corporate outcomes?
+Figure 1: Timing of NGO campaigns and media attention around AGMs
+(a) Campaigns peak during AGMs
+(b) Media coverage rises during AGMs
+Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,544 NGO campaigns.
+Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms.
+The NGO data come from Sigwatch; the news data are from Ravenpack.
+Examining campaign timing reveals that less experienced NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020).
+These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents.
+For instance, as the NGO Public Eye put it in its 2025 campaign timed on the AGM date of Inditex: “As shareholders, you have the power to put Inditex on a sustainable course” (Clean Clothes Campaign, 2025).
+Indeed, we find little evidence that other stakeholders are moved by these campaigns.
+For example, we find no evidence that sales are differently affected based on campaign timing relative to the AGM date, indicating that consumers may not be the main audience of NGOs.
+While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves.
+Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively inexperienced.
+These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times.
+This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
+As NGOs accumulate visibility through AGM-date campaigns, their strategic calculus changes.
+Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes.
+By contrast, campaigns launched in the months leading up to the AGM substantially raise the likelihood that related shareholder proposals succeed.
+Intervening one to three months before the meeting boosts average shareholder support by 8%, while even earlier campaigns often prompt firms to settle before proposals reach a vote, leading to measurable improvements in environmental and social performance.
+Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns forward in time as they gain experience and reputational capital.
+First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions.
+This pattern is especially pronounced among NGOs with stronger reputations, proxied by higher search visibility, which campaign several months ahead of the AGM rather than on the day itself.
+Repeated targeting of AGMs also plays a distinct role.
+By returning to these focal moments, NGOs build credibility and reputational standing, which in turn increases the effectiveness of later campaigns.
+Yet such returns are not indefinite: once credibility is established, additional AGM-day campaigns yield diminishing benefits.
+Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—shows how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for firms’ social and environmental impact.
+In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
+As NGOs accumulate reputational capital (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their strategy shifts: campaigns are launched months before the AGM, when they can still sway shareholder support and discipline managers.
+This dynamic lifecycle highlights NGOs’ role as coordinators that align stakeholder actions and channel them into constructive corporate change, as their strategies shift over time to balance visibility and influence (Briscoe and Gupta, 2016), with campaign tactics evolving dynamically across the life cycle to shape both corporate responses and stakeholder engagement (Cabral and Hortacsu, 2010).
+By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022).
+More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
+Prior studies highlight the role of shareholders influencing firms via voice through voting or exit via divestments (e.g., Cuñat et al., 2016; Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), and of consumers through willingness to pay (Gneezy et al., 2010).
+We instead show that NGOs act as coordination mechanisms, enabling diverse stakeholders to collectively discipline firms (Koenig and Poncet, 2019; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023).
+In doing so, NGOs expand the set of channels through which civil society shapes corporate conduct.
+Finally, our methodology relates to the literature on the strategic timing of actions to exploit fluctuations in media attention.
+Prior work shows that politicians and governments time executive orders, military operations, and even disaster relief to coincide with or avoid competing news (Eisensee and Strömberg, 2007; Durante and Zhuravskaya, 2018; Djourelova and Durante, 2022), and that social movements strategically leverage focal dates to amplify visibility (Madestam et al., 2013).
+We show that NGOs adopt similar tactics in the corporate domain.
 The paper is organized as follows.
-Section 2 outlines the data construction and provides summary statistics.
-The empirical approach is described in Section 3.
-Sections 4 and 5 examine media exposure and influence at the current AGM as potential drivers of AGM actions.
-Section 6 explores the role of NGO reputation in supporting related shareholder proposals.
+Section 2 outlines the data, and Section 3 provides background on AGM dates.
+Sections 4 and 5 examine NGO campaigns as tools to gain media exposure or to influence firms through key stakeholders, respectively.
+Section 6 brings these channels together by analyzing how NGO reputation enhances the success of shareholder proposals.
 Section 7 discusses the results and concludes.
 2 Data
-Our dataset integrates three sources.
+Our dataset integrates four sources: NGO campaigns, online attention, shareholder proposals, and firm-level news.
 First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020).
-The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative.
-We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline.
-[^2] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ.
-Those single-firm campaigns represent 93% of the original sample.
-The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms.
-The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace.
-The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil.
-The most common causes are “environment/pollution,” “human & labor rights,” “animal rights,” and “health.”
-Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms by market capitalization.
-For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
-We select environmental and social (E&S) proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total).
-Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote.
-[^3] Evidence shows that shareholders generally agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023).
-We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021.
-Appendix Figure B1 illustrates these trends.
-Lastly, we source firm-level news data from Ravenpack.
-RavenPack is a data analytics platform that transforms unstructured news content into structured data, offering sentiment analysis and event detection.
-The data allows us to count the news related to every firm headquartered in the US.
-[^4] We only retain news with a Relevance Score of at least 75%, which ensures the relevance of the news for the firm.
-We include only US-listed firms from 2010 to 2020.
-The sample includes 5,212 US firms.
-We link these three datasets using standard identifiers, including year, firms’ CUSIP, and ISIN numbers.
-Appendix Table A1 presents summary statistics for NGO-related variables (Panel a).
-On average, NGOs launch 3.40 campaigns during the sample period, but the distribution is skewed, with over 50% of NGOs initiating only one campaign.
-Most campaigns focus on social issues rather than environmental ones.
-Panel (b) displays the variation in key variables for the targeted firms, including market capitalization, assets, and news coverage.
-The following three sections examine an NGO’s payoffs in connection with the timing of its campaigns, and the reactions of central stakeholders.
-We present the key takeaways from each analysis at the end of each section.
-Before that, we detail our empirical approach.
-3 AGM Dates
+The data record the campaign’s public release date, cause (e.g., “GMOs in food”), the targeted company and its prominence, and sentiment (ranging from –2 to +2, typically negative).
+We focus on single-firm, negative-sentiment campaigns where the firm appears in the headline.
+[^1] Since our analysis links campaign timing with a firm’s AGM date, we exclude multi-firm campaigns as their AGM dates may differ.
+Single-firm campaigns represent 93% of the original sample.
+The final dataset spans 2,512 campaigns launched by 731 NGOs against 523 firms.
+Second, we collect monthly Google Trends scores for all NGOs in our sample from 2010 to 2022.
+Scores are normalized on a 0–100 scale, either (i) relative to the NGO’s own peak searches (within-NGO) or (ii) relative to a benchmark NGO in the same month (between-NGOs).
+We use Greenpeace—the most searched NGO—as the benchmark for comparability across organizations.
+Third, we use data from Institutional Shareholder Services (ISS) on proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms.
+We select 4,483 environmental and social (E&S) proposals filed between 2010 and 2021 (837 firms in total).
+About one-third of these proposals are successful, either withdrawn before the AGM (after firm commitments) or passed in a vote.
+Both the number of E&S proposals and their success rate increased steadily over the period.
+Lastly, we obtain firm-level news from RavenPack, which transforms unstructured news into structured data with sentiment and event tagging.
+We count news items related to US-listed firms from 2010 to 2020, retaining only items with a Relevance Score of at least 75%.
+The final sample includes 5,212 firms.
+We link Sigwatch, ISS, and RavenPack data using standard identifiers (CUSIP, ISIN, year).
+Appendix Table A1 reports descriptive statistics.
+Panel (a) shows that, on average, NGOs launch 3.2 campaigns during the sample, 24% of which are on the target AGM date, though over half of NGOs appear only once.
+Most campaigns focus on social rather than environmental issues, but the distribution is highly skewed: a small set of NGOs accounts for a disproportionate share of activity, while the median NGO runs only a single campaign.
+Panel (b) documents substantial variation across targeted firms in size and news coverage.
+Targets range from very large firms with extensive media presence to smaller firms with limited coverage, suggesting that NGOs are not exclusively focused on high-visibility corporations.
+53% of the campaigns are on social causes, with the remaining being environmental.
+We return to the remaining panels of Table A1 in Section 4, after first providing background on AGM dates.
+3 Institutional Background on AGM Dates
 Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations.
@@
 [^2] Results remain robust excluding the covid period (2020–2021).
-While external events may influence NGOs’ campaign decisions, they do not affect AGM timing.
-Thus, AGMs allow us to isolate NGOs’ strategic trade-off from firms’ responses.
-Target selection.
-While our analysis centers on campaign timing—leaving NGOs’ target selection to future research—Appendix Table A2 shows that NGOs campaigning on AGM dates do not select firms that differ in observable characteristics (market capitalization, assets, sales, media coverage, or environmental and social scores) from those targeted at other times, and campaign sentiment is likewise comparable across both groups.
-Moreover, Table A1 demonstrates that target choice is independent of prior campaign experience: Panel (c) reveals no significant differences in market cap, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs, and Panel (d) finds no correlation between firm characteristics and NGOs’ past campaigns.
-However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.
-4 Timing Campaigns for Media Exposure
-NGOs are substantially more likely to launch their campaigns around AGM dates, on average.
-According to Figure 1(a), NGO campaigns are about 4 to 6 times more common on AGM weeks compared to other weeks; while a typical week sees 28 campaigns, AGM weeks have more than 100.666Appendix Figure B3 confirms this finding showing a significant drop in campaign density immediately after AGM dates (p-value $\leq 1\%$) (Cattaneo et al., 2020).
-Figure 1: Timing of NGO campaigns and media attention around AGMs
-(a) Campaigns peak during AGMs
-(b) Media coverage rises during AGMs
-Note: Panel (a) displays the number of NGO campaigns launched relative to the date of the Annual General Meeting of the targeted firm (week 0), based on 1,489 NGO campaigns.
-Panel (b) compares the weekly average count of news during the AGM week (gray bars) and outside of the AGM week (white bars), based on weekly observations for 5,211 US firms.
-The NGO data come from Sigwatch; the news data are from Ravenpack.
-Calendar weeks are re-centered around the firm’s AGM.
-This finding could be related to the heightened media coverage during AGMs (Fioretti et al., 2023), as illustrated in Figure 1(b), where news coverage peaks in the 10 days around an AGM (gray bars) compared to other periods (white bars).
-To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
-$\text{News}_{n,w,y}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}{\{M_{n,w,y}=\tau\}}\,\cdot\,\text{NGO Campaign}_{n,w,y}\right)$ (1)
-$+\mathbf{X}_{n,w,y}\mathbf{\gamma}^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y}\,\,,$
-where “$\text{News}_{n,w,y}$” is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event.
-The independent variables include a set of dummies “$\mathbf{1}\{M_{n,w,y}=\tau\}$” indicating the distance in months from the closest AGM.
-The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise.777 The analysis aligns the calendar year with the corporate year, starting five months before the AGM month (month 0) and ending five months later, with months -5 and +5 slightly longer.
-Since SEC regulations make AGM months stable over time (Section 3), this relabeling does not alter the distribution of months by firms or double-count adjacent years.
-We interact these event-study indicators, with the variable “$\text{NGO Campaign}_{n,m,y}$,” an indicator variable equal to 1 if firm $n$ experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise.
-We set the reference month to four months before the AGM, following SEC rules that require the AGM date to be set at least four months in advance.
-Figure 2: Firms’ negative media coverage increases if NGOs campaign during AGMs
-Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Standard errors are clustered at the firm level.
-The vector “$\mathbf{X}_{n,w,y}$” includes “$\mathbf{1}{\{M_{n,w,y}=\tau\}}$” and “$\text{NGO Campaign}_{n,m,y}$” separately, and controls for (i) the occurrence of negative news and (ii) all E&S news over the previous four weeks to account for a firm’s recent media coverage.
-[^8] Adding controls for the current count of E&S news to account for NGOs targeting AGMs of high-coverage firms does not qualitatively change the results.
-Since (1)—as well as the following regressions—involves stacking multiple difference-in-differences regressions across years, we also include firm-by-year ($\alpha_{n,y}$) and industry-by-week-by-year fixed effects ($\rho_{w,y,s(n)}$) to control for yearly changes within firms as well as time-varying industry-relevant changes (e.g., policies or tariffs).
-Standard errors are clustered at the firm level.
-Figure 2 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients.
-Negative news follow a bell-shaped pattern around the AGM, peaking significantly above zero in the AGM month, highlighting the benefits of targeting AGMs.
-[^9] In section 6.2, we show that the spike on AGM date is particularly strong for an NGO’s first few campaigns but it fades away for later ones (Appendix Figure B4).
-To quantify the magnitude of the AGM peak, an NGO campaign launched in the AGM week doubles the probability of a negative E&S news compared to a campaign initiated four months earlier, on average.
-After the AGM, coverage returns to baseline levels, indicating that NGOs attract more media attention when campaigning closer to the AGM.
+While external events may shape NGOs’ incentives to campaign, they do not alter the timing of AGMs.
+This stability makes AGMs a useful benchmark: NGOs can choose to react immediately to a negative event concerning a firm, or alternatively wait and align their campaign with its AGM, when the firm is most exposed to scrutiny, as shown in Panel (b) of Figure 1, which shows that news coverage spikes in AGM weeks.
+Consistent with this view, we show in the Appendix that AGM-timed campaigns generate more negative news coverage than those launched at other times (Appendix Figure B2).
+Taken together, these patterns highlight why AGMs provide a natural focal point to study the strategic trade-off between maximizing visibility and exerting direct influence.
+4 Timing Campaigns for Visibility
+Having shown that NGO campaigns cluster around AGM dates and that they increase a target’s negative media coverage, we next ask whether such timing enhances NGOs’ visibility.
+We use Google search activity as a proxy for NGO visibility and estimate event-study regressions comparing campaigns launched on the AGM date with those launched at other times:
+$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$ (1)
+$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$
+$\quad+\alpha_{c,y}+\alpha_{m,y}+u_{g,m,y}.$
+Because NGOs typically campaign multiple times, we follow Baker et al.
+(2022a) and stack all campaign events.
+Each event is denoted $c(i,n,g,y,m)$, where campaign $c(\cdot)$ is identified by cause $i$ (e.g., pollution), NGO $g$ (e.g., Greenpeace), target firm $n$ (e.g., Shell), calendar year $y$, and month $m$.
+We then analyze variation in the Google Trends score for NGO $g$ within a five‐month window around the start of $g$’s campaign; this window may spill into years y1y-1 or $y+1$ if the campaign occurs early or late in year $y$.
+The main explanatory variables are event-time dummies, $\mathbf{1}_{\{M_{c,m,y}=\tau\}}$, equal to one if month $m$ is exactly $\tau$ months from campaign $c$, with $\tau\in[-5,5]$.
+The omitted category is the month immediately preceding the campaign ($\tau=-1$).
+These indicators are interacted with $\text{AGM Campaign}_{c}$, a dummy equal to one if campaign $c$ was launched on the target firm’s AGM date.
+Thus $\beta^{AGM}_{\tau}$ estimates the increase in visibility for NGO campaigning on an AGM date, over the direct effect of campaigning at any date, $\beta^{Non-AGM}$.
+The remaining terms include fixed effects at the NGO–firm–year level ($\alpha_{c,y}$), which absorb all variation across NGO-firm pairs over years, as well as month-by-year fixed effects ($\alpha_{m,y}$).
+Our main coefficients of interest ($\beta^{AGM}_{\tau}$) are therefore identified from within-pair variation across months.
+Standard errors are clustered at the NGO level.
+Under the hypothesis that AGM campaigns affect the visibility of experienced and inexperienced NGOs differently, we estimate (1) separately for an NGO’s first few campaigns and for later ones.
+Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and its subsequent campaigns (red triangles).
+The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months.
+For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established.
+These patterns suggest that inexperienced NGOs rely more heavily on AGM timing to build name recognition, whereas established NGOs gain less from repeating the same tactic.
+Figure 2: Increased NGO visibility around an AGM campaign
+(a) The first AGM campaigns bring visibility
+(b) No evidence of Selection
+Note: Panel (a) plots the estimated $\beta^{AGM}$ coefficients from (1).
+The dependent variable is the monthly Google Trends score for the NGO.
+Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns.
+Standard errors are clustered at the campaign level.
+Panel (b) reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects.
+Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
+We show that these results are robust to addressing three main concerns.
+First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months.
+This misalignment can distort estimates when treatment effects are heterogeneous across groups or over time.
+Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other.
+To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year.
+Both approaches yield similar results, with much larger gains for “unexperienced” NGOs.
+Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility—for example, by focusing on firms that are more vulnerable to pressure.
+Panel (b) of Figure 2 addresses this concern by regressing an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the cause of the campaign, with year fixed effects and firm-clustered standard errors.
+[^3] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), the firm’s Environmental and Social scores (Env.
+score, Soc.
+score), how negative the campaign is towards the firm (Campaign sentiment), and an indicator for the campaign cause being related to an environmental topic (Env.
+campaign).
+Note that no coefficient is significantly different from zero.
+We find that a one standard deviation increase in any of these variables moves the likelihood of an AGM campaign by less than 5% of a standard deviation, indicating little selection on observables.
+Further evidence on selection based on NGO experience comes from Table A1.
+Panel (c) compares firms targeted by NGOs with high versus low campaign experience and finds no significant differences across the same variables studied in Panel (b) of Figure 2.
+Panel (d) corroborates this by regressing firm characteristics on NGO experience (i.e., number of past campaigns), showing no correlation once fixed effects are included.
+The only systematic trend is that NGOs increasingly focus on environmental rather than social issues in later years, reflecting their rising importance.
+Thus, the evidence presented so far points to:
 Result 1.
-Media coverage of NGO campaigns is substantially greater during the target’s AGM.
-Importantly, we rule out the hypothesis that NGOs select more negative causes when targeting AGMs.
-Without an adequate ranking of causes by negativity, we estimate (1) with firm $n$’s average E&S news sentiment in week $w$ as the dependent variable.
-[^10] RavenPack scores the tone of a firm’s news coverage between -1 (adverse) and 1 (favorable).
-Examples of positive and negative news can be found in Appendix Figure B5.
-The estimated coefficients show that campaigns launched during AGM months are no different than other campaigns (Appendix Figure B6).
+Gains in visibility are greater for unexperienced NGOs campaigning on the target’s AGM day than for experienced NGOs.
+For clarity, each of the next sections highlights its main findings in a similar manner.
 4.1 Stakeholders’ Reactions
-This section explores how campaign timing affects critical stakeholders of a targeted firm, such as shareholders and consumers, while Section 4.2 examines its ramifications for the target’s operations.
-The relevant exhibits are in the Online Appendix.
+We next ask how campaign timing matters for stakeholders and firms’ social and environmental outcomes. We focus on shareholders and consumers, with supporting exhibits in the Online Appendix.
 4.1.1 Shareholders
-Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022).
+Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022b).
 Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest.
 We define as “related” a shareholder proposal on the same cause of the campaign in the previous year.
-Appendix C describes the matching procedure.
-For each campaign in our sample, we look at the probability that the next AGM has a related shareholder proposal, depending on the timing of the campaign:
-[…]
-$+\alpha_{i,s(n),g,y}+u_{i,n,y}$
-[…]
+Appendix D describes the matching procedure.
+For each campaign in our sample, we examine the probability that the next AGM includes a related shareholder proposal, depending on the campaign’s timing.
+To this end, we recalendarize time so that month 0 corresponds to the AGM month for each firm.
+As discussed in Section 3, AGM months are stable over time; thus, this recentering does not alter the distribution of months across years or risk double counting.
+[^4] All subsequent analyses use the same recalendarization.
+Months –5 and +5 are defined as 45-day intervals to ensure the recalendarized year covers the entire calendar.
+We estimate the following regression:
+$\text{Proposal on Same Cause}_{i,n,y}$ $=\sumop\displaylimits^{\tau=-4}_{\begin{subarray}{c}\tau=-15\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM${}_{c(i,n,g,y,m)}$}=-\tau\}}\right)$ (2)
+$+\alpha_{i,s(n),g,m,y}+\gamma_{n}+u_{i,n,y}$
+For each campaign $c(i,n,g,y,m)$ launched by NGO $g$ against firm $n$ on cause $i$ in year $y$ and month $m$, we analyze the likelihood that the next AGM of firm $n$ includes a shareholder proposal on the same cause.
+Formally, $\text{Proposal on Same Cause}_{i,n,y}$ equals one if such a proposal is observed within 4 to 15 months after campaign $c(\cdot)$, and zero otherwise, with timing measured by the indicator $\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=-\tau\}}$.
+Since SEC 14a8(e)(2) rule prevents shareholders from including proposals in the proxy statement within 120 days of its release, we cap the analysis at four months before the previous year’s AGM and set the reference month to $\tau=-4$ because of SEC rules.
+Although firm-by-time fixed effects are not possible in this setting as they would absorb the variation used to estimate $\beta_{\tau}$, this analysis purges all the variation from cause-by-industry-by-NGO-by-month-by-year through $\alpha_{i,s(n),g,m,y}$ and firm invariant characteristics through $\gamma_{n}$, which we view as the main confounding factor since issue salience varies systematically across firms, industries, causes, and NGOs over time.
+We cluster standard errors by firm.
 Figure B3 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue.
 The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline.
-Since SEC 14a8(e)(2) rule prevents shareholders from submitting proposals within 120 days of the proxy statement’s release, the x-axis is capped at four months before the previous year’s AGM.
-This timing suggests that campaigns shortly after an AGM may spur proposals at the next one, possibly due to heightened media attention.
+This timing suggests that campaigns around an AGM may spur proposals at the next one, possibly due to heightened media attention.
 Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome.
-If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately gain broad support on the shareholder’s proposal, without having to wait a year to advocate for change at the next AGM.
+If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partners to immediately gain broad support on the proposal, without having to wait a year to advocate for change at the next AGM.
 4.1.2 Consumers
 Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
-$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
-where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively.
+$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n,y}+\alpha_{s(n),q}+u_{n,q},$ (3)
+where $q$ indicates a quarter and $\alpha_{n,y}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm-by-year level or the industry-by-quarter-by-year level, respectively.
 The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample.
 Standard errors are clustered at the firm level.
-Appendix Figure B8 shows the estimated ${\hat{\beta}\tau}\tau$ coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b).
-We find no consumer response to the latter, but sales drop by 0.7pp (about 5%) for the former—a magnitude consistent with Conway and Boxell (2024); Saint-Jean (2024).
-In Panel (c), the sales gap between AGM and non-AGM campaigns emerges one quarter before the event and continues to widen, possibly due to the greater media coverage of AGM dates.
+Appendix Figure B4 shows the estimated $\hat{\beta}_{\tau}$} coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b).
+We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns.
+[^5] Unreported results show also no difference between experienced and unexperienced NGOs.
 4.2 Does AGM-Targeting Create Positive Change?
-If public shaming were an effective strategy to influence firm behavior, targeting the AGM month should yield the largest improvements in E&S scores.
-However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by 2–3% over the following year.
+If public shaming effectively influenced firm behavior, the AGM month should deliver the largest improvements in E&S scores, given that it triggers heightened negative news coverage of the target firm (Appendix Figure B2).
+However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year.
 By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
 Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms.
-Therefore,
+Therefore, the results in this section suggest that:
 Result 2.
-NGO campaigns on an AGM date impact stakeholders but do not lead to improvements in E&S scores.
-(i) Shareholders are more likely to make related proposals the following year.
-(ii) The target’s sales drop in the months after a campaign.
-5 Timing Campaigns to Influence AGM Votes
+NGO campaigns on AGM dates spur shareholders to place related proposals at future AGMs of the targeted firm, but they produce no contemporaneous improvements in the firm’s E&S scores.
+5 Timing Campaigns to Influence AGM Outcomes
 Early campaigns could benefit an NGO aiming to influence current AGM outcomes.
@@
 We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
-$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
-where “$\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and “$\text{Campaign}_{i,n,m+l,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $l$ months before the AGM.
-The reference period covers the months that are more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$.
+$\text{Success}_{i,n,m,y}=\sumop\displaylimits^{0}_{\begin{subarray}{c}\tau=-6\end{subarray}}\beta_{\tau}\cdot\text{Campaign}_{i,n,m+\tau,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
+where $\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+\tau,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $\tau$ months before the AGM.
+The reference period covers the months that are more than six months before the AGM of year yy—i.e., $m\in[-11,-7]$, as campaigns launched during that period are more likely to be in reaction to the previous year’s AGM results.
 We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs.
@@
 Figure 3: NGOs’ campaigns before the AGM help E&S shareholder proposals succeed
-(a) Proposals accepted or withdrawn
+(a) Proposal accepted or withdrawn
 (b) Vote share in favor of proposal
-Note: E&S denote environmental and social proposals, respectively.
-The figure plots the estimated coefficients from (4).
+Note: The figure plots the estimated coefficients from (4).
 The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b).
 Standard errors are clustered at the firm level.
-Figure 3(a) plots the estimated $\{\hat{\beta}_{l}\}_{l}$.
+Panel (a) Figure 3 plots the estimated ${\beta}_{\tau}$.
 We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals.
-They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more six months earlier.
-To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable.
-Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”).
-Institutional investors often follow proxy advisor recommendations, particularly from ISS (Shu, 2024).
-We include management and ISS recommendations as covariates in (4).
-Figure 3(b) shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points.
+They increase, respectively, the probability of success by 2 percentage points and 3 percentage points more than a campaign launched more than six months earlier, corresponding to a relative increase between 6 and 8%.
+To understand why campaigns launched six months before the AGM are also somewhat successful, we re-estimate (4) on the subset of proposals that went to a vote, using the percentage of votes in favor as the dependent variable.
+[^6] Since this analysis focuses on voted proposals, we also control for the voting recommendations of management and proxy advisors such as ISS (“For,” “Against,” or “Abstain”).
+Panel (a) of Figure 3 shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points.
 Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
-In contrast to Figure 3(a), Figure 3(b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM.
+In contrast to Panel (a), Panel (b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM.
 This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm.
 As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
-Moreover, the fact that campaigns launched one to three months before an AGM help secure more votes aligns with the timing of proxy voting.
-Shareholders typically receive voting materials within 90 days of the AGM, a period when AGM visibility in the news may increase as well as the salience of the campaign for shareholders (ISS, 2022).
 Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015).
-For this reason, we investigate all E&S proposals in Figure 3(a), which also include withdrawn proposals.
-Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote.
+Given the generally low support of shareholder proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote.
 However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011).
+This reinforces the importance of analyzing all E&S proposals, including those withdrawn as in Panel (a).
+The evidence in this section points to:
 Result 3.
-NGO campaigns are more effective at influencing a firm’s decisions when launched before the AGM.
+Pre-AGM campaigns raise shareholder support and proposal success, while AGM-day campaigns have little effect.
 6 Exposure vs. Influence
-Sections 4 and 5 outline a tradeoff: late campaigns expose a firm’s wrongdoing to its stakeholders and boost the NGO’s visibility, while early ones are more effective at directly influencing a firm’s decisions.
-How do NGOs navigate this tradeoff?
-In this section, we first explore whether NGOs, after gaining visibility by targeting AGMs in the past, switch to campaigning before the AGM.
-We then examine whether this “visibility capital” from previous campaigns increases the likelihood of success at the AGM for shareholder proposals related to the NGO campaign.
-6.1 Accumulating Visibility
-Targeting AGM-dates could enahnce an NGO’s visibility and its reputation for commitment to a cause.
-In this section, we explore whether NGOs, after accumulating some visibility, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
-We investigate this mechanism by:
-$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}$ (5)
-$+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$
-which examines “$\text{Distance to AGM}_{c(i,n,g),y}$,” the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using an indicator, “$\text{Campaign \#}_{c,g,y}$,” that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
-Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010.
-To control for this truncation as well as time varying confounders such as NGO size, we include NGO-by-year fixed effects.
-We also include firm-by-year and cause-by-year fixed effects to account for firm and cause-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign.
-Hence, the identification of $\{\beta_{\tau}\}_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^11] In Appendix Figure B10, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
-Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample.
-The standard errors are clustered at the NGO level.
-To interpret the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM.
-Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date.
-[^12] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year.
-We postpone the interpretation of this result to Section 7, which sums up and discusses the results.
-Alternative mechanism—learning.
-This result is unlikely due to NGOs learning that campaigning on the AGM date does not influence voting decisions, as most shareholders have already voted by proxy.
-Appendix Figure B11 shows the same downward trend observed in Figure 4 when subsetting to NGOs whose first campaign took place within 2 weeks after the AGM, when the NGO surely knew that it could not affect votes.
-Alternative mechanism—selection.
-NGO do not increasingly launch campaigns against firms that may be more easily influenced over time.
-Appendix Figures B12 and B13 exclude selection on observables.
-For instance, target sizes do not covary (market capitalization) with NGO experience.
-In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
+Sections 4 and 5 outlined a tradeoff: campaigns launched close to an AGM maximize exposure, while those launched earlier maximize influence.
+In this section, we examine how NGOs navigate this tradeoff as their visibility increases.
+We first study whether greater experience or reputation leads NGOs to shift campaign timing from the AGM date to earlier months.
+We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
+6.1 NGO Experience and Campaign Timing
+We first examine whether NGOs, as they gain experience, shift from targeting AGMs to launching earlier campaigns. Specifically, we estimate
+$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}$ (5)
+$\quad+\alpha_{t(\cdot),y}+\alpha_{g,y}+\alpha_{n,y}+u_{c(i,n,g,y,m)}.$
+where the dependent variable measures the number of months between campaign $c(\cdot)$ and the targeted firm’s AGM, and the key explanatory variable is the sequence number of campaigns by NGO $g$ by year $y$.
+$\alpha_{t(\cdot),y}$, $\alpha_{g,y}$, and $\alpha_{n,y}$ are topic-by-year, NGO-by-year, and firm-by-year fixed effects, respectively.
+Identification thus comes from NGOs launching multiple campaigns in the same year.
+[^7] Appendix Figure B6 shows similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
+Panel (a) of Figure 4 shows that first campaigns tend to occur right at the AGM, while subsequent ones are launched progressively earlier.
+This pattern is consistent with NGOs initially using AGMs to gain visibility and later exploiting their visibility to influence voting outcomes.
+These patterns are not driven by alternative explanations.
+First, they are unlikely to reflect NGOs learning that campaigning on the AGM date does not affect voting, since most shareholders have already voted by proxy.
+Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7).
+Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO experience (Appendix Figures B8 and B9).
+6.2 Reputation and Campaign Timing
+We next use Google Trends scores to proxy for an NGO’s reputation.
+We compute the average search score for NGO $g$ in year $y$ relative to Greenpeace in year y1y-1, discretized into deciles and regress:
+$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Relative Google Score}_{g,y-1}=\tau\!\right\}}\right)$ (6)
+$\quad+\alpha_{t(\cdot)}+\alpha_{y}+u_{c(i,n,g,y,m)},$
+which is similar to model (6) with $\text{Relative Google Score}_{g,y-1}$ as a measure of NGO visibility.
+$\alpha_{t(\cdot)}$ and $\alpha_{y}$ are topic and year fixed effects.
+Panel (b) of Figure 4 shows that only the most visible NGOs—those in the top deciles of relative searches—campaign several months ahead of AGMs, while less visible NGOs continue to concentrate on AGM dates.
+Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation or experience grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
 Result 4.
-NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of shareholder proposals.
-Figure 4: More experienced NGOs campaign earlier
+NGOs initially target AGM dates to gain visibility. As their experience or reputation increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
+Figure 4: NGO reputation and campaign timing
+(a) More experienced NGOs campaign earlier
+(b) More visible NGOs campaign earlier
 Note: This figure plots the estimated coefficients from (5).
 The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date.
-Only firms with at least two campaigns are considered (1,136 observations).
 Standard errors are clustered at the NGO level.
-6.2 NGO Success and Reputation
-We investigate whether NGO reputation, as proxied by past campaigns on AGM dates, makes NGOs more effective at producing positive corporate change.
-We assess this channel with the following regression:
-$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g),y}+\beta_{2}\,\text{Experience}_{g,y}$ (6)
-$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
+6.3 NGO Reputation and AGM Influence
+We now ask whether NGO experience and reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
+$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{Experience}_{g,y-1}$ (7)
+$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{Experience}_{g,y-1}$
 $+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
-where “$\text{Success}_{i,n,y}$” is defined as in (4)—namely, either a successful vote or a withdrawn proposal.
-On the right-hand side, “$\text{Campaign on Same Cause}_{i,n,y}$” is one if there has been a campaign by NGO $g$ against firm $n$ on cause $i$ in the ten months preceding the AGM and zero otherwise, and “$\text{Experience}_{g,y}$” indicates the number of campaigns NGO $g$ had launched in the past.
-We control for all confounding variations at the month, firm-by-year, NGO-by-year, and cause-by-year through fixed effects.
-Standard errors are clustered by firm.
+where $\text{Success}_{i,n,y}$” indicates whether a proposal is withdrawn or passes a vote, as in equation (4).
+The key variable $\text{Campaign on Same Cause}_{i,n,y}$” equals one if NGO $g$ campaigned against firm $n$ on cause $i$ in the ten months preceding the AGM.
+$\text{Experience}_{g,y-1}$” counts NGO $g$’s past campaigns.
+We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
 Table 1: NGO experience at targeting AGMs and success at supporting current shareholder proposals
@@
 Model: (1) (2) (3) (4)
-Main Variables
-Campaign on Same Topic 0.008 0.076 0.032 0.070
+Campaign on Same Topic -0.0003 -0.013 -0.005 -0.001
 (0.026) (0.028) (0.025) (0.030)
-Experience 0.018
-Campaign on Same Topic $\times$ Experience 0.026
-AGM Experience (0/1) -0.042 -0.041
+Experience -0.034
+Campaign on Same Topic $\times$ Experience 0.078∗∗∗
+Experience square 0.008∗∗
+Campaign on Same Topic $\times$ Experience square -0.015∗∗
+Non-AGM Experience 0.0008
+Campaign on Same Topic $\times$ Non-AGM Experience -0.011
+AGM Experience 0.005 -0.034
 (0.020) (0.023)
-Campaign on Same Topic $\times$ AGM Experience (0/1) 0.169∗∗ 0.150∗∗
+Campaign on Same Topic $\times$ AGM Experience 0.048 0.091∗∗∗
 (0.026) (0.032)
-Non-AGM Experience (0/1) 0.005
-Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.065
-AGM Experience -0.018
-Squared AGM Experience 0.002∗∗
-Campaign on Same Topic $\times$ AGM Experience 0.074∗∗
-Campaign on Same Topic $\times$ Squared AGM Experience -0.005∗∗∗
+AGM Experience square 0.010∗∗
+Campaign on Same Topic $\times$ AGM Experience square -0.013∗∗
+Campaign on Same Topic $\times$ Google Search Popularity 0.071
+Campaign on Same Topic $\times$ Google Search Popularity square -0.021
 Fixed-effects
-Firm x Year
-Topic x Year
+CUSIP x Year
 NGO x Year
-Month
-Fit statistics
-Observations 38,613 38,613 38,613 38,613
-R2 0.321 0.321 0.321 0.323
-*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
+Topic x Month-Year
+Observations 36,543 35,778 36,543 33,142
+R2 0.61894 0.61698 0.61995 0.62078
+*–p<0.1<0.1; **–p<0.05<0.05; ***–p<0.01<0.01.
   • Note: This table reports the estimated coefficients from (7) in Column 1.
-In Column 1, Experience is capped at 5 to avoid outliers–as 85% of NGOs never reach that level of experience.
-In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic.
-In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic.
-All regressions also include firm-by-year, NGO-by-year, topic-by-year, and month fixed effects.
+In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (Non-AGM Experience) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), and their interaction with Campaign on Same Topic.
+In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square, and their interaction with Campaign on Same Topic.
+In Column 4, we replace Experience from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign.
+All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects.
 Clustered (firm) standard errors are reported in parentheses.
-Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$.
-Campaigning on the same cause as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero.
-Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a cause being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
-Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns.
-We replace “$\text{Experience}_{g,y}$” for the variable “$\text{AGM Experience}_{g,y}$,” which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise.
-The interaction coefficient is significant and economically important.
-While a campaign on the same cause as the shareholder proposal increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
-[…]
-Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates.
-Column (4) replaces “$\text{Experience}_{g,y}$” for the cumulative count of the campaigns when NGO $g$ targeted an AGM date (“$\text{AGM Experience}_{g,y}$”) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and “$\text{Campaign on Same Cause}_{i,n,g,y}$” imply a concave relationship between past AGM-targeting and proposals’ success.
-Therefore, the support of NGO’s to a shareholder proposal display diminishing returns.
-[^13] Note that the combined effect of AGM Experience and AGM Experience Squared never yields a negative probability of proposal success—even at the highest observed levels of AGM Experience—suggesting our reputation proxy does not exhibit negative returns.
-This finding supports the results in Appendix Figure B4, discussed in Section 4, which demonstrate that AGM-targeting increases the probability that target receives negative media coverage only if the NGO has little campaign experience (i.e., having run less than four campaigns).
-Later campaigns do not generate additional awareness on AGM dates compared to other dates.
-In fact, “mature NGOs” seem to be better off campaigning one or two months earlier, indicating diminishing returns from campaigning on AGM dates, even in terms of visibility.
-We conclude that,
+Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$).
+What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has accumulated experience ($\beta_{3}>0$).
+A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average.
+Column (2) distinguishes between campaigns held on AGM dates and those held at other times.
+Only the former contribute to proposal success.
+In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
+Column (3) tests for diminishing returns by including squared terms.
+The results confirm a concave pattern: additional AGM campaigns continue to improve proposal success, but at a decreasing rate.
+Importantly, the combined effect remains non-negative even for the most active NGOs.
+[^8] The marginal effect of AGM Experience and its square never turns negative, so our proxy for reputation does not imply adverse returns.
+We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends.
+Specifically, Column (4) replaces experience by Google Search Popularityg,y, which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.999The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects.
+The interaction between this measure and campaigning on the same cause is positive, with a concave shape.
+Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
+These results also resonate with the media evidence in Section 4: AGM targeting increases negative media coverage only for relatively inexperienced NGOs.
+More seasoned NGOs gain little additional visibility from campaigning exactly on AGM dates and may be more effective when campaigning slightly earlier.
+These results support our view that:
 Result 5.
-Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but displays decreasing returns.
+Targeting AGMs builds NGO reputation and increases their ability to raise the supported of related shareholder proposals, but with diminishing returns.
 7 Discussion and Conclusion
@@
 Over time, NGOs shift from visibility-driven exposure tactics to influence current strategies, highlighting how the temporal dimension of their strategy channels pressure that leads firms to internalize social and environmental externalities.
-We first show that NGOs tend to launch their campaigns on the AGM date, as media coverage is higher during AGMs (Result 1).
-By doing so, NGOs elicit reactions from key stakeholders, such as shareholders and consumers (Result 2).
-The immediate goal of these campaigns is to influence a firm’s strategy by damaging its reputation and voicing stakeholder dissatisfaction.
-However, campaigning before the AGM offers direct advantages.
-Specifically, when an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of acceptance—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3).
-Given that proposals typically attract low support, early NGO interventions can effectively discipline managers (Levit and Malenko, 2011).
-Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results.
-Furthermore, we show that selection issues are not a central concern: Panel d of Appendix Table A1 shows target‐firm characteristics are invariant to an NGO’s past campaign experience, and Appendix Table A2 confirms NGOs campaigning at AGMs do not systematically target different firms than those campaigning at other times.
-Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time.
-NGOs with limited activism histories tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
+We first show that NGOs tend to launch their campaigns on the AGM date, likely motivated by the higher media coverage surrounding these events (Result 1).
+By doing so, NGOs elicit reactions from key stakeholders, such as shareholders.
+However, the immediate purpose of these campaigns is primarily to gain visibility by voicing stakeholder concerns; they have no direct effect on firms’ current E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
+Hence, campaigning before the AGM offers direct advantages.
+When an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of success—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3).
+Given the typically low baseline support for such proposals, early NGO interventions can effectively discipline managers who must contend with heightened shareholder pressure (Levit and Malenko, 2011).
+Importantly, all analyses include NGO-by-year and cause-by-year fixed effects, thereby accounting for time-varying confounders at both the NGO level—such as experience or reputation—and the cause level—such as specific news events.
+This ensures that globally renowned NGOs (e.g., PETA) do not drive the results.
+Furthermore, the analyses in Sections 4 and 6 show little evidence of selection on observables, alleviating concerns that NGOs disproportionately target firms that are easier to influence or more likely to generate visibility.
+Our results suggest that NGOs face a trade-off, which they navigate by adapting their strategies over time.
+NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
 One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
@@
 First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device.
-This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following ”reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future.
+This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following “reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future.
 [^10] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
-Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics.
-Consistent with this, mature NGOs—those with multiple prior campaigns—do not inflict greater reputation harm by campaigning on the AGM date, whereas they do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B4).
-In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
+Second, stakeholders may come to expect more constructive approaches from established NGOs, such as supporting related shareholder proposals.
+Consistent with this hypothesis, AGM-date campaigns by mature NGOs—those with multiple prior campaigns—do not generate more negative media coverage than campaigns at other times.
+By contrast, the very same NGOs do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B10).
+Hence, we observe not only a shift in strategy as NGOs mature (Figure 4) but also a change in their capacity to inflict reputational harm on firms.
 Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause.
@@
 In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied.
-This paper analyzes NGO campaigns, leveraging the exogeneity of the corporate calendar to uncover their trade-offs.
-Similar approaches have identified stakeholder preferences, showing that AGMs are more than voting venues—they also serve as forums for firm-stakeholder negotiations, often unseen by researchers but inferred from underlying incentives (Fioretti et al., 2023).
+We leverage the exogeneity of corporate events to uncover a trade-off between visibility and influence that NGOs face when deciding when to campaign, and show that how this trade-off is resolved determines which stakeholders are mobilized (e.g., shareholders today versus shareholders tomorrow).
+Hence, NGOs act as a coordination device, with distinct implications for the social and environmental impact of targeted firms.
 We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010).
 Future research could capitalize on similar exogenous variation to quantitatively study the incentives for collaboration.
+[…]
 Appendix A Omitted Tables
@@
 Number of campaigns by NGO 3.20 1 1 3 13
-      – E campaigns 1.51 0 1 1 5
-      – S campaigns 1.70 0 1 1 6
+     – Campaigns on Env. topic 1.51 0 1 1 5
+     – Campaigns on Soc. topic 1.70 0 1 1 6
      – On AGM date 0.24 0 0 0 1
@@
 Yearly Number of News 1,183 328 616 1,046 1,695.25
-E score 5.66 4 5.58 7.3 2.22
-S score 4.44 3.30 4.70 5.60 1.58
-E campaign 0.47 0 0 1 0.50
-S campaign 0.53 0 1 1 0.50
+Environmental score 5.66 4 5.58 7.3 2.22
+Social score 4.44 3.30 4.70 5.60 1.58
+Campaign on Env. topic 0.47 0 0 1 0.50
+Campaign on Soc. topic 0.53 0 1 1 0.50
 Panel c: Variables varying by target firm and NGO experience
-High Experience Low Experience Test difference
+NGO w/ AGM Experience NGO w/o AGM Experience Test difference
 Mean S.E. Mean S.E. p-value
@@
 Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
-E score 5.64 0.12 5.72 0.06 0.58
-S score 4.44 0.09 4.45 0.04 0.89
-E campaign 0.46 0.03 0.5 0.01 0.15
-S campaign 0.54 0.03 0.49 0.01 0.12
-Panel d: OLS Regression of $y_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
+Environmental score 5.64 0.12 5.72 0.06 0.58
+Social score 4.44 0.09 4.45 0.04 0.89
+Campaign on Env. topic 0.46 0.03 0.5 0.01 0.15
+Campaign on Soc. topic 0.54 0.03 0.49 0.01 0.12
+Panel d: OLS Regression of $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
 $\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
-Dependent Variable ($y_{ny}$): (1) (2) (3) (4)
+Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
 Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
@@
 Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
-E score 0.004 (0.009) 0.041 (0.041) 1,338
-S score -0.0006 (0.006) 0.019 (0.032) 1,338
-E campaign 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
-S campaign -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
-  • Note: This table reports summary statistics on NGO campaigns (Panel a, NGO campaign data from Sigwatch and AGM dates from ISS), firms targeted by the campaigns (Panel b, accounting data from Compustat, ESG scores from MSCI, and news count based on RavenPack data), as well as an analysis of the relationship between NGO experience and targeted firms’ characteristics (Panels c and d).
-In Panel c, the Market Cap, Total Assets, and Sales deciles are based on the universe of US firms in Compustat on the quarter of the campaign.
-We use deciles instead of dollar values because highly ranked firms saw a large expansion in market cap and assets over the last years.
-Ranking measures like deciles avoid such a mechanical correlation with NGO Experience, which, being defined as the number of campaigns launched by NGO g in previous years, also grows over time.
-In Panel d, the fixed-effects are: NGO×$\times$×Year, Firm×$\times$×Year and Topic×$\times$×Year, except for the last two rows where Topic×$\times$×Year fixed effects are not included because of perfect collinearity with the dependent variables.
-Standard errors are clustered by firm.
-Table A2: Campaign timing and characteristics of target firms
-Characteristics of Target Firm
-Market Cap Assets Sales Yearly Environmental Social Campaign
-Dependent Variables: (decile) (decile) (decile) News Score Score Sentiment
-Model: (1) (2) (3) (4) (5) (6) (7)
-Variables
-Campaign on AGM day (0/1) 0.469 0.495 0.374 83.2 0.469 -0.522 0.022
-(0.628) (0.695) (0.550) (185.7) (0.919) (0.715) (0.383)
-Fixed-effects
-Industry-Year
-NGO-Year
-Topic-Year
-Fit statistics
-Observations 1,338 1,338 1,337 1,299 1,338 1,338 1,338
-R2 0.93364 0.92669 0.90388 0.99232 0.95420 0.92167 0.84525
-*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
-  • Note: This table reports the estimated coefficients from regressions of different target firm characteristics on a dummy indicating whether the campaign took place on the target firm’s AGM day.
-All regressions also include Industry-by-Year, NGO-by-Year, and Topic-by-Year fixed effects.
-The Market Cap, Assets, and Sales deciles are based on the entire universe of US firms in Compustat in the quarter of the campaign.
-Yearly News refers to the four-quarter moving average of news mentioning the target firm.
-Accounting data from Compustat.
-NGO campaign (including Sentiment) data from Sigwatch.
-News data from RavenPack.
-Environmental and Social scores are from MSCI.
-Standard errors are clustered by firm.
+Environmental score 0.004 (0.009) 0.041 (0.041) 1,338
+Social score -0.0006 (0.006) 0.019 (0.032) 1,338
+Campaign on Env. topic 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
+Campaign on Soc. topic -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
+[…]
 Appendix B Omitted Figures
-Figure B1: Count of E&S shareholder proposals and their success
-Note: This figure plots the number of E&S proposals at Russell 3,000 firms from 2006 to 2021 (solid line) and the number of successful E&S proposals (blue area) in the same period.
-A successful proposal is defined as a proposal that is either withdrawn before the AGM, or that goes to vote and passes.
-Data from ISS.
-Based on 5,937 shareholder proposals filed at 979 firms.
 Figure B1: Persistence of the AGM month over time
@@
 Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
-Figure B3: Testing NGO campaign timing around the AGM date
-Note: This figure plots the density test for the occurrence of NGO campaigns around firms’ AGM, following Cattaneo et al.
-(2020).
-The test is based on 521 NGO campaigns launched within 50 days of the firm’s closest AGM.
-The NGO data come from Sigwatch; the AGM dates are from ISS.
-Figure B4: Media exposure of NGO campaigns around an AGM
-(a) Main specification
-(b) Controlling also for current E&S news
-Note: The figure presents estimates from (1).
-The dependent variable is an indicator that a firm receives at least one negative E&S news in a week.
-Panel (b) also controls for the current count of E&S news to account for firm-week specific events that increase news coverage (not necessarily negative coverage).
-Standard errors are clustered at the firm level.
-Figure B5: Examples of positive and negative E&S news
-(a) Example of a positive E or S news
-(b) Example of a negative E or S news
-Notes: This figure displays examples of a positive (top) and negative (bottom) E&S news. News data from RavenPack.
-Figure B6: Media sentiment of NGO campaigns around an AGM
-Note:The figure presents estimates from (1). The dependent variable is the average sentiment of E&S news received by firm in a week. Standard errors are clustered at the firm level.
+Figure B2: Firms’ negative media coverage increases if NGOs campaign during AGMs
+Note: The figure plots the coefficients $\beta_{\tau}$ from the following event study regression where we re-calendarize time around an AGM date so that the AGM month is month 0 in each year for each firm:
+$News_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\!\left(\beta_{\tau}\cdot 1\!\left\{M_{n,w,y}=\tau\!\right\}\cdot NGO\,Campaign_{n,w,y}\!\right)$
+$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
+where $News_{n,w,y}$ is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event.
+The set of dummies $1\{M_{n,w,y}=\tau\}$ indicate the distance in months from the closest AGM.
+The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise.
+We interact these event-study indicators, with the variable $NGO\,Campaign_{n,w,y}$, an indicator variable equal to 1 if firm n experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise.
+The vector $X_{n,w,y}$ includes the direct effect of these variables.
+The reference month is $\tau=-4$ following SEC rules on shareholder proposals.
+Note that, because of the stability of AGM dates across years within firms (Appendix Figure B1), the distribution of months fixed effects is unaffected by the re-calendarization, so that months are not double counted across years.
+Standard errors are clustered by firm.
 Figure B3: NGO campaigns and shareholder proposals
 Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead.
-The regression also includes month-by-year-by-industry-by-topic fixed effects.
+The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects.
 The regression is based on 1,227 NGO campaigns.
 Standard errors are clustered at the firm level.
-Figure B8: NGO campaigns and consumer backlash
-(a) NGO Campaigns on AGM
+Figure B4: NGO campaign timing and consumer backlash
+(a) NGO Campaigns on AGM dates
 (b) NGO Campaigns on Other Days
-(c) Estimated Effect of Campaigning on AGM
-Notes: Panels (a) and (b) plot the estimated coefficients of the regression of sales over (lagged) total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$ from (3).
-Panel(c) plots the estimated $\hat{\beta}$ from
-$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign on AGM}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)$
-$+\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\tilde{\beta}_{\tau}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
-which interacts the time dummy, the occurrence of a campaign, and a dummy indicating that the campaign happened on AGM day.
-The regressions also include firm and quarter-by-industry fixed effects.
+Notes: The figure plots the estimated ${\beta}_{\tau}$ from (3).
+Panel (a) focuses on NGO campaigns on AGM dates, while Panel (b) focuses on campaigns on dates other than AGM dates.
 The reference period is -1 quarter, consistent with the other analyses.
-The number of observations is 10,188, 351,351, and 361,539, respectively.
+The number of observations is 239,818 in Panel (a) and 239,925 in Panel (b).
 Standard errors are clustered at the firm level.
@@
 Note: This figure plots the estimated coefficients of the following regression:
-$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
-where $\Delta_{4}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter $q$ and quarter $q+4$, and $\sum_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM.
-The regression also controls for firm size (logged total assets), current average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$).
+$_{5}\text{E\&S Score}_{n,q}=\sumop\displaylimits_{l=-6}^{5}\text{Campaign}_{n,q,l}+X_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
+where $_{5}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter q1q-1 and quarter $q+4$, and $\sumop\displaylimits_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM.
+The regression also controls ($X_{n,q}$) for firm size (logged total assets), lagged average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$).
 ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
@@
 The number of observations is 636 and 698, respectively.
-Appendix C Matching Proposals And NGO Campaigns
+Figure B10: Media coverage of NGO-target firms around their AGM
+Note: To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
+$\text{News}_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\left(\beta_{\tau}\cdot\mathbf{1}_{\{M_{n,w,y}=\tau\}}\cdot\text{NGO Campaign}_{n,w,y}\right)$
+$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
+[…]
+Appendix C Robustness of Google Trends Results
+This section discusses the robustness of the results presented in Figure 2 to alternative specifications.
+In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (Sun and Abraham, 2021), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
+C.1 Staggered Difference-in-Differences Specification
+We first replicate Figure 2, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (Sun and Abraham, 2021).
+Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1.
+The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%.
+By contrast, experienced NGOs see a more modest increase (33%) in search intensity, and only in the month of the campaign.
+Figure C1: Staggered difference-in-differences specification
+Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (1) using a staggered difference-in-differences estimator (Sun and Abraham, 2021).
+The dependent variable is the monthly Google Trends score for the NGO.
+Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns.
+Standard errors are clustered at the campaign level.
+C.2 Distributed Lags Specification
+In the baseline event-study, if an NGO launches several campaigns in close succession, their effects may overlap in calendar time.
+For example, if one campaign takes place at $t=0$ and another at $t=2$, the months around $t=2$ capture the lingering impact of both events.
+This overlap can bias the estimated effects in the previous analyses.
+To address this, we estimate a distributed lags specification that explicitly includes leads and lags of both AGM and non-AGM campaigns:
+$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{g,m+\tau}\right)$ (C1)
+$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\text{Non-AGM Campaign}_{g,m+\tau}\right)$
+$\quad+\alpha_{m,y}+\alpha_{g,y}+u_{g,m,y},$
+By doing so, the regression attributes observed variation in search activity in each month to the full sequence of campaigns occurring in that window, rather than assigning the entire effect to whichever campaign happens to be closest in event time.
+Put differently, the distributed lags approach disentangles the impact of overlapping campaigns by allowing each campaign to contribute separately to visibility over several months.
+This ensures that the estimated $\beta^{AGM}_{\tau}$ captures the marginal visibility effect of a single campaign, net of any other campaigns that may have occurred nearby.
+Figure C2: Distributed lags specification
+(a) First three campaigns
+(b) Later campaigns
+Note: The two plots present the estimated $\beta^{AGM}$ with blue circles and $\beta^{Non-AGM}$ with red triangles coefficients from (C1) on two different samples.
+The first panel subsets the data to each NGO’s first three campaigns, while the second panel examines later campaings.
+The dependent variable is the monthly Google Trends score for the NGO.
+Standard errors are clustered at the campaign level.
+Figure C2 reports the estimated coefficients.
+When restricting attention to an NGO’s first three campaigns, the results confirm that AGM-timed campaigns lead to significant short-run increases in visibility (blue dots), while later campaigns generate little or no effect.
+Whether early or late, non-AGM campaings have a limited effect on an NGO’s visibility.
+Overall, these robustness checks confirm that our main results are not driven by campaign timing heterogeneity across NGOs or by overlapping campaigns within a year.
+Appendix D Matching Proposals And NGO Campaigns
 This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign cause.
@@
 We start by assigning proposals to causes based on keywords in their descriptions and then manually classify any remaining proposals.
-Table C1: Proposal Causes
+Table D1: Proposal Causes
 NGO Campaign Category Example of Shareholder Proposals

### v7

@@
 Using data on 2,500 U.S.
-campaigns, we find that NGOs frequently launch campaigns on the targeted firm’s AGM date, which boosts media exposure and triggers future shareholder proposals but does not sway current AGM votes or affect sales.
-As NGOs gain experience and reputational capital, they shift to earlier pre-AGM campaigns that increase shareholder support at the AGM and lead to improvements in corporate behavior.
-These patterns reveal a strategic trade-off between maximizing exposure and achieving direct influence.
+campaigns, we find that NGOs often campaign on the AGM date, which increases media exposure and future shareholder proposals but does not affect current AGM votes or consumers.
+As NGOs gain reputational capital, they shift to earlier pre-AGM campaigns to support related shareholder proposals and directly improve corporate behavior.
+The evidence underscores a strategic trade-off between visibility and influence, showing how NGO tactics evolve with NGOs’ credibility and the changing expectations of socially conscious stakeholders about NGO activities.
 JEL classifications: L21, L31
@@
 The NGO data come from Sigwatch; the news data are from Ravenpack.
-Examining campaign timing reveals that less experienced NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020).
+Examining campaign timing reveals that younger NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020).
 These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents.
@@
 While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves.
-Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively inexperienced.
+Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively young.
 These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times.
 This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
-As NGOs accumulate visibility through AGM-date campaigns, their strategic calculus changes.
+As NGOs accumulate visibility through AGM-date campaigns, their strategy changes.
 Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes.
-By contrast, campaigns launched in the months leading up to the AGM substantially raise the likelihood that related shareholder proposals succeed.
-Intervening one to three months before the meeting boosts average shareholder support by 8%, while even earlier campaigns often prompt firms to settle before proposals reach a vote, leading to measurable improvements in environmental and social performance.
-Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns forward in time as they gain experience and reputational capital.
+By contrast, campaigns in the months preceding the AGM raise the likelihood of proposal success.
+Campaigns one to three months before the meeting increase support by 8%, while earlier ones often induce settlements that improve firms’ environmental and social performance.
+Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns earlier in time as they gain AGM exposure.
 First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions.
-This pattern is especially pronounced among NGOs with stronger reputations, proxied by higher search visibility, which campaign several months ahead of the AGM rather than on the day itself.
+This pattern is especially pronounced among NGOs with high search intensity, which may not drive substantial visibility benefits from AGM-day campaigns.
 Repeated targeting of AGMs also plays a distinct role.
-By returning to these focal moments, NGOs build credibility and reputational standing, which in turn increases the effectiveness of later campaigns.
-Yet such returns are not indefinite: once credibility is established, additional AGM-day campaigns yield diminishing benefits.
-Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—shows how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for firms’ social and environmental impact.
-In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
-As NGOs accumulate reputational capital (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their strategy shifts: campaigns are launched months before the AGM, when they can still sway shareholder support and discipline managers.
+By returning to these focal moments, NGOs build a “reputation for toughness”—demonstrating their ability to amplify negative media coverage when firms are most exposed—which can enhance the effectiveness of subsequent campaigns.
+Yet, once reputation is established, further AGM-day campaigns yield diminishing benefits, generating smaller increases in NGO search intensity, lower success rates for related shareholder proposals, and less negative media coverage on the target firm.
+This pattern reflects NGOs’ changing incentives because of evolving stakeholder expectations.
+As NGOs become recognized and credible actors (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their actions are interpreted differently: what initially signals resolve later signals routine.
+Established NGOs are therefore expected to engage in more constructive tactics, such as supporting shareholder proposals, rather than relying on visible confrontation.
+Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—illustrates how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for their social and environmental impact.
+In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982), but the meaning and effectiveness of those signals evolve as stakeholders update their expectations about each NGO’s toughness and legitimacy.
 This dynamic lifecycle highlights NGOs’ role as coordinators that align stakeholder actions and channel them into constructive corporate change, as their strategies shift over time to balance visibility and influence (Briscoe and Gupta, 2016), with campaign tactics evolving dynamically across the life cycle to shape both corporate responses and stakeholder engagement (Cabral and Hortacsu, 2010).
-By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022).
+By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022; Allcott et al., 2022).
 More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
@@
 Standard errors are clustered at the NGO level.
-Under the hypothesis that AGM campaigns affect the visibility of experienced and inexperienced NGOs differently, we estimate (1) separately for an NGO’s first few campaigns and for later ones.
-Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and its subsequent campaigns (red triangles).
+Under the hypothesis that AGM campaigns affect NGO visibility differently depending on their reputation, we estimate (1) separately for an NGO’s first few campaigns, when its recognition is still limited, and for later ones.
+Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and subsequent campaigns (red triangles).
 The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months.
 For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established.
-These patterns suggest that inexperienced NGOs rely more heavily on AGM timing to build name recognition, whereas established NGOs gain less from repeating the same tactic.
+These results are estimated within NGOs: early in their lifecycle, NGOs benefit more from AGM timing to build visibility, but the returns to this tactic decline as their public recognition grows.
 Figure 2: Increased NGO visibility around an AGM campaign
@@
 To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year.
-Both approaches yield similar results, with much larger gains for “unexperienced” NGOs.
+Both approaches yield similar results, with substantially larger gains for early campaigns.
 Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility—for example, by focusing on firms that are more vulnerable to pressure.
@@
 We find that a one standard deviation increase in any of these variables moves the likelihood of an AGM campaign by less than 5% of a standard deviation, indicating little selection on observables.
-Further evidence on selection based on NGO experience comes from Table A1.
-Panel (c) compares firms targeted by NGOs with high versus low campaign experience and finds no significant differences across the same variables studied in Panel (b) of Figure 2.
-Panel (d) corroborates this by regressing firm characteristics on NGO experience (i.e., number of past campaigns), showing no correlation once fixed effects are included.
-The only systematic trend is that NGOs increasingly focus on environmental rather than social issues in later years, reflecting their rising importance.
+Further evidence on potential selection based on the number of past NGO campaigns is presented in Table A1.
+Panel (c) compares firms targeted by NGOs that had previously campaigned on a target’s AGM date with those that had not, finding no significant differences across the same variables analyzed in Panel (b) of Figure 2.
+This suggests that NGOs do not appear to select targets based on their past campaigning strategies.
+Panel (d) corroborates this result by regressing firm characteristics on the number of past NGO campaigns and showing no systematic relationship once fixed effects are included.
+The only consistent pattern is that NGOs increasingly focus on environmental rather than social issues over time, reflecting their growing salience.
 Thus, the evidence presented so far points to:
 Result 1.
-Gains in visibility are greater for unexperienced NGOs campaigning on the target’s AGM day than for experienced NGOs.
+Visibility gains from AGM‐day campaigns are greatest during an NGO’s initial campaigns.
 For clarity, each of the next sections highlights its main findings in a similar manner.
 4.1 Stakeholders’ Reactions
-We next ask how campaign timing matters for stakeholders and firms’ social and environmental outcomes. We focus on shareholders and consumers, with supporting exhibits in the Online Appendix.
+We next examine how campaign timing affects stakeholders and firms’ ESG outcomes.
 4.1.1 Shareholders
@@
 We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns.
-[^5] Unreported results show also no difference between experienced and unexperienced NGOs.
+[^5] Unreported results show also no difference between early and late campaigns.
 4.2 Does AGM-Targeting Create Positive Change?
@@
 However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year.
+[^6] Acknowledging the limits of ESG scores as measures of social and environmental impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), we use MSCI to mitigate signal risk: its ratings are more time-consistent and widely used by sustainable investors (Berg et al., 2021, 2024).
 By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
-Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms.
-Therefore, the results in this section suggest that:
+This finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore, the results in this section suggest that:
 Result 2.
@@
 In this section, we examine how NGOs navigate this tradeoff as their visibility increases.
-We first study whether greater experience or reputation leads NGOs to shift campaign timing from the AGM date to earlier months.
+We first study whether greater past exposure leads NGOs to shift campaign timing from the AGM date to earlier months.
 We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
-6.1 NGO Experience and Campaign Timing
-We first examine whether NGOs, as they gain experience, shift from targeting AGMs to launching earlier campaigns. Specifically, we estimate
+6.1 The Timing of Past and Current Campaigns
+We first examine whether NGOs shift from targeting AGMs to launching earlier campaigns as they campaign more. Specifically, we estimate
 $\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}$ (5)
@@
 Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7).
-Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO experience (Appendix Figures B8 and B9).
-6.2 Reputation and Campaign Timing
+Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO past AGM exposure (Appendix Figures B8 and B9).
+6.2 Visibility and Campaign Timing
 We next use Google Trends scores to proxy for an NGO’s reputation.
@@
 Panel (b) of Figure 4 shows that only the most visible NGOs—those in the top deciles of relative searches—campaign several months ahead of AGMs, while less visible NGOs continue to concentrate on AGM dates.
-Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation or experience grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
+Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
 Result 4.
-NGOs initially target AGM dates to gain visibility. As their experience or reputation increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
+NGOs initially target AGM dates to gain visibility. As their exposure increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
 Figure 4: NGO reputation and campaign timing
-(a) More experienced NGOs campaign earlier
+(a) Later campaigns pre-date the target’s AGM date
 (b) More visible NGOs campaign earlier
@@
 6.3 NGO Reputation and AGM Influence
-We now ask whether NGO experience and reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
-$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{Experience}_{g,y-1}$ (7)
-$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{Experience}_{g,y-1}$
+We now ask whether NGO reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
+$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{\# Past Campaigns}_{g,y-1}$ (7)
+$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{\# Past Campaigns}_{g,y-1}$
 $+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
@@
 The key variable $\text{Campaign on Same Cause}_{i,n,y}$” equals one if NGO $g$ campaigned against firm $n$ on cause $i$ in the ten months preceding the AGM.
-$\text{Experience}_{g,y-1}$” counts NGO $g$’s past campaigns.
+“# $\text{\# Past Campaigns}_{g,y-1}$” counts NGO $g$’s past campaigns.
 We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
-Table 1: NGO experience at targeting AGMs and success at supporting current shareholder proposals
+Table 1: NGO’s strategy and support for related shareholder proposals
 Dependent Variable: Proposal Passes (0/1)
@@
 (0.026) (0.028) (0.025) (0.030)
-Experience -0.034
-Campaign on Same Topic $\times$ Experience 0.078∗∗∗
-Experience square 0.008∗∗
-Campaign on Same Topic $\times$ Experience square -0.015∗∗
-Non-AGM Experience 0.0008
-Campaign on Same Topic $\times$ Non-AGM Experience -0.011
-AGM Experience 0.005 -0.034
+# Past Campaigns -0.034
+Campaign on Same Topic $\times$ # Past Campaigns 0.078∗∗∗
+# Past Campaigns squared 0.008∗∗
+Campaign on Same Topic $\times$ # Past Campaigns squared -0.015∗∗
+# Past Non-AGM Campaigns 0.0008
+Campaign on Same Topic $\times$ # Past Non-AGM Campaigns -0.011
+# Past AGM Campaigns 0.005 -0.034
 (0.020) (0.023)
-Campaign on Same Topic $\times$ AGM Experience 0.048 0.091∗∗∗
+Campaign on Same Topic $\times$ # Past AGM Campaigns 0.048 0.091∗∗∗
 (0.026) (0.032)
-AGM Experience square 0.010∗∗
-Campaign on Same Topic $\times$ AGM Experience square -0.013∗∗
+# Past AGM Campaigns squared 0.010∗∗
+Campaign on Same Topic $\times$ # Past AGM Campaigns squared -0.013∗∗
 Campaign on Same Topic $\times$ Google Search Popularity 0.071
@@
   • Note: This table reports the estimated coefficients from (7) in Column 1.
-In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (Non-AGM Experience) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), and their interaction with Campaign on Same Topic.
-In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square, and their interaction with Campaign on Same Topic.
-In Column 4, we replace Experience from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign.
+In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (# Past Non-AGM Campaigns) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), and their interaction with Campaign on Same Topic.
+In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), its square, and their interaction with Campaign on Same Topic.
+In Column 4, we replace # Past Campaigns from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign.
 All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects.
@@
 Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$).
-What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has accumulated experience ($\beta_{3}>0$).
+What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has already campaigned ($\beta_{3}>0$).
 A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average.
-Column (2) distinguishes between campaigns held on AGM dates and those held at other times.
-Only the former contribute to proposal success.
+The regression also includes the squared term of “# $\text{\# Past Campaigns}_{g,y-1}$,” showing decreasing marginal returns of campaigning.
+We unpack the mechanism behind these coefficient estimates by distinguishing between prior AGM campaigns and non-AGM campaigns in Column (2).
+We find that only the former contribute to proposal success.
 In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
@@
 Importantly, the combined effect remains non-negative even for the most active NGOs.
-[^8] The marginal effect of AGM Experience and its square never turns negative, so our proxy for reputation does not imply adverse returns.
+[^9] The marginal effect of AGM campaigning never turns negative in our sample since no firm launches more than five AGM-date campaigns, so our proxy for reputation does not imply adverse returns.
 We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends.
-Specifically, Column (4) replaces experience by Google Search Popularityg,y, which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.999The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects.
+Specifically, Column (4) replaces “# $\text{\# Past Campaigns}_{g,y-1}$” by “Google Search Popularityg,y-1,” which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.101010The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects.
 The interaction between this measure and campaigning on the same cause is positive, with a concave shape.
 Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
-These results also resonate with the media evidence in Section 4: AGM targeting increases negative media coverage only for relatively inexperienced NGOs.
-More seasoned NGOs gain little additional visibility from campaigning exactly on AGM dates and may be more effective when campaigning slightly earlier.
-These results support our view that:
+These results support the following result:
 Result 5.
-Targeting AGMs builds NGO reputation and increases their ability to raise the supported of related shareholder proposals, but with diminishing returns.
+Targeting AGMs builds NGO reputation and increases their ability to raise the support of related shareholder proposals, but with diminishing returns.
 7 Discussion and Conclusion
-Examining NGO campaign timing reveals that not just the decision to intervene, but its timing, can materially affect firm behavior and shape the course of future activism.
-Over time, NGOs shift from visibility-driven exposure tactics to influence current strategies, highlighting how the temporal dimension of their strategy channels pressure that leads firms to internalize social and environmental externalities.
-We first show that NGOs tend to launch their campaigns on the AGM date, likely motivated by the higher media coverage surrounding these events (Result 1).
-By doing so, NGOs elicit reactions from key stakeholders, such as shareholders.
-However, the immediate purpose of these campaigns is primarily to gain visibility by voicing stakeholder concerns; they have no direct effect on firms’ current E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
+Examining NGO campaign timing, we show that not just the decision to intervene but also its timing can materially affect firm behavior and shape the course of future activism.
+Over time, NGOs shift from visibility-driven exposure tactics to those designed to influence corporate strategies, highlighting how the temporal dimension of their actions channels pressure that leads firms to internalize social and environmental externalities.
+We first show that NGOs’ early campaigns often coincide with the target’s AGM date, when media attention is highest, and that such timing significantly increases their visibility (Result 1).
+These visibility gains are concentrated in NGOs’ first few campaigns, indicating that AGM-day actions serve primarily to build name recognition rather than to influence firms directly.
+By contrast, later AGM-day campaigns have no immediate effect on firms’ E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
 Hence, campaigning before the AGM offers direct advantages.
@@
 Furthermore, the analyses in Sections 4 and 6 show little evidence of selection on observables, alleviating concerns that NGOs disproportionately target firms that are easier to influence or more likely to generate visibility.
-Our results suggest that NGOs face a trade-off, which they navigate by adapting their strategies over time.
+Our results suggest that NGOs face a trade-off that they navigate by adapting their strategies over time.
 NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
-One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
-Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5).
-First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device.
-This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following “reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future.
-[^10] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
-Second, stakeholders may come to expect more constructive approaches from established NGOs, such as supporting related shareholder proposals.
-Consistent with this hypothesis, AGM-date campaigns by mature NGOs—those with multiple prior campaigns—do not generate more negative media coverage than campaigns at other times.
-By contrast, the very same NGOs do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B10).
-Hence, we observe not only a shift in strategy as NGOs mature (Figure 4) but also a change in their capacity to inflict reputational harm on firms.
+As a result, having targeted AGMs in the past increases the likelihood that an NGO can help a shareholder proposal pass related to its campaign.
+However, repeatedly targeting AGMs leads to weaker influence over firm behavior (Result 5).
+One interpretation is that NGOs, having already demonstrated commitment to their cause, face weaker incentives to keep signaling through costly AGM-day campaigns.
+In this view, declining returns simply reflect a reduced need for visibility once reputational capital is established.
+Yet the evidence points to a different mechanism.
+Within the same NGO, AGM-day campaigns progressively lose their marginal effect on media attention (Figure 2) and their ability to damage a firm’s reputation (Appendix Figure B10), suggesting that the audience’s response—rather than the NGO’s effort—changes over time.
+As NGOs become recognized and credible actors, stakeholders interpret their campaigns differently: what initially signals toughness later signals redundancy or routine.
+Established NGOs are expected to engage in more constructive dialogue, such as supporting shareholder proposals, rather than resorting to public confrontation.
+Consequently, the same action produces less reputational harm for firms and fewer visibility gains for NGOs.
+The diminishing returns to AGM targeting thus stem from evolving stakeholder expectations and a changing perception of NGO legitimacy, rather than from reduced internal pressure to seek exposure.
+[^11] This interpretation does not rule out signaling motives altogether.
+Early AGM-day campaigns are consistent with “reputation-for-toughness” models, in which actors bear short-term costs to establish credibility that can be leveraged in future interactions (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
 Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause.
-For instance, organizational aging (e.g., older NGOs becoming larger and more complex) is accounted for through fixed effects.
+For instance, organizational aging (e.g., older NGOs becoming larger and more complex) or experience in managing AGM-related information is accounted for through NGO-by-year fixed effects.
 These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
-In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied.
-We leverage the exogeneity of corporate events to uncover a trade-off between visibility and influence that NGOs face when deciding when to campaign, and show that how this trade-off is resolved determines which stakeholders are mobilized (e.g., shareholders today versus shareholders tomorrow).
-Hence, NGOs act as a coordination device, with distinct implications for the social and environmental impact of targeted firms.
-We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010).
-Future research could capitalize on similar exogenous variation to quantitatively study the incentives for collaboration.
+In conclusion, NGOs without an established reputation coordinate stakeholders over time by campaigning on AGM dates, drawing attention to issues that spur future engagement.
+Once reputable, they coordinate stakeholders within the same AGM cycle by supporting existing shareholder proposals and directly shaping firm decisions.
+Thus, NGOs act as a coordination device throughout their lifecycle, with the form of coordination evolving with their perceived toughness and credibility.
+In this way, NGOs’ incentives shape their strategies and, ultimately, the social and environmental impact of firms.
+We also acknowledge that NGOs increasingly engage not only in campaigns but also in collaborations with corporations (Dahan et al., 2010).
+Future research could examine when confrontation shifts to cooperation and how such collaboration reshapes the balance between visibility, credibility, and influence.
 […]
@@
 Campaign on Soc. topic 0.53 0 1 1 0.50
-Panel c: Variables varying by target firm and NGO experience
-NGO w/ AGM Experience NGO w/o AGM Experience Test difference
+Panel c: Selection—Avg. target characteristics’ based on whether the campaigning NGO campaigned on AGM dates in the past
+NGO w/ Past AGM Campaigns NGO w/o Past AGM Campaigns Test Difference
 Mean S.E. Mean S.E. p-value
@@
 Campaign on Soc. topic 0.54 0.03 0.49 0.01 0.12
-Panel d: OLS Regression of $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
+Panel d: Selection—Target’s characteristics based on NGO’s past campaigns regressing $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{\# Past Campaign}_{g(n)y}+\varepsilon_{ny}$
 $\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
@@
 Campaign on Soc. topic -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
+  • Note: This table reports summary statistics and robustness checks across four panels.
+  • Panel (a). Summary statistics on NGO campaigns (Sigwatch); AGM dates (ISS).
+  • Panel (b). Target-firm statistics: accounting (Compustat), ESG (MSCI), news counts (RavenPack).
+  • Panel (c).
+No-selection-on-observables test comparing targets where the NGO had previously campaigned on at least one of its targets’ AGM dates.
+Market Cap, Total Assets, and Sales are deciles using all U.S.
+Compustat firms in the campaign quarter.
+Deciles avoid mechanical correlation with # Past NGO Campaigns, which grows over time.
 […]
@@
 Results remain similar if we instead set month –4 as the reference period, as in the other analyses.
-Figure B6: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects
-Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
+Figure B6: NGO past campaigns and the timing of current NGO campaigns – Alternative set of Fixed Effects
+Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the number of campaigns the focal NGO had already launched.
 The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year.
-Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in NGO experience over years.
+Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in the numbers of NGO campaigns over years.
 Standard errors are clustered at the NGO level.
 Figure B7: Campaign timing for NGOs whose first campaign is after the AGM
-Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before.
+Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the number of past campaigns of the focal NGO.
 The panel focuses on NGOs whose first campaign was within the two weeks following the AGM.
 Standard errors are clustered at the NGO level.
-Figure B8: NGO experience and target firm’s size
-Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the experience of the NGO, measured by the number of campaigns it had launched in previous year as in (5).
+Figure B8: Past campaigns and target firm’s size
+Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the number of campaigns the focal NGO had launched in previous years as in (5).
 Standard errors are clustered at the firm level.
 The number of observations is 1,338.
-Figure B9: NGO experience and target firm’s E&S characteristics
+Figure B9: Past campaigns and target firm’s E&S characteristics
 (a) Target firm’s E score for Environmental campaigns
 (b) Target firm’s S score for Social campaigns
-Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating the experience of the NGO, measured by either the number of environmental (top panel) or social (bottom panel) campaigns it had launched in previous year as in (5).
+Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating either the number of environmental (top panel) or social (bottom panel) campaigns the focal NGO had launched in previous years as in (5).
 Standard errors are clustered at the firm level.
@@
 The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%.
-By contrast, experienced NGOs see a more modest increase (33%) in search intensity, and only in the month of the campaign.
+By contrast, NGOs see a more modest increase (33%) in search intensity for later campaigns, and only in the month of the campaign.
 Figure C1: Staggered difference-in-differences specification

### v8

@@
 Abstract
-We examine how NGOs time campaigns to influence corporate decisions at Annual General Meetings (AGMs).
-Using data on 2,500 U.S.
-campaigns, we find that NGOs often campaign on the AGM date, which increases media exposure and future shareholder proposals but does not affect current AGM votes or consumers.
-As NGOs gain reputational capital, they shift to earlier pre-AGM campaigns to support related shareholder proposals and directly improve corporate behavior.
-The evidence underscores a strategic trade-off between visibility and influence, showing how NGO tactics evolve with NGOs’ credibility and the changing expectations of socially conscious stakeholders about NGO activities.
+This paper studies how the timing of NGO activism shapes its effectiveness in influencing corporate behavior.
+Using data on 2,500 campaigns targeting U.S.
+firms, we show that campaigns timed at annual general meetings (AGMs) generate large visibility gains but little contemporaneous influence, while campaigns launched before the AGM significantly increase shareholder proposal success and improve firms’ environmental and social performance.
+We develop a dynamic model in which NGOs trade off awareness building and credibility formation, generating a lifecycle in activism from visibility-seeking to influence-oriented engagement.
+Therefore, NGOs’ objectives evolve endogenously to coordinate stakeholder pressure and shape corporate behavior.
 JEL classifications: L21, L31
@@
 1 Introduction
-Non-governmental organizations (NGOs) are increasingly prominent actors in shaping firms’ environmental and social policies (Baron, 2001; Egorov and Harstad, 2017; Yaziji and Doh, 2009).
-Yet, their strategic choices remain poorly understood.
-Much of the literature treats NGOs as either principled advocates motivated by social good (Besley and Ghatak, 2005; Bénabou and Tirole, 2006; Daubanes and Rochet, 2019) or as dysfunctional entities prone to corporate capture and strategic manipulation (Grossman and Helpman, 1994; Bertrand et al., 2020; Akey et al., 2021).
-We argue and demonstrate empirically that NGOs adapt their strategies in response to incentives, shaping which stakeholders they engage and ultimately affecting the social and environmental performance of targeted firms.
-To study how NGOs respond to incentives and their implications for firms, we focus on a setting where firms are especially exposed to outside pressure: the Annual General Meeting (AGM).
-AGMs are predictable focal points at which executives present results, shareholders vote on proposals, and media attention peaks (Fioretti et al., 2025).
-Figure 1 shows that NGO campaigns cluster heavily around AGM weeks: incidence is four to six times higher than in non-AGM weeks (Panel a), and media coverage spikes simultaneously (Panel b).
-These striking patterns raise a central question: do NGOs target AGMs mainly to maximize exposure, or do they also use campaigns to exert influence over corporate outcomes?
+Non-governmental organizations (NGOs) have become influential actors in shaping firms’ environmental and social policies (e.g., baron2001private; egorov2017private; yaziji2009ngos).
+Yet, we know remarkably little about their strategic behavior.
+NGOs are highly opaque: they disclose almost nothing about internal decisions, incentives, or tactics, making most of their actions effectively unobservable (maxwell2012economic).
+As a result, the literature largely portrays NGOs either as principled advocates driven by social missions (e.g., besley2005competition; benabou2006incentives; daubanes2019rise) or as organizations vulnerable to strategic misuse and capture (e.g., grossman1994protection; bertrand2020tax; akey2021hacking).
+We argue instead that NGOs respond systematically to incentives generated by stakeholder attention and expectations, shaping NGOs’ strategic choices, with consequences for corporate behavior and social and environmental outcomes.
 Figure 1: Timing of NGO campaigns and media attention around AGMs
@@
 The NGO data come from Sigwatch; the news data are from Ravenpack.
-Examining campaign timing reveals that younger NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020).
-These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents.
-For instance, as the NGO Public Eye put it in its 2025 campaign timed on the AGM date of Inditex: “As shareholders, you have the power to put Inditex on a sustainable course” (Clean Clothes Campaign, 2025).
-Indeed, we find little evidence that other stakeholders are moved by these campaigns.
-For example, we find no evidence that sales are differently affected based on campaign timing relative to the AGM date, indicating that consumers may not be the main audience of NGOs.
-While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves.
-Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively young.
-These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times.
-This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
-As NGOs accumulate visibility through AGM-date campaigns, their strategy changes.
-Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes.
-By contrast, campaigns in the months preceding the AGM raise the likelihood of proposal success.
-Campaigns one to three months before the meeting increase support by 8%, while earlier ones often induce settlements that improve firms’ environmental and social performance.
-Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns earlier in time as they gain AGM exposure.
-First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions.
-This pattern is especially pronounced among NGOs with high search intensity, which may not drive substantial visibility benefits from AGM-day campaigns.
-Repeated targeting of AGMs also plays a distinct role.
-By returning to these focal moments, NGOs build a “reputation for toughness”—demonstrating their ability to amplify negative media coverage when firms are most exposed—which can enhance the effectiveness of subsequent campaigns.
-Yet, once reputation is established, further AGM-day campaigns yield diminishing benefits, generating smaller increases in NGO search intensity, lower success rates for related shareholder proposals, and less negative media coverage on the target firm.
-This pattern reflects NGOs’ changing incentives because of evolving stakeholder expectations.
-As NGOs become recognized and credible actors (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their actions are interpreted differently: what initially signals resolve later signals routine.
-Established NGOs are therefore expected to engage in more constructive tactics, such as supporting shareholder proposals, rather than relying on visible confrontation.
-Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—illustrates how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for their social and environmental impact.
-In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982), but the meaning and effectiveness of those signals evolve as stakeholders update their expectations about each NGO’s toughness and legitimacy.
-This dynamic lifecycle highlights NGOs’ role as coordinators that align stakeholder actions and channel them into constructive corporate change, as their strategies shift over time to balance visibility and influence (Briscoe and Gupta, 2016), with campaign tactics evolving dynamically across the life cycle to shape both corporate responses and stakeholder engagement (Cabral and Hortacsu, 2010).
-By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022; Allcott et al., 2022).
-More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
-Prior studies highlight the role of shareholders influencing firms via voice through voting or exit via divestments (e.g., Cuñat et al., 2016; Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), and of consumers through willingness to pay (Gneezy et al., 2010).
-We instead show that NGOs act as coordination mechanisms, enabling diverse stakeholders to collectively discipline firms (Koenig and Poncet, 2019; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023).
-In doing so, NGOs expand the set of channels through which civil society shapes corporate conduct.
-Finally, our methodology relates to the literature on the strategic timing of actions to exploit fluctuations in media attention.
-Prior work shows that politicians and governments time executive orders, military operations, and even disaster relief to coincide with or avoid competing news (Eisensee and Strömberg, 2007; Durante and Zhuravskaya, 2018; Djourelova and Durante, 2022), and that social movements strategically leverage focal dates to amplify visibility (Madestam et al., 2013).
-We show that NGOs adopt similar tactics in the corporate domain.
-The paper is organized as follows.
-Section 2 outlines the data, and Section 3 provides background on AGM dates.
-Sections 4 and 5 examine NGO campaigns as tools to gain media exposure or to influence firms through key stakeholders, respectively.
-Section 6 brings these channels together by analyzing how NGO reputation enhances the success of shareholder proposals.
-Section 7 discusses the results and concludes.
+A rare opportunity to observe these strategic responses arises from campaign timing.
+NGO campaigns must be public to be effective, and their timing provides one of the few visible choices NGOs make.
+We focus on how NGOs align campaigns with a moment when firms are uniquely exposed to outside scrutiny: the Annual General Meeting (AGM).
+AGMs are predictable focal events at which executives report results, shareholders vote, and media attention peaks (fioretti2024shared).
+Panel (a) of Figure 1 shows that NGO campaigns are five to six times more likely to occur in AGM weeks than at other times, and Panel (b) shows that media coverage of targeted firms peaks at the same moment.
+These patterns raise a central question: do NGOs target AGMs primarily to maximize exposure, or do they use campaign timing to exert influence on firms?
+Is there a tradeoff between exposure and influence?
+More broadly, what incentives shape NGO strategy, and how do their choices affect the firms they target?
+Leveraging the AGM as a predictable focal point, we address these questions and uncover a lifecycle in NGO activism.
+Younger NGOs concentrate on AGM-day campaigns that build visibility, while more established NGOs shift to pre-AGM campaigns that influence shareholder decisions and firms’ environmental and social performance.
+A simple dynamic model, in which NGOs trade off awareness building against credibility formation, rationalizes this timing pattern.
+Taken together, our results show that NGO behavior is governed by incentives that determine which stakeholders are mobilized and when, and thus whether campaigns generate symbolic exposure or substantive corporate change.
+In particular, because most shareholder votes are cast before the AGM, campaigns timed on the meeting date primarily affect visibility, whereas campaigns launched earlier can influence voting outcomes – linking NGO effectiveness to institutional features such as AGM and proxy rules and disclosure timing.
+Imagine an NGO uncovers troubling information about a firm and seeks to change its behavior.
+Because Annual General Meetings (AGMs) occur at predictable times, the NGO can choose whether to campaign before the AGM—when shareholder votes and negotiations are still open—or at the AGM itself, when public and media attention peak but most proxy votes have already been cast.
+If NGOs simply campaigned as soon as concerns arose, we would not expect the sharp clustering of campaigns around AGMs shown in Figure 1.
+The pronounced spike instead suggests that campaign timing is a strategic choice.
+To study this choice, we assemble a novel dataset of roughly 2,500 campaigns by international NGOs targeting U.S.
+firms, linked to media coverage, Google search activity, NGO financial statements, shareholder proposals, and environmental and social (E&S) outcomes.
+SEC rules tie proxy filings and AGM dates to the fiscal year, so each firm’s AGM recurs within a narrow window each year, creating a predictable surge in visibility alongside a fixed deadline for shareholder voting.
+This institutional structure generates sharp variation in when campaigns occur relative to when shareholder decisions are made, which we exploit to distinguish exposure-oriented from influence-oriented activism and their drivers.
+Our first set of results shows that campaign timing is tightly linked to how NGOs affect firms.
+In their first campaigns, NGOs tend to target the AGM date, generating large visibility shocks – raising media coverage, online search activity, donations, and the likelihood that related shareholder proposals are filed at the next AGM – but having little effect on contemporaneous proposal success or firms’ E&S performance.
+This pattern suggests that shaming campaigns alone may have limited influence on corporate decisions.
+As NGOs gain experience and reputation, campaigns shift into the months before the AGM, when proxy voting and negotiations are still open.
+These earlier actions significantly increase the success and vote share of E&S proposals and are followed by improvements in firms’ E&S scores, consistent with genuine changes in corporate policy rather than pure reputational damage.
+These patterns are not explained by NGOs gradually moving from easier to harder targets: observable characteristics of target firms do not change systematically as NGOs accumulate campaigns.
+To address selection on unobservables, we exploit quasi-exogenous variation from the publication date of the Fashion Transparency Index, which is fixed by the Rana Plaza anniversary and thus pins down both when campaigns occur and which firms are targeted.
+Using the distance between index publication and firms’ AGM dates as a source of exogeneous timing variation, we find similar effects both in signs and magnitudes to those estimated in the whole dataset.
+We then show that NGOs’ use of AGM visibility versus pre-AGM influence evolves systematically over their organizational life cycle.
+NGOs with little prior AGM exposure or public visibility concentrate campaigns on the AGM date, where media attention peaks and visibility gains are largest.
+As NGOs accumulate visibility – proxied by past AGM campaigns, search intensity, and organizational size – they progressively shift campaigns earlier in the AGM cycle, when shareholder votes have not yet been cast and managerial concessions are more likely.
+NGO income statement data further indicate that early campaigns are costlier and less visibility-oriented than AGM-day campaigns.
+For younger NGOs, AGM-day actions generate sizable short-run increases in donations and returns on assets, consistent with shaming campaigns being relatively inexpensive and effective visibility tools.
+For more established NGOs, the marginal fundraising benefit of AGM campaigns is small, and non-AGM campaigns do not produce comparable financial gains for any group.
+The empirical patterns point to a sharp distinction between campaigns launched before the AGM and those timed on the meeting date.
+Campaigns before the AGM are associated with greater influence on shareholder votes and firm outcomes, while AGM-day campaigns primarily generate visibility.
+We interpret these findings through a dynamic model in which an NGO, shareholders, and the firm interact repeatedly over time, and in which campaign timing is the NGO’s central strategic choice.
+In each period, the public state is summarized by two variables.
+The first is an awareness stock, which captures how salient the issue is among dispersed shareholders and the broader public.
+Awareness increases with highly visible actions—most notably campaigns timed on the AGM—and decays over time in the absence of renewed attention.
+The second is a credibility belief, which reflects stakeholders’ assessment of the NGO’s resolve and effectiveness in sustaining pressure on firms.
+This belief evolves endogenously as shareholders and firms observe the NGO’s actions and their outcomes.
+Given the current state, the NGO chooses whether to campaign before shareholder votes are cast or at the AGM.
+AGM-day campaigns maximize visibility and generate warm-glow benefits for the NGO, but they occur after most proxy votes are locked in and therefore have little contemporaneous influence on firm decisions.
+Early campaigns, by contrast, are less visible and forego some warm-glow benefits, but they can affect shareholder voting and negotiations with the firm.
+Importantly, early campaigns are also more informative: they require sustained engagement and, when successful, reveal persistence and competence, strengthening the NGO’s credibility.
+AGM-day campaigns are comparatively cheap and visible, but once an NGO is established they convey little new information about its type.
+Shareholders observe campaign timing and, when relevant, campaign outcomes.
+Based on the current levels of awareness and credibility, they decide whether to file a proposal and how much support to give it.
+Higher awareness increases the likelihood that shareholders notice and coordinate around an issue, while higher credibility raises the expected effectiveness of activism and the probability of success.
+The firm, anticipating these responses, adjusts its behavior accordingly.
+This dynamic interaction yields a simple equilibrium structure.
+When credibility and awareness are low, the NGO optimally campaigns at the AGM: visibility and warm-glow benefits dominate, and early engagement is unlikely to be persuasive.
+As awareness accumulates and credibility improves, the NGO crosses a threshold at which the influence and reputational gains from early campaigning outweigh the loss in visibility.
+At that point, it switches to pre-AGM campaigns to exert direct pressure on shareholder decisions and firm policies.
+This cutoff behavior reflects evolving stakeholder expectations: actions that initially signal resolve later appear routine, leading established NGOs to rely less on public shaming and more on early, substantive engagement (baur2011moral; mitchell2017reputations).
+A purely visibility-driven model would predict persistent AGM-day campaigning.
+Matching the observed shift from late, high-visibility actions to earlier, influence-oriented engagement instead requires a dynamic reputation-building mechanism in the spirit of intertemporal signaling models (kreps1982reputation; milgrom1982limit; cabral2010dynamics), combined with a deterministic awareness stock that separates visibility-driven exposure from credibility-driven influence.
+This framework provides a unified explanation for why NGOs initially seek attention and later convert reputation into corporate impact.
+Our analysis yields three main contributions.
+First, we offer a new quantitative framework to examine NGO strategy and to infer how NGOs respond to incentives, which stakeholders they prioritize at different stages, and how their objectives evolve.
+Classic corporate-finance approaches tie organizational objectives to ownership structure (e.g., grossman1977value; grossman1979theory), while stakeholder theory emphasizes salience based on power, legitimacy, and urgency (freeman1984strategic; mitchell1997toward).
+Our results show that NGO objectives shift with how donors and investors interpret NGOs’ actions and with credibility concerns about timing.
+In this sense, we provide a quantitative setting to study how organizational objectives emerge from interactions with multiple stakeholders rather than being fixed by a single principal—for example, an entrepreneur defining the firm’s goals (e.g., scott2002love; romer2006firms; duarte2025competitive; fioretti2024shared).
+[^1] Our approach also relates to work on the strategic timing of actions around news cycles and focal dates (eisensee2007news; madestam2013do; durante2018attack; djourelova2022media), but shifts the focus to NGOs in the corporate domain and to the management of reputational capital over time.
+Second, we contribute to the corporate-finance and stakeholder-governance literatures by identifying mechanisms through which firms partially internalize social and environmental externalities (fioretti2022caring; allcott2022economic).
+Our setting complements work on shareholder activism and gadflies (e.g., brav2008hedge; gantchev2021costs; malenko2024information): NGOs do not vote but act as external coordinators that put issues on the ballot and help align dispersed investors’ behavior, shaping both the filing and success of shareholder proposals (see also chuah2024shareholder; mazet2023some; mcgahan2023new).
+Third, our approach has implications for policy.
+For boards, early, credibility‐backed campaigns are informative signals that warrant engagement rather than dismissal (dimson2015active; kolbel2017media).
+For investors and stewardship codes, credible NGOs supply information that helps coordinate dispersed shareholders on E&S issues, complementing traditional voice and exit (oehmke2020theory; gans2021exit; broccardo2020exit; saint2024exit; bonnefon2025moral).
+For regulators, rules on AGM timing, proposal access, and disclosure shape the returns to NGO campaigns and thus the scope for private activism to substitute for, or complement, formal regulation.
+The paper proceeds as follows.
+Section 2 presents the model.
+Section 3 describes the data, and Section 4 outlines the empirical strategy.
+Sections 5 and 6 examine NGO campaigns as tools for gaining media exposure and influencing firms through key stakeholders.
+Section 7 brings these channels together by analyzing how NGO reputation enhances the success of shareholder proposals, and Section 8 provides evidence from the quasi-experimental case study of the Fashion Transparency Index.
+Section 9 discusses implications for policy and future research, and Section 10 concludes.
+2 Model
+This section develops a model in which the timing of an NGO campaign conveys information about the NGO’s credibility and shifts the salience of the underlying issue among shareholders.
+Two forces interact: warm glow, which makes high‑visibility AGM campaigns intrinsically valuable, and reputation, which evolves by Bayesian learning as timing and proposal outcomes are observed.
+2.1 Environment and Timing
+Time is discrete, $t=0,1,2,\ldots$.
+The public state at the start of period $t$ is $(A_{t},\pi_{t})$, where $A_{t}$ denotes the stock of awareness, or issue salience among dispersed shareholders on the topic championed by the NGO, and $\pi_{t}\in[0,1]$ is the belief that the NGO is the credible type ($\theta=C$) rather than the weak one ($\theta=W$).
+A credible NGO can coordinate investors, sustain pressure, and translate outreach into voting support, whereas a weak NGO lacks this capacity and therefore exerts little direct influence on the target firm.
+Timeline.
+Given $(A_{t},\pi_{t})$, this is the order of play in each period $t$:
+  1. 1.
+    Filing decision.
+A (representative) shareholder decides whether to file a proposal for the AGM that takes place at the end of period $t$.
+A proposal is filed if the perceived probability of success covers a fixed cost $f$.
+  2. 2.
+    Campaign timing decision. The NGO chooses $d_{t}\in\{\mathrm{early},\mathrm{AGM}\}$. Early campaigns occur before proxy votes are locked; AGM campaigns occur when most votes are already cast.
+  3. 3.
+    Outcome. If a proposal was filed, the proposal is voted and an outcome $y_{t}\in\{0,1\}$ is realized at the end of period $t$.
+  4. 4.
+    Updating. Observers see $d_{t}$ and, if a proposal was on the ballot, $y_{t}$. Awareness and beliefs update to $(A_{t+1},\pi_{t+1})$.
+Awareness law of motion.
+The awareness stock $A_{t}$ summarizes issue salience among dispersed shareholders and evolves according to
+$A_{t+1}=\rho\cdot A_{t}+(1-\rho)\cdot v(d_{t}),\qquad A_{0}=0,\quad\rho\in(0,1),\quad\text{ and }\quad v(\text{AGM})>v(\text{early}),$ (1)
+so that awareness is a convex combination between the visibility of the chosen timing, $v(d_{t})$, and past awareness.
+The function $v(d_{t})$ captures the visibility of the issue–firm link among shareholders: the condition $v(\text{AGM})>v(\text{early})$ implies that awareness increases when the NGO campaign aligns with the AGM, when firms are most exposed to media and investor scrutiny, while the others are regularity conditions.
+NGO preferences.
+Warm glow from visibility and contemporaneous influence on shareholder outcomes jointly determine one-period payoffs. The per-period utility is
+$u(d_{t};\pi_{t},A_{t})=\alpha\cdot v(d_{t})+B\cdot{}_{d_{t}}(\pi_{t},A_{t})-C,$ (2)
+[…]
+$_{\mathrm{early}}(\pi,A)>{}_{\mathrm{AGM}}(\pi,A)=0,\qquad\frac{\partial{}_{\mathrm{early}}}{\partial\pi}(\pi,A)>0,\qquad\frac{\partial{}_{\mathrm{early}}}{\partial A}(\pi,A)\geq 0.$ (3)
+Hence, AGM influence is normalized to zero because most shareholders have already cast their votes, leaving little scope for NGOs to affect outcomes.
+By contrast, early campaigns have greater influence: higher credibility $\pi$ strengthens their effect since credible NGOs exert greater pressure on firms and investors, whereas awareness $A$ enhances influence indirectly by broadening the audience.
+The effect of awareness is therefore weak and may disappear once the issue is already widely known (i.e., $\frac{\partial^{2}{}_{\text{early}}}{\partial A^{2}}\leq 0$).
+[…]
+[…]
+[…]
+[…]
+Success probability and realized outcome.
+When a shareholder proposal is on the ballot, its likelihood of passing depends on three factors: how aware shareholders are of the issue, how credible they perceive the NGO to be, and whether the campaign occurs early enough to influence votes.
+We model the probability of a successful proposal for an NGO of type $\theta$ as a logistic function,
+$S^{\theta}(A_{t},\pi_{t},d_{t})=\frac{\exp\big(h_{0}+h_{A}A_{t}+h_{\pi}\pi_{t}+h\,{}_{d_{t}}^{\theta}(A_{t},\pi_{t})\big)}{1+\exp\big(h_{0}+h_{A}A_{t}+h_{\pi}\pi_{t}+h\,{}_{d_{t}}^{\theta}(A_{t},\pi_{t})\big)},$ (4)
+where $h_{0}$ is a baseline term, $h_{A}>0$ measures how greater awareness $A_{t}$ increases the baseline support among shareholders, $h_{\pi}>0$ captures the effect of credibility $\pi_{t}$ on the likelihood of success, and $h>0$ scales the influence due to the campaign’s timing.
+[…]
+The realized proposal outcome is the only random element in the model:
+$y_{t}=\begin{cases}1,&\text{with probability }S(A_{t},\pi_{t},d_{t}),\\[3.0pt] 0,&\text{with probability }1-S(A_{t},\pi_{t},d_{t}).\end{cases}$ (5)
+Thus, $y_{t}=1$ represents a successful proposal – one that either passes or achieves a favorable settlement before the AGM – and occurs more frequently when campaigns are early, credibility is high, or awareness is widespread.
+In contrast, unsuccessful proposals ($y_{t}=0$) provide weaker evidence of toughness and slow reputation accumulation.
+Filing rule.
+Before the AGM takes place, shareholders decide whether to file a proposal.
+Filing is worthwhile when the expected benefit of passage exceeds the fixed cost $f>0$.
+Let $V_{P}>0$ denote the value of a successful proposal to a shareholder, and let $\hat{d}_{t}$ denote the NGO’s anticipated timing implied by its policy.
+Because timing is a deterministic function of the state $(A_{t},\pi_{t})$, shareholders can forecast it.
+A proposal is filed if and only if
+$V_{P}\cdot S(A_{t},\pi_{t},\hat{d}_{t})\geq f,$ (6)
+where since the shareholder does not know the true type $\theta$, $S(\cdot)$ is the perceived success probability, a belief-weighted average of the type-specific probabilities $S^{C}(\cdot)$ and $S^{W}(\cdot)$ in (4), i.e., $S(\cdot)\ =\ \pi_{t}\cdot S^{C}(\cdot)+(1-\pi_{t})\cdot S^{W}(\cdot)$.
+This rule makes filing completely determined by the current state: greater awareness and credibility increase $S(A_{t},\pi_{t},\hat{d}_{t})$ and therefore the likelihood that filing is worthwhile.
+AGM timing builds awareness and encourages future filing, whereas early timing boosts contemporaneous success without necessarily raising next-period awareness.
+2.2 Belief Updating
+After observing the NGO’s campaign timing $d_{t}$ and the outcome $y_{t}$ of the shareholder proposal, shareholders update their belief $\pi_{t+1}$ about the NGO’s credibility.
+The posterior depends on two sources of information: the timing choice itself, which is strategically informative, and the realized outcome, which is stochastic but type dependent.
+Let $p^{C}(d_{t},y_{t}\mid\pi_{t},A_{t})$ and $p^{W}(d_{t},y_{t}\mid\pi_{t},A_{t})$ denote the likelihood of observing $(d_{t},y_{t})$ when the NGO is of the credible type ($C$) or weak type ($W$), respectively.
+By Bayes’ rule, the posterior belief that the NGO is credible is
+$\pi_{t+1}=\frac{\pi_{t}\cdot p^{C}(d_{t},y_{t}\mid\pi_{t},A_{t})}{\pi_{t}\cdot p^{C}(d_{t},y_{t}\mid\pi_{t},A_{t})+(1-\pi_{t})\cdot p^{W}(d_{t},y_{t}\mid\pi_{t},A_{t})}.$
+Dividing by $1-\pi_{t+1}$ and taking logs reveals the following belief updating rule, [^2] Note that the joint likelihood factors as $p^{\theta}(d_{t},y_{t}\mid\pi_{t},A_{t})=p^{\theta}(y_{t}\mid d_{t},\pi_{t},A_{t})\cdot p^{\theta}(d_{t}\mid\pi_{t},A_{t})$ for $\theta\in(W,C)$.
+$\ln\left(\frac{\pi_{t+1}}{1-\pi_{t+1}}\right)-\ln\left(\frac{\pi_{t}}{1-\pi_{t}}\right)=\ln\frac{p^{C}(d_{t}\mid\pi_{t},A_{t})}{p^{W}(d_{t}\mid\pi_{t},A_{t})}+\ln\frac{p^{C}(y_{t}\mid d_{t},\pi_{t},A_{t})}{p^{W}(y_{t}\mid d_{t},\pi_{t},A_{t})}.$ (7)
+The first term on the right-hand side captures what observers can infer from the NGO’s timing choice $d_{t}$ alone.
+Because the payoff advantage of early campaigning increases with credibility ($\partial{}_{\mathrm{early}}/\partial\pi>0$), a more credible NGO finds early timing relatively more attractive.
+In equilibrium, therefore, early campaigns occur more often when credibility is high, making timing itself an informative signal.
+We parameterize the informativeness of this timing signal as
+[…]
+[…]
+[…]
+[…]
+The second likelihood term in (7) captures the information contained in the realized outcome.
+Conditional on $(A_{t},\pi_{t},d_{t})$, the probability of proposal success is given by (4).
+For notational convenience, write $S^{\theta}(A_{t},\pi_{t},d_{t})\equiv S(A_{t},\pi_{t},d_{t}\mid\theta)$ for the success probability under type $\theta$.
+In the underlying outcome process, early campaigning is more effective for credible NGOs, so that $S^{C}(A_{t},\pi_{t},\mathrm{early})>S^{W}(A_{t},\pi_{t},\mathrm{early})$, while at the AGM, where $_{\mathrm{AGM}}\equiv 0$, both types perform identically and $S^{C}=S^{W}$.
+Hence, outcomes are informative about credibility only when the campaign occurs early.
+Given this Bernoulli data-generating process, Appendix A.1 derives the outcome component of the likelihood ratio that drives belief updating as,
+$\ln\frac{p^{C}(y_{t}\mid d_{t},A_{t},\pi_{t})}{p^{W}(y_{t}\mid d_{t},A_{t},\pi_{t})}=y_{t}\cdot\ln\frac{S^{C}(A_{t},\pi_{t},d_{t})}{S^{W}(A_{t},\pi_{t},d_{t})}+(1-y_{t})\cdot\ln\frac{1-S^{C}(A_{t},\pi_{t},d_{t})}{1-S^{W}(A_{t},\pi_{t},d_{t})}.$
+Since $S^{C}=S^{W}$ at the AGM, outcomes then carry no additional information, and beliefs update only through the timing term.
+When the campaign is early, $S^{C}>S^{W}$, so a successful proposal ($y_{t}=1$) increases $\pi_{t+1}$, while failure ($y_{t}=0$) lowers it.
+Intuitively, early campaigns are reputation-building experiments: they can signal credibility, but they also expose the NGO to reputational risk if they fail.
+Appendix A.2 shows that posterior beliefs are consistent under standard regularity conditions.
+2.3 Dynamic Program
+At each date $t$, the NGO chooses its campaign timing to balance the short-run benefit of visibility against the expected influence on shareholder behavior and the evolution of its reputation.
+Given beliefs $(\pi_{t},A_{t})$, the organization anticipates that an AGM campaign will provide the warm-glow payoff from visibility but no direct effect on voting, while an early campaign will generate a smaller visibility payoff today but can influence votes immediately and improve its reputation through a more informative outcome.
+Let $V(\pi_{t},A_{t})$ denote the NGO’s continuation value. An NGO’s recursive problem is
+$V(\pi_{t},A_{t})=\max_{d_{t}\in\{\mathrm{early},\mathrm{AGM}\}}\Big\{\alpha\,v(d_{t})+B\,{}_{d_{t}}(\pi_{t},A_{t})-C+\beta\,\mathbb{E}_{y}\big[V(\pi_{t+1},A_{t+1})\mid\pi_{t},A_{t},d_{t}\big]\Big\},$ (8)
+where $\beta\in(0,1)$ is the discount factor, and expectations are taken with respect to the only stochastic element in the model – the realized proposal outcome $y_{t}$.
+Awareness follows (1), while beliefs evolve through Bayesian updating as explained below.
+2.4 Equilibrium Timing and Cutoffs
+The NGO’s equilibrium policy specifies, for every state $(A_{t},\pi_{t})$, whether it campaigns early or at the AGM. Define the action-value difference
+[…]
+$\quad+\beta\cdot\big(\mathbb{E}_{y}\big[V(\pi_{t+1},A_{t+1})\mid\pi_{t},A_{t},\mathrm{early}\big]-\mathbb{E}_{y}\big[V(\pi_{t+1},A_{t+1})\mid\pi_{t},A_{t},\mathrm{AGM}\big]\big),$
+where expectations are taken over next-period beliefs and awareness determined by (1) and (7).
+The first term in brackets is the current payoff wedge: it weighs the incremental influence and lower visibility of early campaigning.
+The second term is the continuation-value wedge: early campaigning improves future payoffs by shifting beliefs and, through reputation, increasing expected influence in later periods.
+[…]
+Proposition 1 (Separation by Campaign Timing).
+There exists a unique cutoff function $\bar{\pi}(A)\in(0,1)$ such that the NGO’s optimal campaign timing is
+$d_{t}^{\star}=\begin{cases}\mathrm{AGM},&\pi_{t}<\bar{\pi}(A_{t}),\\ \mathrm{early},&\pi_{t}>\bar{\pi}(A_{t}).\end{cases}$
+The cutoff is weakly decreasing in awareness, $\bar{\pi}^{\prime}(A)\leq 0$. Moreover, in equilibrium:
+  1. (i)
+    For $\pi_{t}<\bar{\pi}(A_{t})$, both credible and weak NGOs pool on AGM campaigns.
+  2. (ii)
+    For $\pi_{t}>\bar{\pi}(A_{t})$, the credible NGO separates by campaigning early, while the weak NGO continues to campaign at the AGM.
+Proof. See Appendix A.3.
+The intuition is straightforward.
+When credibility is low, early campaigning is too costly to be persuasive: stakeholders do not expect weak NGOs to follow through, and AGM visibility provides immediate warm-glow benefits for both types.
+As a result, both credible and weak NGOs pool on AGM timing.
+As credibility rises, however, early campaigning becomes informative.
+For the credible NGO, the influence and reputational benefits of early action outweigh the visibility advantage of AGM campaigns, leading it to switch to early timing.
+The weak NGO, which derives no direct influence payoff from early engagement ($^{W}_{\mathrm{early}}=0$), does not find this deviation profitable and remains at the AGM.
+Importantly, although at very high beliefs early campaigning would be optimal for any NGO in a static sense, only the credible NGO reaches this region in equilibrium.
+Its reputation is built through successful early engagement, after which timing serves a purely instrumental role.
+The weak NGO, by contrast, remains trapped in the low-credibility region where visibility dominates and continues to rely on AGM campaigns.
+Thus, the equilibrium exhibits persistent separation: early campaigns are a credible signal of influence, not a pooling outcome.
+2.5 Discussion and Empirical Implications
+The model links the strategic timing of NGO campaigns to the joint behavior of NGOs and shareholders.
+The cutoff policy $\bar{\pi}(A)$ determines when an NGO switches from AGM to early campaigning: as credibility $\pi$ and awareness $A$ rise, the marginal value of early influence increases and the threshold $\bar{\pi}(A)$ declines.
+This dynamic generates a lifecycle pattern in which young NGOs focus on visibility and awareness building, while mature NGOs campaign earlier to exert direct influence.
+The model yields several testable implications for both NGOs and shareholders.
+AGM campaigns, which maximize visibility, expand awareness ($A_{t}$) but do not directly affect contemporaneous success rates since $_{\mathrm{AGM}}(\pi,A)\equiv 0$.
+Their impact on future shareholder filings depends on both the awareness and the credibility channels as captured by the success function $S(A_{t},\pi_{t},d_{t})$ in (4).
+Because $S$ increases in both $A_{t}$ and $\pi_{t}$, higher awareness raises the baseline likelihood that shareholders notice and support related proposals, while greater credibility increases the perceived probability that the NGO’s engagement will be effective.
+The relative strength of these channels is governed by the parameter vector $h$, which is left unspecified.
+Hence, whether shareholder filings are primarily driven by awareness or by credibility remains an open question within the model, depending on how visibility and reputation interact to shape $S(A_{t},\pi_{t},d_{t})$.
+Early campaigns, by contrast, work through the credibility channel.
+They generate informative outcomes and update beliefs via (7), raising $\pi_{t}$ when $y_{t}=1$ and thereby increasing expected proposal success in subsequent periods.
+Because the effect of timing enters success probabilities through $_{\mathrm{early}}(\pi,A)>0$, early campaigns also enhance contemporaneous influence: credible NGOs that campaign early obtain higher vote support and a greater likelihood of success within the same AGM cycle but do not affect future filing.
+Over time, credibility accumulates endogenously.
+The average success rate of an NGO’s campaigns therefore rises with its reputation $\pi_{t}$, which reflects the history of its visible and early actions.
+New NGOs that have campaigned mostly at AGMs build awareness but remain less credible, whereas established NGOs that have undertaken successful early campaigns achieve higher influence on both current and future outcomes.
+As awareness and credibility jointly accumulate, the cutoff $\bar{\pi}(A)$ declines, prompting NGOs to shift gradually from AGM to early timing.
+This transition generates a lifecycle pattern in which young NGOs cluster around AGM dates to gain visibility, while mature NGOs increasingly campaign early to exert direct influence.
+Early successes accelerate this process by raising $\pi_{t}$, whereas early failures slow reputation building, producing persistent heterogeneity in influence and effectiveness across NGOs.
+In this simple model, only credible NGOs switch to early campaigns.
+Our framework abstracts from a potential third region in which both credible and weak NGOs would eventually choose early campaigning.
+Such a region can arise in richer structures where weak NGOs derive some direct influence payoff or gain utility from being perceived as credible.
+Whether and when this occurs, however, depends entirely on functional-form assumptions about influence and visibility [^3] In richer specifications, weak NGOs could obtain a small influence payoff (e.g.
+$^{W}_{\mathrm{early}}(A)>0$) or benefit from being perceived as credible, implying that both types might eventually choose early timing.
+The location of such a region, however, depends on functional-form details we cannot discipline empirically.
+We therefore impose the simplifying assumption $^{W}=0$ and focus on the empirically relevant transition in which only credible NGOs switch to early campaigns.
+By imposing $^{W}=0$, we emphasize the empirically relevant transition from pooled AGM campaigning at low credibility to early campaigning by credible NGOs as reputation accumulates.
+Interpretation.
+The model follows the spirit of reputation-for-toughness models (kreps1982reputation; milgrom1982limit), where players of uncertain type use costly actions to build a reputation for persistence.
+Here, early campaigning is the costly signal: it forgoes short-term visibility and warm-glow utility (or donations) for long-term influence and reputational gains.
+AGM campaigns, by contrast, are cheap – they provide exposure but little information.
+Introducing the deterministic awareness stock extends this logic by separating visibility-driven exposure (awareness) from credibility-driven reputation.
+Both dimensions are necessary to reproduce the observed timing shift: with only warm glow and awareness, NGOs would campaign persistently at AGMs; adding reputation explains the empirical transition from AGM visibility to early influence as NGOs mature.
 3 Data
 Our dataset integrates four sources: NGO campaigns, online attention, shareholder proposals, and firm-level news.
-First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020).
+First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations from 2010 to 2021 (hatte2020geography).
 The data record the campaign’s public release date, cause (e.g., “GMOs in food”), the targeted company and its prominence, and sentiment (ranging from –2 to +2, typically negative).
@@
 Second, we collect monthly Google Trends scores for all NGOs in our sample from 2010 to 2022.
-Scores are normalized on a 0–100 scale, either (i) relative to the NGO’s own peak searches (within-NGO) or (ii) relative to a benchmark NGO in the same month (between-NGOs).
-We use Greenpeace—the most searched NGO—as the benchmark for comparability across organizations.
+Scores are normalized on a 0–100 scale, either (i) relative to the NGO’s own peak searches or (ii) relative to a benchmark NGO in the same month.
+We use Greenpeace – the most searched NGO – as the benchmark for comparability across organizations.
 Third, we use data from Institutional Shareholder Services (ISS) on proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms.
@@
 Both the number of E&S proposals and their success rate increased steadily over the period.
-Lastly, we obtain firm-level news from RavenPack, which transforms unstructured news into structured data with sentiment and event tagging.
+Fourth, we obtain firm-level news from RavenPack, which transforms unstructured news into structured data with sentiment and event tagging.
+[^5] RavenPack News Analytics assigns tone scores to news articles drawn from over 40,000 sources, such as Dow Jones Newswires, the Wall Street Journal, MarketWatch, Factiva, and Barron’s, and identifies articles linked to specific events.
+The validity of this score has been established in prior research (e.g., bushman2024influence).
+The dataset also differentiates between media-generated coverage and firm-issued press releases, enabling us to focus on the former.
 We count news items related to US-listed firms from 2010 to 2020, retaining only items with a relevance score of at least 75%.
 The final sample includes 5,212 firms.
+Lastly, we collect balance sheet and income statement data for US NGOs in the sample from their Internal revenue Service form 990, available on the National Center for Charitable Statistics website.
+Form 990 is the annual information return that most tax-exempt nonprofit organizations in the United States must file with the IRS to disclose their finances, governance practices, and programmatic activities.
+Filing obligations are primarily determined by organizational size and type: tax-exempt entities with gross receipts over \$200,000 and assets over \$500,000 file the full Form 990.
+From our starting sample, we are able to link 115 NGOs to IRS filed forms.
 We link Sigwatch, ISS, and RavenPack data using standard identifiers (CUSIP, ISIN, year).
-Appendix Table A1 reports descriptive statistics.
+Table 1 reports descriptive statistics.
 Panel (a) shows that, on average, NGOs launch 3.2 campaigns during the sample, 24% of which are on the target AGM date, though over half of NGOs appear only once.
@@
 53% of the campaigns are on social causes, with the remaining being environmental.
-We return to the remaining panels of Table A1 in Section 4, after first providing background on AGM dates.
-3 Institutional Background on AGM Dates
-Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations.
-Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014).
-Appendix Figure B1 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs.
-[^2] Results remain robust excluding the covid period (2020–2021).
-While external events may shape NGOs’ incentives to campaign, they do not alter the timing of AGMs.
-This stability makes AGMs a useful benchmark: NGOs can choose to react immediately to a negative event concerning a firm, or alternatively wait and align their campaign with its AGM, when the firm is most exposed to scrutiny, as shown in Panel (b) of Figure 1, which shows that news coverage spikes in AGM weeks.
-Consistent with this view, we show in the Appendix that AGM-timed campaigns generate more negative news coverage than those launched at other times (Appendix Figure B2).
-Taken together, these patterns highlight why AGMs provide a natural focal point to study the strategic trade-off between maximizing visibility and exerting direct influence.
-4 Timing Campaigns for Visibility
-Having shown that NGO campaigns cluster around AGM dates and that they increase a target’s negative media coverage, we next ask whether such timing enhances NGOs’ visibility.
-We use Google search activity as a proxy for NGO visibility and estimate event-study regressions comparing campaigns launched on the AGM date with those launched at other times:
-$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$ (1)
-$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$
-$\quad+\alpha_{c,y}+\alpha_{m,y}+u_{g,m,y}.$
-Because NGOs typically campaign multiple times, we follow Baker et al.
-(2022a) and stack all campaign events.
+We return to the remaining panels of Table 1 in Section 5, after first providing background on AGM dates.
+Table 1: Summary Statistics
+Panel a: Variables varying by NGO
+Mean 25% 50% 75% 95%
+Number of campaigns by NGO 3.20 1 1 3 13
+     – Campaigns on Env. topic 1.51 0 1 1 5
+     – Campaigns on Soc. topic 1.70 0 1 1 6
+     – On AGM date 0.24 0 0 0 1
+Campaign sentiment -1.72 -2 -2 -1 -1
+Form 990-filing U.S. NGO 0.15 0 0 0 1
+     – Number of campaigns 10.19 1 3 14 33
+     – Campaigns on Env. topic 4.36 0 1 4 19
+     – Campaigns on Soc. topic 2.83 0 1 4 12
+     – On AGM date 0.28 0 0 0 1
+Campaign sentiment -1.75 -2 -2 -2 -1
+NGO Total Assets (USD mn) 50.02 2.80 7.95 53.91 248.02
+Yearly donations received (USD mn) 38.03 3.18 10.13 59.14 143.09
+NGO age (years) 42.79 21 32 47 126
+Panel b: Variables varying by target firm
+Mean 25% 50% 75% SD
+Market Cap (USD bn) 175.66 21.71 75.05 209.47 263.66
+Total Assets (USD bn) 98.81 21.80 59.27 138.68 124.36
+Yearly Sales (USD bn) 37.26 6.85 17.35 42.41 150.58
+Yearly Number of News 1,183 328 616 1,046 1,695.25
+Environmental score 5.66 4 5.58 7.3 2.22
+Social score 4.44 3.30 4.70 5.60 1.58
+Campaign on Env. topic 0.47 0 0 1 0.50
+Campaign on Soc. topic 0.53 0 1 1 0.50
+  • Note: This table reports summary statistics and robustness checks across four panels.
+  • Panel (a). Summary statistics on NGO campaigns (Sigwatch); AGM dates (ISS).
+  • Panel (b). Target-firm statistics: accounting (Compustat), ESG (MSCI), news counts (RavenPack).
+4 Empirical Strategy
+Our empirical strategy exploits the timing of NGO campaigns relative to AGMs, leveraging the fact that AGM dates are largely exogenous and stable within firms over time due to SEC regulations.
+Rule 14a-8 requires shareholder proposals to be filed at least 120 days before the proxy statement, effectively anchoring AGM dates across years (glac2014influence; fioretti2024shared).
+Figure 2 illustrates this stability by comparing each firm’s AGM month across consecutive years.
+Most observations lie along the diagonal, indicating strong persistence of the AGM month.
+This predictability is central to our identification strategy.
+Because AGM dates are fixed well in advance and rarely change, NGOs face an exogenous schedule of heightened public attention: they can either react immediately to external events or strategically align their campaigns with these predetermined meetings, when firms are most exposed to shareholder and media scrutiny.
+In this section, we exploit this feature to study how campaigning around AGMs affects NGOs’ visibility (Section 5), stakeholder reactions (Section 5.1), and firms’ subsequent outcomes (Section 6.1).
+Figure 2: Persistence of AGM month over time
+Note: The figure plots the percentage of AGMs held in each month (y-axis) given the month of the previous year’s AGM (x-axis).
+Larger and darker circles indicate a higher proportion of meetings.
+Data cover 38,471 meetings of 5,745 U.S.
+companies from 2010–2020.
+4.1 Target Selection
+A potential concern is that NGOs might select target firms strategically to maximize either visibility or credibility rather than choosing campaign timing exogenously.
+For instance, NGOs could preferentially target firms that are already more visible, frequently covered in the media, or more responsive to shareholder pressure.
+Table 2 addresses this concern by testing for selection on observables o
+Panels (a) and (b) examine whether NGOs with prior AGM campaigns systematically target different types of firms.
+Panel (a) compares average firm characteristics across NGOs with and without previous AGM campaigns.
+In addition, Appendix Figure B1 regresses an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the campaign’s cause, controlling for year fixed effects and clustering standard errors at the firm level.
+[^6] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), environmental and social scores (Env.
+score, Soc.
+score), campaign sentiment, and an indicator for whether the campaign concerns environmental topics (Env.
+campaign).
+None of these variables significantly predicts AGM targeting.
+A one–standard-deviation increase in any firm characteristic shifts the likelihood of an AGM campaign by less than 5 percent of a standard deviation, indicating little selection on observables regarding on whether target firms of NGOs targeting AGM campaigns systematically differ from other NGOs.
+Panel (b) further regresses target characteristics on the number of an NGO’s past campaigns (the results are similar for past AGM campaigns).
+We include standard balance-sheet variables (market capitalization, total assets, and sales) as well as media and ESG indicators (news coverage, environmental and social scores, and topic dummies).
+None of these coefficients is economically and statistically significant, except for a mild increase in environmental campaigns over time, consistent with a post–Paris Agreement shift in NGO focus toward environmental issues.
+Panels (c) and (d) turn to NGO-level selection.
+Panel (c) focuses on 541 U.S.
+NGOs with balance-sheet data and regresses firm characteristics on NGO size, proxied by deciles of total assets, while Panel (d) correlates NGO financial ratios – assets, donations-to-assets, and expenses-to-assets – with the number of past campaigns.
+In both analyses, coefficients are small and statistically insignificant, showing no systematic link between NGO scale or financial capacity and target-firm characteristics.
+Taken together, these results indicate that neither NGO nor firm characteristics predict AGM-timed activism.
+The evidence supports our interpretation that campaign timing reflects strategic decisions by NGOs rather than endogenous target selection.
+However, while our analysis rules out selection on observables, unobserved factors may still affect campaign timing – a possibility we address in Section 8, where we exploit quasi-exogenous variation in campaign dates.
+Table 2: Selection on Observables
+Panel a: Avg. target characteristics’ based on whether the campaigning NGO campaigned on AGM dates in the past
+NGO w/ Past AGM Campaigns NGO w/o Past AGM Campaigns Test Difference
+Mean S.E. Mean S.E. p-value
+Market Cap (decile) 9.23 0.10 9.07 0.04 0.13
+Total Assets (decile) 9.26 0.08 9.23 0.04 0.72
+Sales (decile) 9.31 0.09 9.14 0.04 0.07
+Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
+Environmental score 5.64 0.12 5.72 0.06 0.58
+Social score 4.44 0.09 4.45 0.04 0.89
+Campaign on Env. topic 0.46 0.03 0.5 0.01 0.15
+Campaign on Soc. topic 0.54 0.03 0.49 0.01 0.12
+Panel b: Target’s characteristics based on NGO’s past campaigns regressing $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{\# Past Campaign}_{g(n)y}+\varepsilon_{ny}$
+$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
+Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
+Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
+Total Assets Decile 0.003 (0.004) -0.002 (0.012) 1,338
+Sales Decile 0.001 (0.004) -0.002 (0.03) 1,337
+Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
+Environmental score 0.004 (0.009) 0.041 (0.041) 1,338
+Social score -0.0006 (0.006) 0.019 (0.032) 1,338
+Campaign on Env. topic 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
+Campaign on Soc. topic -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
+Panel c: Target’s characteristics based on NGO size regressing $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{Size Decile}_{g(n)y}+\varepsilon_{ny}$
+$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
+Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
+Market Cap Decile 0.015 (0.016) -0.006 (0.022) 541
+Total Assets Decile 0.030 (0.023) 0.010 (0.025) 541
+Sales Decile -0.008 (0.019) -0.002 (0.018) 541
+Yearly Number of News -0.020 (0.036) 0.022 (0.028) 508
+Environmental score 0.061 (0.068) -0.501∗∗ (0.189) 521
+Social score 0.016 (0.039) 0.042 (0.120) 521
+Campaign on Env. topic 0.026∗∗ (0.012) -0.006 (0.036) 541
+Campaign on Soc. topic -0.024∗∗ (0.012) 0.016 (0.037) 541
+Panel d: NGO characteristics based on past campaigns regressing $z_{gy}=\beta_{0}+\beta_{1}\cdot\textit{\# Past Campaign}_{gy}+\varepsilon_{gy}$
+$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
+Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
+Total Assets (log) 0.046∗∗∗ (0.013) 0.003 (0.005) 1,079
+Donations to assets -0.023 (0.025) -0.008 (0.016) 1,079
+Expenses to assets -0.027 (0.025) -0.006 (0.015) 1,079
+  • Note: This table reports summary statistics and robustness checks across four panels.
+  • Panel (a).
+No-selection-on-observables test comparing targets where the NGO had previously campaigned on at least one of its targets’ AGM dates.
+Market Cap, Total Assets, and Sales are deciles using all U.S.
+Compustat firms in the campaign year.
+[…]
+[…]
+  • Panel (d).
+Change in NGO characteristics as an NGO campaigns multiple times.
+Each row reports $\beta_{1}$ from the panel-header regression of a dependent variable (e.g., Total assets (log)) on the NGO’s number of past campaigns up to the previous year.
+Columns (3)–(4) include NGO and Year fixed effects.
+Standard errors clustered by NGO.
+5 Campaign Timing and Exposure
+The model predicts that campaigns timed on AGM dates generate a larger visibility shock than early campaigns – formally, that $v(\mathrm{AGM})>v(\mathrm{early})$ in the awareness law of motion (1).
+Because neither the latent visibility inputs $v(\cdot)$ nor the awareness stock $A_{t}$ are observed directly, we test this implication using observable proxies that capture different components of real-world awareness formation.
+Google search activity for the NGO reflects demand-based attention, measuring how many individuals actively seek information about the organization.
+Negative news coverage of the target firm reflects a supply-side media response, capturing how much information is produced and disseminated about the campaign.
+While the model remains agnostic about the microfoundations of awareness, it implies that AGM-timed campaigns should generate larger visibility responses in these variables than campaigns launched at other times.
+To test this, we estimate the following stacked event-study specification:
+$\text{Awareness}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$ (10)
+$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)+\alpha_{c,y}+\alpha_{m,y}+u_{g,m,y}.$
+Because NGOs campaign multiple times, we follow baker2022much and stack all campaign events.
 Each event is denoted $c(i,n,g,y,m)$, where campaign $c(\cdot)$ is identified by cause $i$ (e.g., pollution), NGO $g$ (e.g., Greenpeace), target firm $n$ (e.g., Shell), calendar year $y$, and month $m$.
-We then analyze variation in the Google Trends score for NGO $g$ within a five‐month window around the start of $g$’s campaign; this window may spill into years y1y-1 or $y+1$ if the campaign occurs early or late in year $y$.
-The main explanatory variables are event-time dummies, $\mathbf{1}_{\{M_{c,m,y}=\tau\}}$, equal to one if month $m$ is exactly $\tau$ months from campaign $c$, with $\tau\in[-5,5]$.
-The omitted category is the month immediately preceding the campaign ($\tau=-1$).
-These indicators are interacted with $\text{AGM Campaign}_{c}$, a dummy equal to one if campaign $c$ was launched on the target firm’s AGM date.
-Thus $\beta^{AGM}_{\tau}$ estimates the increase in visibility for NGO campaigning on an AGM date, over the direct effect of campaigning at any date, $\beta^{Non-AGM}$.
-The remaining terms include fixed effects at the NGO–firm–year level ($\alpha_{c,y}$), which absorb all variation across NGO-firm pairs over years, as well as month-by-year fixed effects ($\alpha_{m,y}$).
-Our main coefficients of interest ($\beta^{AGM}_{\tau}$) are therefore identified from within-pair variation across months.
+The event-time indicator $\mathbf{1}\{M_{c,m,y}=\tau\}$ equals one if month $m$ is exactly $\tau$ months from the campaign, with $\tau=-1$ serving as the reference period.
+The coefficients $\beta_{\tau}^{AGM}$ capture the incremental visibility response to AGM-timed campaigns relative to the general effect of campaigning captured by $\beta_{\tau}^{Non\text{-}AGM}$.
+NGO-firm-year fixed effects ($\alpha_{c,y}$) absorb all time-invariant characteristics of each NGO-target pair within a year, while month-year fixed effects ($\alpha_{m,y}$) absorb seasonality and aggregate monthly shocks.
 Standard errors are clustered at the NGO level.
-Under the hypothesis that AGM campaigns affect NGO visibility differently depending on their reputation, we estimate (1) separately for an NGO’s first few campaigns, when its recognition is still limited, and for later ones.
-Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and subsequent campaigns (red triangles).
-The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months.
-For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established.
-These results are estimated within NGOs: early in their lifecycle, NGOs benefit more from AGM timing to build visibility, but the returns to this tactic decline as their public recognition grows.
+The model implies that the impact of visibility shocks on awareness depends on an NGO’s position in the $(A_{t},\pi_{t})$ state space.
+To capture this heterogeneity, we estimate (10) separately for younger and more mature NGOs.
+NGOs with fewer than four total campaigns are classified as “young,” while those with four or more campaigns are treated as “mature,” reflecting higher accumulated awareness and reputation.
+[^7] Appendix Figure B2 replicates the analysis on the full sample.
+Comparing its estimates with those in Figure 3 (discussed below) shows that the increases in Google Trends scores and in negative news coverage are largely driven by NGOs in their first few campaigns (i.e., “young” NGOs).
+Panel (a) of Figure 3 reports the estimated $\beta^{AGM}_{\tau}$ using Google Trends as a proxy for demand-based attention.
+AGM campaigns by young NGOs generate a sizeable rise in search activity, though the estimates are imprecise.
+For mature NGOs, the AGM-related coefficients are small and statistically indistinguishable from zero, consistent with saturation of awareness: once $A_{t}$ is high, additional visibility shocks produce little incremental response.
+[…]
 Figure 3: Increased NGO visibility around an AGM campaign
-(a) The first AGM campaigns bring visibility
-(b) No evidence of Selection
-Note: Panel (a) plots the estimated $\beta^{AGM}$ coefficients from (1).
-The dependent variable is the monthly Google Trends score for the NGO.
+(a) Campaign timing and campaigning NGO’s Google Trends scores
+(b) Campaign timing and number of target firms’ negative news (standardized)
+Note: Each panel plots the estimated $\beta^{AGM}$ coefficients from (10).
+In Panel (a), the dependent variable is the monthly Google Trends score for the NGO.
+In Panel (b), the dependent variable is the (normalized) number of monthly negative media coverage received by a target firm.
 Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns.
-Standard errors are clustered at the campaign level.
-Panel (b) reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects.
-Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
-We show that these results are robust to addressing three main concerns.
+Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
+The contrast between early and later campaigns also sheds light on the persistence of awareness.
+If awareness evolves according to $A_{t+1}=\rho A_{t}+(1-\rho)v(d_{t})$, then the first few campaigns produce the largest increases in $A_{t}$ unless $\rho$ is very close to one.
+The empirical pattern – strong AGM effects for young NGOs and no detectable response for mature NGOs – accordingly points to an awareness process that converges relatively quickly.
+[^8] To see this formally, suppose an NGO always chooses $d_{t}=\mathrm{AGM}$ and normalize $v(\mathrm{AGM})=1$.
+Then awareness follows $A_{t+1}=\rho A_{t}+(1-\rho)$, which solves to $A_{t}=1-\rho^{\,t}$.
+Awareness therefore converges quickly when $\rho$ is moderate: for example, with $\rho=0.5$, the first three campaigns raise $A_{t}$ to $1-0.5^{3}=0.875$, leaving little room for additional campaigns to generate detectable increases in visibility.
+This illustrates why strong AGM effects appear for early campaigns but not for later ones.
+Given that the average NGO conducts 3.2 campaigns, treating NGOs with four or more campaigns as “mature” is thus a reasonable way to proxy for having reached a high-awareness region of the state space.
+Note that the estimates are obtained within NGOs: the sample split is not based on the total number of campaigns between 2010 and 2020, but on each NGO’s early versus later campaigns: that is, all NGOs appearing in the red regressions also appear in the blue regressions.
+Thus, early in their lifecycle, NGOs benefit more from AGM timing to build visibility, but the returns to this tactic decline as their public recognition grows.
+Robustness
+These results are robust to addressing three main concerns.
 First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months.
@@
 Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other.
-To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year.
-Both approaches yield similar results, with substantially larger gains for early campaigns.
-Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility—for example, by focusing on firms that are more vulnerable to pressure.
-Panel (b) of Figure 2 addresses this concern by regressing an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the cause of the campaign, with year fixed effects and firm-clustered standard errors.
-[^3] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), the firm’s Environmental and Social scores (Env.
-score, Soc.
-score), how negative the campaign is towards the firm (Campaign sentiment), and an indicator for the campaign cause being related to an environmental topic (Env.
-campaign).
-Note that no coefficient is significantly different from zero.
-We find that a one standard deviation increase in any of these variables moves the likelihood of an AGM campaign by less than 5% of a standard deviation, indicating little selection on observables.
-Further evidence on potential selection based on the number of past NGO campaigns is presented in Table A1.
-Panel (c) compares firms targeted by NGOs that had previously campaigned on a target’s AGM date with those that had not, finding no significant differences across the same variables analyzed in Panel (b) of Figure 2.
-This suggests that NGOs do not appear to select targets based on their past campaigning strategies.
-Panel (d) corroborates this result by regressing firm characteristics on the number of past NGO campaigns and showing no systematic relationship once fixed effects are included.
-The only consistent pattern is that NGOs increasingly focus on environmental rather than social issues over time, reflecting their growing salience.
-Thus, the evidence presented so far points to:
-Result 1.
-Visibility gains from AGM‐day campaigns are greatest during an NGO’s initial campaigns.
-For clarity, each of the next sections highlights its main findings in a similar manner.
+To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of sun2021estimating, which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year.
+Both approaches yield similar results, with substantially larger gains for AGM campaigns by younger NGOs.
+Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility.
+Section 4.1 examines this concern in detail.
+There we show that target-firm characteristics are unrelated to an NGO’s past campaigning behavior and that NGO size and financial capacity similarly fail to predict which firms are targeted.
+The only systematic pattern is a gradual shift toward environmental campaigns, consistent with broader post–Paris Agreement trends.
+While these findings rule out selection on observables, selection on unobservables remains possible.
+To address this, Section 8 exploits quasi-exogenous variation in the timing of firms’ inclusion in public shaming lists, providing an external validation of our main identification strategy.
+Next, we turn to the reaction of stakeholders to NGOs.
 5.1 Stakeholders’ Reactions
-We next examine how campaign timing affects stakeholders and firms’ ESG outcomes.
+We next examine how key stakeholders respond to NGO campaigns.
+In the model, shareholders react by filing proposals when awareness and credibility increase (Section 5.1.1), while donors respond through warm-glow motives that translate into higher contributions (Section 5.1.2).
+In addition, although consumer behavior lies outside the model, we explore whether campaign timing affects target firms’ revenues as a proxy for potential consumer backlash (Section 5.1.3).
 5.1.1 Shareholders
-Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022b).
-Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest.
-We define as “related” a shareholder proposal on the same cause of the campaign in the previous year.
-Appendix D describes the matching procedure.
-For each campaign in our sample, we examine the probability that the next AGM includes a related shareholder proposal, depending on the campaign’s timing.
-To this end, we recalendarize time so that month 0 corresponds to the AGM month for each firm.
-As discussed in Section 3, AGM months are stable over time; thus, this recentering does not alter the distribution of months across years or risk double counting.
-[^4] All subsequent analyses use the same recalendarization.
-Months –5 and +5 are defined as 45-day intervals to ensure the recalendarized year covers the entire calendar.
-We estimate the following regression:
+We begin with shareholders.
+We test whether they are more likely to submit proposals on the same cause when an NGO campaigned around the firm’s AGM date in the previous year, as suggested by the filing rule in (6).
+A proposal is defined as “related” when it addresses the same cause as the NGO campaign.
+As detailed in Appendix D, we identify related proposals by matching each campaign topic in the Sigwatch data to the topic of an AGM proposal at the same target firm in the following year, using ISS topic definitions.
+Whenever such a match occurs, we classify the shareholder proposal as being on the same topic as the earlier NGO campaign.
+To examine this, we recalendarize time so that month 0 corresponds to each firm’s AGM month, taking advantage of the stability of AGM dates over time.
+This creates a common event-time scale on which to compare campaigns launched near versus far from the AGM.
+We then estimate the following regression:
 $\text{Proposal on Same Cause}_{i,n,y}$ $=\sumop\displaylimits^{\tau=-4}_{\begin{subarray}{c}\tau=-15\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM${}_{c(i,n,g,y,m)}$}=-\tau\}}\right)$ (11)
@@
 For each campaign $c(i,n,g,y,m)$ launched by NGO $g$ against firm $n$ on cause $i$ in year $y$ and month $m$, we analyze the likelihood that the next AGM of firm $n$ includes a shareholder proposal on the same cause.
-Formally, $\text{Proposal on Same Cause}_{i,n,y}$ equals one if such a proposal is observed within 4 to 15 months after campaign $c(\cdot)$, and zero otherwise, with timing measured by the indicator $\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=-\tau\}}$.
-Since SEC 14a8(e)(2) rule prevents shareholders from including proposals in the proxy statement within 120 days of its release, we cap the analysis at four months before the previous year’s AGM and set the reference month to $\tau=-4$ because of SEC rules.
-Although firm-by-time fixed effects are not possible in this setting as they would absorb the variation used to estimate $\beta_{\tau}$, this analysis purges all the variation from cause-by-industry-by-NGO-by-month-by-year through $\alpha_{i,s(n),g,m,y}$ and firm invariant characteristics through $\gamma_{n}$, which we view as the main confounding factor since issue salience varies systematically across firms, industries, causes, and NGOs over time.
+Formally, $\text{Proposal on Same Cause}_{i,n,y}$ equals one if such a proposal is observed within 4 to 15 months after campaign $c(\cdot)$, and zero otherwise, with timing measured by the indicator $\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=-\tau\}}$, which is one if campaign $c(\cdot)$ occurs $\tau$ months from the the target’s AGM date, and zero otherwise.
+Since SEC 14a8(e)(2) rule prevents shareholders from including proposals in the proxy statement within 120 days of its release, we cap the analysis at four months before the previous year’s AGM and set the reference month to $\tau=-4$.
+Although firm-by-time fixed effects are not possible in this setting as they would absorb the variation used to estimate $\beta_{\tau}$, this analysis purges all the variation from cause-by-industry-by-NGO-by-month-by-year through $\alpha_{i,s(n),g,m,y}$ and firm invariant characteristics through $\gamma_{n}$, which we view as the main confounding factor since issue salience varies systematically across industries, causes, and NGOs over time.
 We cluster standard errors by firm.
-Figure B3 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue.
-The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline.
-This timing suggests that campaigns around an AGM may spur proposals at the next one, possibly due to heightened media attention.
-Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome.
-If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partners to immediately gain broad support on the proposal, without having to wait a year to advocate for change at the next AGM.
+Panel (a) of Figure 4 compares an NGO’s first three campaigns (blue dots) with its later campaigns (red triangles).
+AGM-timed campaigns (identified at $\beta_{-12}$) by more established NGOs increase the probability of a related shareholder proposal by roughly two percentage points, whereas AGM campaigns by younger NGOs generate no significant response.
+This pattern is consistent with the success function $S(\cdot)$ in (6): for more mature NGOs, beliefs $\pi_{t}$ are higher, so the expected success probability $\pi_{t}S^{C}+(1-\pi_{t})S^{W}$ lies closer to the credible-type benchmark $S^{C}$, making shareholders more responsive to visibility shocks.
+Although fully credible NGOs would eventually shift entirely to early campaigns in the model, the mature NGOs in our data need not have reached that limiting region.
+AGM-timed campaigns by higher $\pi$ NGOs are therefore consistent with NGOs still transitioning toward higher credibility and awareness, for whom AGM visibility remains sufficiently informative to affect filing decisions.
+In contrast, the youngest NGOs are not credible enough to push shareholder to submit proposals.
+[^9] Appendix Figure 3(a) replicates the analysis on the full sample, suggesting that more “mature” NGOs drive the effect we observe in the full sample.
+Panel (b) focuses on the subset of U.S.
+NGOs for which we observe balance-sheet data and incorporation year, allowing us to compare younger and more mature organizations based on age at incorporation.
+We classify as “young” (blue dots) those founded after 2000, corresponding to NGOs younger than ten years at the start of the sample.
+The pattern is similar to Panel (a): a clear spike in filing rates appears at t=12t=-12 for mature NGOs (red triangles), whereas effects for younger NGOs are smaller and imprecisely estimated.
+Note that these NGOs are generally larger than those in Panel (a), which explains why we see some actions around the AGM also for younger NGOs.
+Figure 4: NGO campaigns and shareholder proposals
+(a) By past experience
+(b) By NGO age
+Notes: Note: These figures plot the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead.
+The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects.
+Panel (a) regression is run on the full sample (1,227 NGO campaigns), while the one in Panel (b) is based on U.S.
+NGOs with balance sheet data (541 campaigns).
+Error bars shows 95% confidence intervals using standard errors that are clustered by firm.
+5.1.2 Donors
+Focusing on the smaller sample of U.S.
+NGOs for which balance-sheet data are available, we examine whether campaigns timed around firms’ AGMs translate into higher donations for the campaigning NGO, and whether this response differs across NGO maturity.
+Because donations are observed only annually, we cannot align them precisely with the timing of individual campaigns, and many NGOs conduct multiple campaigns within the same year.
+Stacking event windows would therefore generate overlapping observations without adding variation on the left-hand side.
+Instead, we estimate a distributed-lag specification that relates the timing of campaigns relative to AGMs to annual financial outcomes:
+$\frac{\text{Donations}_{g,y}}{\text{Assets}_{g,y-1}}=\sumop\displaylimits_{\tau=-4}^{4}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=\tau\}}\right)+\alpha_{g}+\alpha_{y}+u_{g,y},$ (12)
+where “Distance to AGM” measures how NGO $g$’s campaigns in year $y$ are distributed relative to the target firms’ AGMs.
+NGO fixed effects ($\alpha_{g}$) absorb persistent differences in size or donor base, while year fixed effects ($\alpha_{y}$) capture aggregate shocks.
+[^10] We treat total reported revenues as donations.
+Panel (a) of Figure 5 shows that young NGOs (blue dots) experience a clear increase in donations when campaigning on the AGM date, while older NGOs (red triangles) show no meaningful response.
+Panel (b) presents similar patterns for return on assets, with sharp AGM effects only for young NGOs.
+Because ROA nets out expenditures, these findings suggest that AGM-day campaigns are relatively inexpensive yet effective in generating additional funding for less established NGOs.
+These donation patterns should not be interpreted as estimates of the model’s warm-glow term $\alpha v(d_{t})$.
+Instead, they provide reduced-form evidence on how the exposure created by campaign timing translates into financial support.
+As shown in Figure 3, AGM campaigns generate substantially larger visibility shocks for young NGOs, making it natural that only these organizations exhibit a measurable donation response, while the same exposure increment has little effect for mature NGOs.
+[^11] Appendix Figure B4 finds a similar AGM-day effect on the full sample, highlighting the important need for visibility of “younger” NGOs.
+Figure 5: NGO campaign timing and donations
+(a) Donations
+(b) Operating Margin
+Notes: The figure plots the estimated ${\beta}_{\tau}$ from (12).
+Panel (a) focuses on donations over (lagged) assets, while Panel (b) focuses on donations minus expenditures over (lagged) assets.
+Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
 5.1.3 Consumers
-Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
-$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n,y}+\alpha_{s(n),q}+u_{n,q},$ (3)
-where $q$ indicates a quarter and $\alpha_{n,y}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm-by-year level or the industry-by-quarter-by-year level, respectively.
-The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample.
+We next examine whether NGO campaigns trigger a consumer response, even though such behavior lies outside the model.
+Using quarterly Compustat data, we study how the target firm’s log of sales varies with the timing of NGO campaigns relative to the firm’s AGM in a specification similar to (10):
+$\log\left(\text{Sales}_{n,q,y}\right)$ $=\sumop\displaylimits^{\tau=4}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,q,y}=\tau\}}\right)+\sumop\displaylimits^{\tau=4}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,q,y}=\tau\}}\right)$ (13)
+$+\log\left(\text{Assets}_{n,q,y}\right)+\alpha_{c,y}+\alpha_{m,y}+u_{n,q,y}.$
+where $q$ denotes the quarter, $\alpha_{c,y}$ are campaign fixed effects, $\alpha_{q}$ are industry-by-quarter fixed effects, and standard errors are clustered at the campaign level.
+Figure 6 reports the estimated $\hat{\beta}_{\tau}$} coefficients, separately for campaigns launched on the AGM date (Panel a) and for all other campaigns (Panel b).
+Across both specifications, we find no meaningful change in sales around NGO campaigns, suggesting that consumers do not systematically adjust their purchasing behavior in response to NGO activism.
+Taken together with earlier results, this evidence indicates that NGO campaigns primarily affect stakeholders other than consumers, especially shareholders and donors, consistent with the model’s focus on visibility and influence within the investor domain.
+Figure 6: NGO campaign timing and sales
+(a) NGO Campaigns on AGM dates
+Notes: The figure plots the estimated ${\beta}_{\tau}$ from (13).
+The reference period is -1 quarter, consistent with the other analyses.
+The number of observations is 239,818 in Panel (a) and 239,925 in Panel (b).
+Error bars shows 95% confidence intervals using standard errors that are clustered by firm.
+Overall, the evidence in this section indicates that AGM timing delivers meaningful visibility gains mainly for younger NGOs.
+We now turn to whether this timing advantage also translates into differences in shareholder proposal outcomes and firm responses.
+6 Campaign Timing and Influence
+The model predicts that early campaigns offer NGOs greater scope to influence contemporaneous AGM outcomes through the term $_{d_{t}}$ in the proposal success equation (4).
+Because most shareholders cast their votes by proxy before the meeting, campaigns launched on the AGM date have limited ability to affect voting outcomes (i.e., $_{AGM_{t}}=0$).
+For example, more than 85% of vote-eligible shares for Microsoft’s 2023 AGM were submitted in advance (microsoft2023), illustrating why visibility at the AGM does not necessarily translate into influence over the vote itself.
+We classify a shareholder proposal as successful if it is withdrawn before the AGM or receives at least 50% + 1 of the votes at the meeting.
+To assess whether NGO campaigns influence the success of related proposals, we estimate the following specification on the full sample of shareholder proposals:
+$\text{Success}_{i,n,m,y}=\sumop\displaylimits_{\tau=-7}^{0}\left(\beta_{\tau}\cdot\text{Campaign}_{i,n,m+\tau,y}\right)+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (14)
+where $\text{Success}_{i,n,m,y}$ is an indicator for the success of proposal $i$ at firm $n$, and $\text{Campaign}_{i,n,m+\tau,y}$ indicates whether an NGO campaign on the same cause occurred $\tau$ months before the AGM.
+Time is recentered so that the AGM month is coded as month 0 for each firm, and the reference period corresponds to months 11 to 7 months before the AGM, which we code as $\tau=-7$ in (14).
+Campaigns in this [11,7][-11,-7] window – roughly three to four quarters before the focal meeting – are more likely to reflect reactions to the previous year’s AGM rather than anticipatory campaigning.
+The regression includes firm-by-year, cause-by-year, and month fixed effects, thereby exploiting within–firm-year and within-cause variation in campaign timing.
 Standard errors are clustered at the firm level.
-Appendix Figure B4 shows the estimated $\hat{\beta}_{\tau}$} coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b).
-We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns.
-[^5] Unreported results show also no difference between early and late campaigns.
-4.2 Does AGM-Targeting Create Positive Change?
-If public shaming effectively influenced firm behavior, the AGM month should deliver the largest improvements in E&S scores, given that it triggers heightened negative news coverage of the target firm (Appendix Figure B2).
-However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year.
-[^6] Acknowledging the limits of ESG scores as measures of social and environmental impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), we use MSCI to mitigate signal risk: its ratings are more time-consistent and widely used by sustainable investors (Berg et al., 2021, 2024).
-By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
-This finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore, the results in this section suggest that:
-Result 2.
-NGO campaigns on AGM dates spur shareholders to place related proposals at future AGMs of the targeted firm, but they produce no contemporaneous improvements in the firm’s E&S scores.
-5 Timing Campaigns to Influence AGM Outcomes
-Early campaigns could benefit an NGO aiming to influence current AGM outcomes.
-Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes.
-For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting, indicating that campaigning around AGM dates may have little impact on shareholders’ votes (Microsoft, 2023).
-We define a proposal as successful if either it is withdrawn by the shareholder before the AGM or it secures at least $50\%+1$ vote at the AGM.
-We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
-$\text{Success}_{i,n,m,y}=\sumop\displaylimits^{0}_{\begin{subarray}{c}\tau=-6\end{subarray}}\beta_{\tau}\cdot\text{Campaign}_{i,n,m+\tau,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
-where $\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+\tau,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $\tau$ months before the AGM.
-The reference period covers the months that are more than six months before the AGM of year yy—i.e., $m\in[-11,-7]$, as campaigns launched during that period are more likely to be in reaction to the previous year’s AGM results.
-We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs.
-The standard errors are clustered by firm.
+[^12] The sample is not split by NGO maturity because matching campaigns and proposals by topic already reduces the number of observations.
 Figure 7: NGOs’ campaigns before the AGM help E&S shareholder proposals succeed
@@
 To understand why campaigns launched six months before the AGM are also somewhat successful, we re-estimate (14) on the subset of proposals that went to a vote, using the percentage of votes in favor as the dependent variable.
-[^7] Since this analysis focuses on voted proposals, we also control for the voting recommendations of management and proxy advisors such as ISS (“For,” “Against,” or “Abstain”).
+[^13] Since this analysis focuses on voted proposals, we also control for the voting recommendations of management and ISS (“For,” “Against,” or “Abstain”).
 Panel (a) of Figure 7 shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points.
@@
 This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm.
-As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
-Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015).
+Taken together, these results indicate that as a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal.
+Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by flammer2015does.
 Given the generally low support of shareholder proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote.
-However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011).
+However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (levit2011nonbinding).
 This reinforces the importance of analyzing all E&S proposals, including those withdrawn as in Panel (a).
-The evidence in this section points to:
-Result 3.
-Pre-AGM campaigns raise shareholder support and proposal success, while AGM-day campaigns have little effect.
+6.1 Campaign Timing and Prosocial Changes
+We next examine whether the timing of NGO campaigns relates to subsequent changes in firms’ environmental and social (E&S) performance.
+The model does not explicitly incorporate firm responses driven by public shaming, yet such a mechanism would imply that visibility generated during the AGM – when exposure is highest – could directly pressure firms, corresponding to $_{\mathrm{AGM}_{t}}>0$.
+By contrast, the model’s influence mechanism operates through stakeholder interactions before the AGM (see equation 3): early campaigns may foster constructive dialogue between shareholders and firms, shaping proposals or prompting managerial adjustments ahead of the meeting as illustrated in the previous section (Figure 7).
+To investigate which mechanism is more consistent with the data, we estimate the following specification:
+[…]
+[…]
+Figure 8: Campaign Timing and Changes in E&S Scores
+Note: The figure plots the estimated coefficients $\beta_{l}$ from equation (15).
+The dependent variable is the yearly change in a firm’s average E&S scores.
+Shaded areas show 95% confidence intervals based on firm-level clustered standard errors.
+Data combine MSCI ESG scores, NGO campaigns from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
+Figure 8 reports the estimated $\beta_{l}$ coefficients from equation (15), distinguishing an NGO’s first three campaigns (blue dots) from its later campaigns (red triangles).
+Only early campaigns by more established NGOs – those launched between six and three months before the AGM – lead to measurable improvements in firms’ E&S performance.
+Following such campaigns, E&S scores rise by about 2% over the subsequent year, whereas campaigns at other times, including during the AGM month, have no statistically or economically significant effect.
+Combined with the results in Figure 7, this pattern suggests that established NGOs can influence firm behavior ahead of the AGM, potentially by shaping shareholder proposals or managerial decisions without requiring a formal vote.
+For younger NGOs, positive effects arise only for campaigns launched close to the AGM, with estimates significant at the 90 percent level.
+Campaigns occurring earlier in the year show no beneficial impact and may even reduce subsequent E&S scores.
+This mirrors the spike in shareholder support observed one month before the AGM in Figure 7: NGOs still building credibility may be perceived as effective only when campaigning in the immediate run-up to the meeting, but lack the standing needed to engage constructively with firms at earlier stages.
+[^15] Appendix Figure B5 replicates the analysis on the full sample, showing that only early campaigns (about 6 months in advance) have a positive impact on ESG scores.
+The coefficients for the other months are close to zero, consistent with them being the average between the red triangles and blue dots in Figure 8, thereby masking important heterogeneity based on NGO maturity.
 7 Exposure vs. Influence
-Sections 4 and 5 outlined a tradeoff: campaigns launched close to an AGM maximize exposure, while those launched earlier maximize influence.
-In this section, we examine how NGOs navigate this tradeoff as their visibility increases.
-We first study whether greater past exposure leads NGOs to shift campaign timing from the AGM date to earlier months.
-We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
-6.1 The Timing of Past and Current Campaigns
+[…]
+7.1 Choosing when to Campaign based on Past Campaigns
 We first examine whether NGOs shift from targeting AGMs to launching earlier campaigns as they campaign more. Specifically, we estimate
-$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}$ (5)
-$\quad+\alpha_{t(\cdot),y}+\alpha_{g,y}+\alpha_{n,y}+u_{c(i,n,g,y,m)}.$
+$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau>1}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}\right)$ (16)
+$\quad+\alpha_{i,y}+\alpha_{g,y}+\alpha_{n,y}+u_{c(i,n,g,y,m)}.$
 where the dependent variable measures the number of months between campaign $c(\cdot)$ and the targeted firm’s AGM, and the key explanatory variable is the sequence number of campaigns by NGO $g$ by year $y$.
-$\alpha_{t(\cdot),y}$, $\alpha_{g,y}$, and $\alpha_{n,y}$ are topic-by-year, NGO-by-year, and firm-by-year fixed effects, respectively.
+$\alpha_{i,y}$, $\alpha_{g,y}$, and $\alpha_{n,y}$ are topic-by-year, NGO-by-year, and firm-by-year fixed effects, respectively.
 Identification thus comes from NGOs launching multiple campaigns in the same year.
@@
 Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7).
-Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO past AGM exposure (Appendix Figures B8 and B9).
-6.2 Visibility and Campaign Timing
-We next use Google Trends scores to proxy for an NGO’s reputation.
-We compute the average search score for NGO $g$ in year $y$ relative to Greenpeace in year y1y-1, discretized into deciles and regress:
-$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Relative Google Score}_{g,y-1}=\tau\!\right\}}\right)$ (6)
-$\quad+\alpha_{t(\cdot)}+\alpha_{y}+u_{c(i,n,g,y,m)},$
-which is similar to model (6) with $\text{Relative Google Score}_{g,y-1}$ as a measure of NGO visibility.
-$\alpha_{t(\cdot)}$ and $\alpha_{y}$ are topic and year fixed effects.
-Panel (b) of Figure 4 shows that only the most visible NGOs—those in the top deciles of relative searches—campaign several months ahead of AGMs, while less visible NGOs continue to concentrate on AGM dates.
+Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor its ESG scores covary with NGO past AGM exposure (Appendix Figures B8 and B9).
+7.2 Choosing When to Campaign Based on NGO Visibility
+We next use Google Trends to proxy for an NGO’s visibility.
+For each NGO $g$, we compute a Relative Google Score equal to its average search score in year y1y-1 divided by the corresponding score for Greenpeace – the most searched NGO in our sample – and discretize this measure into quartiles.
+We then estimate:
+$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=1}^{4}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Relative Google Score Quartile}_{g,y-1}=\tau\right\}}\right)$ (17)
+$\quad+\alpha_{i,y}+\alpha_{g}+\alpha_{n,y}+u_{c(i,n,g,y,m)}$
+which relates campaign timing to lagged NGO visibility.
+Because the key regressor varies only across NGOs and years, we cannot include NGO-by-year fixed effects as in earlier specifications.
+[^17] The Relative Google Score is constructed from the previous year’s Google Trends data, generating variation only across NGOs and years but not within NGO–year cells.
+Panel (b) of Figure 9 shows that only the most visible NGOs – those in the top quartiles of the relative search distribution – campaign several months before AGMs, whereas less visible NGOs continue to concentrate their actions on the AGM date.
 Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
-Result 4.
-NGOs initially target AGM dates to gain visibility. As their exposure increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
 Figure 9: NGO reputation and campaign timing
-(a) Later campaigns pre-date the target’s AGM date
-(b) More visible NGOs campaign earlier
+(a) NGOs tend to campaign before the target’s AGM date as they gain experience
+(b) More visible NGOs campaign before the target’s AGM date
 Note: This figure plots the estimated coefficients from (16).
@@
 7.3 NGO Reputation and AGM Influence
-We now ask whether NGO reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
+We now examine how the evolution of NGO reputation shapes their ability to influence shareholder proposal outcomes, tying together the patterns documented in Sections 7.1 and 7.2.
+In the model, proposal success depends jointly on awareness and credibility through $S(A_{t},\pi_{t},d_{t})$ in (4).
+NGOs that rely heavily on AGM-timed campaigns accumulate visibility but gain little credibility, so repeated AGM activity eventually leads to lower beliefs $\pi_{t}$ and, with it, reduced persuasive power.
+By contrast, NGOs that shift toward earlier, influence-oriented campaigns build credibility over time.
+The model therefore implies that overusing visibility-driven tactics should diminish the effectiveness of AGM campaigns for otherwise credible NGOs, whereas NGOs with stronger reputations should be more effective at shaping proposal outcomes.
+To test this implication, we estimate the following specification:
 $\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{\# Past Campaigns}_{g,y-1}$ (18)
-$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{\# Past Campaigns}_{g,y-1}$
-$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
+$\quad+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{\# Past Campaigns}_{g,y-1}$
+$\quad+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
 where $\text{Success}_{i,n,y}$” indicates whether a proposal is withdrawn or passes a vote, as in equation (14).
@@
 We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
-Table 1: NGO’s strategy and support for related shareholder proposals
+Column (1) of Table 3 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$).
+What matters is the NGO’s past activity on that issue: proposals are more likely to succeed when the NGO has previously campaigned, and a one–standard-deviation increase in past campaigns raises proposal success by about 8%.
+The squared term indicates diminishing marginal effects.
+Table 3: NGO’s strategies and support for related shareholder proposals
 Dependent Variable: Proposal Passes (0/1)
 Model: (1) (2) (3) (4)
-Campaign on Same Topic -0.0003 -0.013 -0.005 -0.001
+Campaign on Same Cause -0.0003 -0.013 -0.005 -0.001
 (0.026) (0.028) (0.025) (0.030)
 # Past Campaigns -0.034
-Campaign on Same Topic $\times$ # Past Campaigns 0.078∗∗∗
+Campaign on Same Cause $\times$ # Past Campaigns 0.078∗∗∗
 # Past Campaigns squared 0.008∗∗
-Campaign on Same Topic $\times$ # Past Campaigns squared -0.015∗∗
+Campaign on Same Cause $\times$ # Past Campaigns squared -0.015∗∗
 # Past Non-AGM Campaigns 0.0008
-Campaign on Same Topic $\times$ # Past Non-AGM Campaigns -0.011
+Campaign on Same Cause $\times$ # Past Non-AGM Campaigns -0.011
 # Past AGM Campaigns 0.005 -0.034
 (0.020) (0.023)
-Campaign on Same Topic $\times$ # Past AGM Campaigns 0.048 0.091∗∗∗
+Campaign on Same Cause $\times$ # Past AGM Campaigns 0.048 0.091∗∗∗
 (0.026) (0.032)
 # Past AGM Campaigns squared 0.010∗∗
-Campaign on Same Topic $\times$ # Past AGM Campaigns squared -0.013∗∗
-Campaign on Same Topic $\times$ Google Search Popularity 0.071
-Campaign on Same Topic $\times$ Google Search Popularity square -0.021
+Campaign on Same Cause $\times$ # Past AGM Campaigns squared -0.013∗∗
+Campaign on Same Cause $\times$ Google Search Popularity 0.071
+Campaign on Same Cause $\times$ Google Search Popularity square -0.021
 Fixed-effects
-CUSIP x Year
-NGO x Year
-Topic x Month-Year
+CUSIP $\times$ Year
+NGO $\times$ Year
+Cause $\times$ Month-Year
 Observations 36,543 35,778 36,543 33,142
@@
   • Note: This table reports the estimated coefficients from (18) in Column 1.
-In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (# Past Non-AGM Campaigns) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), and their interaction with Campaign on Same Topic.
-In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), its square, and their interaction with Campaign on Same Topic.
-In Column 4, we replace # Past Campaigns from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign.
-All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects.
+In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (# Past Non-AGM Campaigns) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), and their interaction with Campaign on Same Cause.
+In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), its square, and their interaction with Campaign on Same Cause.
+In Column 4, we replace # Past Campaigns from model 18 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign.
+All regressions also include firm-by-year, NGO-by-year, Cause-by-month-by-year fixed effects.
 Clustered (firm) standard errors are reported in parentheses.
-Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$).
-What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has already campaigned ($\beta_{3}>0$).
-A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average.
-The regression also includes the squared term of “# $\text{\# Past Campaigns}_{g,y-1}$,” showing decreasing marginal returns of campaigning.
-We unpack the mechanism behind these coefficient estimates by distinguishing between prior AGM campaigns and non-AGM campaigns in Column (2).
-We find that only the former contribute to proposal success.
-In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
-Column (3) tests for diminishing returns by including squared terms.
-The results confirm a concave pattern: additional AGM campaigns continue to improve proposal success, but at a decreasing rate.
-Importantly, the combined effect remains non-negative even for the most active NGOs.
-[^9] The marginal effect of AGM campaigning never turns negative in our sample since no firm launches more than five AGM-date campaigns, so our proxy for reputation does not imply adverse returns.
-We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends.
-Specifically, Column (4) replaces “# $\text{\# Past Campaigns}_{g,y-1}$” by “Google Search Popularityg,y-1,” which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.101010The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects.
-The interaction between this measure and campaigning on the same cause is positive, with a concave shape.
-Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
-These results support the following result:
-Result 5.
-Targeting AGMs builds NGO reputation and increases their ability to raise the support of related shareholder proposals, but with diminishing returns.
-7 Discussion and Conclusion
-Examining NGO campaign timing, we show that not just the decision to intervene but also its timing can materially affect firm behavior and shape the course of future activism.
-Over time, NGOs shift from visibility-driven exposure tactics to those designed to influence corporate strategies, highlighting how the temporal dimension of their actions channels pressure that leads firms to internalize social and environmental externalities.
-We first show that NGOs’ early campaigns often coincide with the target’s AGM date, when media attention is highest, and that such timing significantly increases their visibility (Result 1).
-These visibility gains are concentrated in NGOs’ first few campaigns, indicating that AGM-day actions serve primarily to build name recognition rather than to influence firms directly.
-By contrast, later AGM-day campaigns have no immediate effect on firms’ E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
-Hence, campaigning before the AGM offers direct advantages.
-When an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of success—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3).
-Given the typically low baseline support for such proposals, early NGO interventions can effectively discipline managers who must contend with heightened shareholder pressure (Levit and Malenko, 2011).
-Importantly, all analyses include NGO-by-year and cause-by-year fixed effects, thereby accounting for time-varying confounders at both the NGO level—such as experience or reputation—and the cause level—such as specific news events.
-This ensures that globally renowned NGOs (e.g., PETA) do not drive the results.
-Furthermore, the analyses in Sections 4 and 6 show little evidence of selection on observables, alleviating concerns that NGOs disproportionately target firms that are easier to influence or more likely to generate visibility.
-Our results suggest that NGOs face a trade-off that they navigate by adapting their strategies over time.
-NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
-As a result, having targeted AGMs in the past increases the likelihood that an NGO can help a shareholder proposal pass related to its campaign.
-However, repeatedly targeting AGMs leads to weaker influence over firm behavior (Result 5).
-One interpretation is that NGOs, having already demonstrated commitment to their cause, face weaker incentives to keep signaling through costly AGM-day campaigns.
-In this view, declining returns simply reflect a reduced need for visibility once reputational capital is established.
-Yet the evidence points to a different mechanism.
-Within the same NGO, AGM-day campaigns progressively lose their marginal effect on media attention (Figure 2) and their ability to damage a firm’s reputation (Appendix Figure B10), suggesting that the audience’s response—rather than the NGO’s effort—changes over time.
-As NGOs become recognized and credible actors, stakeholders interpret their campaigns differently: what initially signals toughness later signals redundancy or routine.
-Established NGOs are expected to engage in more constructive dialogue, such as supporting shareholder proposals, rather than resorting to public confrontation.
-Consequently, the same action produces less reputational harm for firms and fewer visibility gains for NGOs.
-The diminishing returns to AGM targeting thus stem from evolving stakeholder expectations and a changing perception of NGO legitimacy, rather than from reduced internal pressure to seek exposure.
-[^11] This interpretation does not rule out signaling motives altogether.
-Early AGM-day campaigns are consistent with “reputation-for-toughness” models, in which actors bear short-term costs to establish credibility that can be leveraged in future interactions (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
-Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause.
-For instance, organizational aging (e.g., older NGOs becoming larger and more complex) or experience in managing AGM-related information is accounted for through NGO-by-year fixed effects.
-These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
-In conclusion, NGOs without an established reputation coordinate stakeholders over time by campaigning on AGM dates, drawing attention to issues that spur future engagement.
-Once reputable, they coordinate stakeholders within the same AGM cycle by supporting existing shareholder proposals and directly shaping firm decisions.
-Thus, NGOs act as a coordination device throughout their lifecycle, with the form of coordination evolving with their perceived toughness and credibility.
-In this way, NGOs’ incentives shape their strategies and, ultimately, the social and environmental impact of firms.
-We also acknowledge that NGOs increasingly engage not only in campaigns but also in collaborations with corporations (Dahan et al., 2010).
-Future research could examine when confrontation shifts to cooperation and how such collaboration reshapes the balance between visibility, credibility, and influence.
+Column (2) separates past AGM and non-AGM campaigns.
+Only AGM campaigns contribute meaningfully: a one–standard-deviation increase in past AGM campaigns raises proposal success by roughly 4.8 percentage points, whereas non-AGM campaigns have no detectable impact.
+Column (3) confirms concavity for AGM campaigns.
+This pattern is consistent with the model’s belief dynamics: AGM campaigns generate visibility but convey limited new information, so repeated AGM activity initially raises beliefs $\pi_{t}$ but eventually adds little, reducing its influence on shareholder decisions.
+Column (4) uses an alternative proxy for reputation – NGO visibility from Google Trends, measured relative to Greenpeace in year y1y\!-\!1.
+The interaction between this visibility measure and campaigning on the same cause is positive but concave, indicating that the effectiveness of AGM campaigns increases with NGO visibility but at a decreasing rate.
+This aligns with the idea that once awareness and credibility are high, additional exposure contributes little to reputation formation.
+Different measures of reputation
+We next examine whether broader measures of NGO scale show similar patterns using data on U.S.
+NGOs.
+Table 4 replaces Google visibility with lagged assets (Column 1), lagged donations (Column 2), and lagged expenditures (Column 3).
+We view these variables as manifestations of an NGO’s reputation: financially larger organizations are more established, have accumulated greater public visibility, and possess the operational capacity to sustain campaigns, all of which are consistent with higher positions in the $(A_{t},\pi_{t})$ state space.
+[^18] A large literature documents that NGO reputation is closely tied to organizational scale, resources, and operational capacity.
+Larger and better-resourced NGOs are viewed as more credible, legitimate, and effective by external stakeholders (e.g., baur2011moral; cooley2002ngo; mitchell2017reputations).
+Across all specifications, proposal success initially rises with NGO size but exhibits diminishing returns.
+This suggests that while more established NGOs are better positioned to influence proposal outcomes, additional scale eventually contributes little new reputational information and may attenuate the marginal effect of further AGM-driven exposure.
+Table 4: NGO’s strategies and support for related shareholder proposals
+Dependent Variable: Proposal Passes (0/1)
+Model: (1) (2) (3)
+Campaign on Same Cause -0.299∗∗ -0.167 -0.208
+(0.145) (0.119) (0.143)
+Campaign on Same Cause $\times$ NGO Size 0.132∗∗∗ 0.073∗∗ 0.089∗∗
+(0.049) (0.032) (0.040)
+Campaign on Same Cause $\times$ Squared NGO Size -0.011∗∗∗ -0.006∗∗ -0.007∗∗
+(0.004) (0.003) (0.003)
+Size is defined by Total Assets Donations Expenses
+Fixed-effects
+CUSIP $\times$ Year
 […]
-Appendix A Omitted Tables
-Table A1: Summary Statistics
-Panel a: Variables varying by NGO
-Mean 25% 50% 75% 95%
-Number of campaigns by NGO 3.20 1 1 3 13
-     – Campaigns on Env. topic 1.51 0 1 1 5
-     – Campaigns on Soc. topic 1.70 0 1 1 6
-     – On AGM date 0.24 0 0 0 1
-Campaign sentiment -1.72 -2 -2 -1 -1
-Panel b: Variables varying by target firm
-Mean 25% 50% 75% SD
-Market Cap (USD bn) 175.66 21.71 75.05 209.47 263.66
-Total Assets (USD bn) 98.81 21.80 59.27 138.68 124.36
-Yearly Sales (USD bn) 37.26 6.85 17.35 42.41 150.58
-Yearly Number of News 1,183 328 616 1,046 1,695.25
-Environmental score 5.66 4 5.58 7.3 2.22
-Social score 4.44 3.30 4.70 5.60 1.58
-Campaign on Env. topic 0.47 0 0 1 0.50
-Campaign on Soc. topic 0.53 0 1 1 0.50
-Panel c: Selection—Avg. target characteristics’ based on whether the campaigning NGO campaigned on AGM dates in the past
-NGO w/ Past AGM Campaigns NGO w/o Past AGM Campaigns Test Difference
-Mean S.E. Mean S.E. p-value
-Market Cap (decile) 9.23 0.10 9.07 0.04 0.13
-Total Assets (decile) 9.26 0.08 9.23 0.04 0.72
-Sales (decile) 9.31 0.09 9.14 0.04 0.07
-Yearly Number of News (lag) 1,132.37 114.74 1,368.67 47.67 0.06
-Environmental score 5.64 0.12 5.72 0.06 0.58
-Social score 4.44 0.09 4.45 0.04 0.89
-Campaign on Env. topic 0.46 0.03 0.5 0.01 0.15
-Campaign on Soc. topic 0.54 0.03 0.49 0.01 0.12
-Panel d: Selection—Target’s characteristics based on NGO’s past campaigns regressing $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{\# Past Campaign}_{g(n)y}+\varepsilon_{ny}$
-$\hat{\beta}_{1}$ (SE) w/o FEs $\hat{\beta}_{1}$ (SE) w/ FEs N
-Dependent Variable ($z_{ny}$): (1) (2) (3) (4)
-Market Cap Decile -0.002 (0.003) 0.003 (0.016) 1,338
-Total Assets Decile 0.003 (0.004) -0.002 (0.012) 1,338
-Sales Decile 0.001 (0.004) -0.002 (0.03) 1,337
-Yearly Number of News 0.996 (12.4) -4.12 (16.5) 1,323
-Environmental score 0.004 (0.009) 0.041 (0.041) 1,338
-Social score -0.0006 (0.006) 0.019 (0.032) 1,338
-Campaign on Env. topic 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
-Campaign on Soc. topic -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
-  • Note: This table reports summary statistics and robustness checks across four panels.
-  • Panel (a). Summary statistics on NGO campaigns (Sigwatch); AGM dates (ISS).
-  • Panel (b). Target-firm statistics: accounting (Compustat), ESG (MSCI), news counts (RavenPack).
-  • Panel (c).
-No-selection-on-observables test comparing targets where the NGO had previously campaigned on at least one of its targets’ AGM dates.
-Market Cap, Total Assets, and Sales are deciles using all U.S.
-Compustat firms in the campaign quarter.
-Deciles avoid mechanical correlation with # Past NGO Campaigns, which grows over time.
 […]
+Observations 12,127 12,367 12,127
+R2 0.59598 0.59441 0.59474
+*–p<0.1<0.1; **–p<0.05<0.05; ***–p<0.01<0.01.
+  • Note: This table reports the estimated coefficients from (18) where, instead of Past Campaign, we use NGO size as measured in terms of its Total Assets (Column 1), Donations (Column 2), and Expenses (Column 3).
+The dataset includes only campaigns run by U.S.
+NGOs for which we observe their balance sheet data.
+These variables are lagged.
+All regressions also include firm-by-year, NGO-by-year, Cause-by-month-by-year fixed effects.
+Clustered (firm) standard errors are reported in parentheses.
+7.4 Summary of Results
+These findings provide clear evidence for the model’s dynamic trade-off between exposure and influence.
+NGOs initially rely on AGM-timed campaigns to build visibility, but as they accumulate reputation and credibility, they shift toward earlier, influence-oriented actions.
+Consistent with the model, repeated AGM campaigning eventually generates diminishing reputational returns: past exposure raises the effectiveness of current campaigns only up to a point, after which additional AGM visibility adds little new information and may even weaken influence.
+Across all empirical measures of reputation – past AGM activity, Google search visibility, and organizational size – we observe the same concave pattern, consistent with credibility saturation and declining informational value of purely visible actions.
+8 Accounting for Selection
+A remaining concern is that NGOs may choose campaign dates strategically, potentially confounding the timing effects documented above.
+To address this, we turn to a setting in which both campaign timing and target identities are anchored to an external event and thus provides quasi-exogenous variation.
+Institutional background
+[…]
+Data
+We collect daily stock prices (form LSEG), AGM dates, and data on environmental and social (E&S) shareholder proposals (from ISS) for 97 listed firms included in the Index between 2016 and 2019.191919The first report included 40 brands globally, and its coverage has increased to 200 by 2020, located in 16 different countries.
+We focus on the firms that have been listed between 2016 and 2019.
+Stock market reactions
+For each firm, we download daily returns on its national stock index (e.g., FTSE 100 for U.K.
+companies, CAC40 for French companies) and compute its daily excess return over the index.
+We then cumulate those returns and plot the results in Figure 10, depending on the report’s publication timing relative to the company’s AGM (Panel a), and depending on the company’s initial FTI score (below or above the median that year, in Panel b).
+The patterns are consistent with the model’s visibility channel: when firms are most exposed to scrutiny, credible activism amplifies the financial costs from a deteriorating reputation (kolbel2017media).
+Figure 10: Stock returns around inclusion
+(a) Stock returns depending on timing
+(b) Stock returns depending on first score
+Note: This figure plots the Cumulative Excess Returns of stocks around the publication of the first Fashion Transparency Index covering the company.
+The excess returns are computed as the difference between the stock’s daily return and the return of its national stock index.
+The sample covers 97 listed stocks between 2016 and 2019.
+Shareholder outcomes
+Figure 11 shows that proposals on the same issues as the Index are 7 percentage points more likely to succeed when the Index is published before the AGM and receive about 8 percentage points more support if voted.
+These magnitudes closely match those estimated in Section 7.3, reinforcing that early, credibility-driven actions enhance influence through coordinated shareholder responses.
+Figure 11: Index publication and support for E&S shareholder proposals
+Note: This figure reports the average success of shareholder proposals at firms targeted by the Fasion Transparency Index, depending on whether the report is published before (blue bars) or after (gray bars) the firm’s AGM.
+The sample covers 40 E&S shareholder proposals, filed at 23 firms between 2016 and 2019.
+Shareholder proposals
+Finally, Figure 12 replicates the timing tests from earlier sections.
+Publication in the month leading up to the AGM raises the probability of a related proposal at the next AGM, whereas publication two months before the AGM – analogous to early campaigning – leads to roughly 10-point improvements in firms’ FTI scores over the following year.
+Figure 12: Publication timing and firm outcomes
+(a) Publication timing and related shareholder proposals in the next AGM
+(b) Publication timing and FTI score improvement
+Note: This Figure plots the estimated coefficients from:
+$y_{n,y}=\sumop\displaylimits_{\tau=-5}^{2}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM = }\tau\}}\cdot\text{Treated}_{n,y}\right)+\mathbf{X}_{n,y}\gamma+\alpha_{n}+\alpha_{y}+u_{n,y},$
+where $y_{n,y}$ is either an indicator for having a shareholder proposal at the AGM in year $y+1$ (panel (a)) or the year-on-year Fashion Transparency Index score improvement (panel (b)), $\mathbf{1}_{\{\text{Distance to AGM = }\tau\}}$ is an indicator for the report being published $\tau$ months away from firm $n$’s closest AGM, and $\text{Treated}_{n,y}$ is an indicator for firm $n$ being included in the index in year $y$.
+$\mathbf{X}_{n,y}$ includes the direct effect and controls for firm size.
+Standard errors are clustered at the firm level.
+Regression in panel (a) has 284 observations and regression in panel (b) has 162 observations.
+These results mirror the model’s main mechanism while accounting for target selection: AGM-timed actions mobilize shareholders, while earlier actions directly influence corporate behavior.
+9 Discussion
+Our findings show that when NGOs campaign can be as important as whether they campaign.
+Early in their life cycle, NGOs concentrate actions on AGM dates, when firms are most visible to investors and the media, echoing work on private politics and focal corporate events (baron2001private; egorov2017private; fioretti2024shared).
+These AGM‑timed campaigns raise media coverage, Google searches, and donations, and increase the likelihood that related shareholder proposals are filed at the next AGM, but they leave current votes and firms’ environmental and social (E&S) performance largely unchanged, aligning with a visibility‑oriented view of NGO activity (yaziji2009ngos).
+As NGOs gain reputation, their timing shifts.
+Campaigns in the months before the AGM raise the success and vote share of related shareholder proposals and are followed by improvements in firms’ E&S scores, consistent with influence operating through informed shareholder coordination and managerial concessions rather than immediate consumer reactions.
+The transition from AGM‑day exposure to pre‑AGM engagement therefore traces a lifecycle in which NGOs first invest in awareness and then leverage credibility to change corporate behavior (e.g., cabral2010dynamics; baur2011moral).
+Conceptually, the paper opens the black box of NGO activism by exploiting one of the few actions that must be public to be effective: campaigns.
+NGOs are otherwise highly opaque organizations, with limited standardized reporting and scarce data on internal decisions (maxwell2012economic).
+We show that the timing of these campaigns relative to the target’s corporate calendar reveals both (i) the NGO’s motive – whether it is primarily seeking visibility or trying to exert influence – and (ii) the main audience of the campaign – donors and the broader public versus shareholders and managers.
+Different audiences imply different mechanisms and map into the heterogeneous firm‑ and shareholder‑level responses we document across sections.
+To interpret these patterns, our model compares two mechanisms: an awareness stock $A_{t}$, built by highly visible actions such as AGM‑day campaigns, and a credibility state $\pi_{t}$, shaped by earlier, costlier engagement that moves beliefs and proposal success.
+Together, these channels rationalize NGOs’ transition from exposure‑oriented tactics to influence‑oriented activism and connect our evidence to work on shareholder voice, exit, and ESG engagement (e.g., cunat2016say; broccardo2020exit; hart2022private).
+Beyond these implications, our results also speak to future research on firm strategy and stakeholder salience.
+We still know relatively little about how firms decide which stakeholders matter in different contexts, despite a large theoretical literature on stakeholder prioritization and corporate objectives that is often qualitative or case-study based (e.g., freeman1984strategic; mitchell1997toward).
+By providing a novel, quasi-exogenous shift in NGO campaign timing – driven by NGOs’ own incentives to campaign early or late – we offer a quantitative framework to study how firms form and update their objective functions and whose welfare is put on the agenda, complementing recent empirical work on firms’ objectives and stakeholder weights (e.g., scott2002love; romer2006firms; duarte2025competitive).
+A further implication of our findings is that firms’ objectives – here, the objectives of NGOs as strategic actors – are not monolithic.
+Classic approaches often tie organizational objectives to ownership structure (e.g., grossman1977value; grossman1979theory), implicitly assuming that preferences flow from residual claimants downward.
+Our results instead show that objectives evolve with the views and expectations of other stakeholders.
+In the case of NGOs, it is not only donors and shareholders who influence their actions, but also the credibility concerns that arise from how these stakeholders interpret the timing of campaigns.
+NGOs shift away from AGM-day campaigning not simply to influence firms more effectively today, but because doing so is instrumental to building credibility, which expands their influence tomorrow.
+In this sense, organizational objectives are shaped endogenously by the incentives created by stakeholder attention and beliefs, rather than solely by ownership or mission.
+This perspective aligns with recent work emphasizing that firms’ (and nonprofits’) objectives emerge from their strategic interaction with multiple stakeholders rather than from a single principal (e.g., fioretti2022caring).
+These results have policy implications for boards, investors, and regulators in the corporate finance tradition.
+For boards, ignoring early, credibility‑backed campaigns is costly: these are precisely the actions that raise proposal success and are associated with subsequent E&S improvements, suggesting that constructive engagement with reputable NGOs can substitute for more confrontational AGM events (dimson2015active; kolbel2017media).
+For investors and stewardship codes, credible NGOs supply information that helps coordinate dispersed shareholders on E&S issues, complementing traditional voice and exit mechanisms.
+For regulators, rules on AGM timing, proposal access, and disclosure shape the returns to NGO campaigns and thereby how far private activism can help firms internalize social and environmental externalities alongside, or in place of, formal regulation.
+A natural question left open by our theory and evidence is how NGOs compete with one another.
+Prior work treats competition largely as a battle for funding—typically among large, global actors—and measures success by donations (e.g., aldashev2009ngos).
+Our framework suggests a broader view: NGOs also compete on visibility and credibility, and the form of competition depends on where organizations sit in that visibility–credibility space.
+When visibility is scarce, NGOs crowd focal moments (AGMs) to attract attention; when reputation matters, competition shifts toward earlier, costlier engagement that builds influence.
+This raises testable questions about the interaction between attention markets and reputation dynamics: How does competing for visibility affect reputation formation?
+Do multiple NGOs targeting the same firm amplify or dilute stakeholder coordination?
+Evidence from related contexts (e.g., microfinance NGOs) indicates competition can change outcomes (ly2012competition); studying its interactions with awareness and credibility is a promising avenue for future work.
+10 Conclusion
+This paper opens the black box of NGO behavior by showing how campaign timing reflects NGOs’ underlying objectives.
+Empirically, we use data on roughly 2,500 campaigns to show that AGM-day actions generate large visibility shocks, raising media coverage, search activity, donations, and the likelihood of related proposals at the next AGM, while having little effect on contemporaneous votes or firms’ E&S performance.
+By contrast, campaigns launched in the months before the AGM significantly increase proposal success, raise vote shares, and are followed by improvements in E&S outcomes.
+A simple model in which NGOs trade off two forces explains these patterns.
+The model features an awareness channel, whereby AGM visibility builds issue salience, and a credibility channel, whereby earlier, costlier actions update beliefs and directly influence shareholder behavior.
+Together, the model and evidence reveal NGOs as strategic actors whose objectives evolve with external incentives, vary across contexts, and reflect the views of multiple stakeholders who shape NGOs’ reputational payoffs.
+These insights matter for policy, showing that NGO campaigns mobilize different classes of investors with distinct implications for firms’ outcomes, whether by exerting direct influence through support for shareholder votes or indirectly, by prompting new proposals at future AGMs.
+Overall, the results demonstrate how civil-society pressure can help firms internalize social and environmental externalities.
+Appendix A Proofs
+A.1 Outcome Likelihood Ratio with Bernoulli Success
+Let $\theta\in\{C,W\}$ denote the NGO’s type. Conditional on state $(A_{t},\pi_{t})$ and timing $d_{t}$, the proposal outcome $y_{t}\in\{0,1\}$ is Bernoulli with type–specific success probabilities
+$S^{\theta}(A_{t},\pi_{t},d_{t})\;\equiv\;\Pr(y_{t}=1\mid A_{t},\pi_{t},d_{t},\theta).$
+Hence the conditional likelihood under type $\theta$ is
+$p^{\theta}(y_{t}\mid d_{t},A_{t},\pi_{t})=\big[S^{\theta}(A_{t},\pi_{t},d_{t})\big]^{y_{t}}\,\big[1-S^{\theta}(A_{t},\pi_{t},d_{t})\big]^{1-y_{t}}.$
+Lemma A1 (Bernoulli outcome log–likelihood ratio).
+For the two types $C$ and $W$, the outcome component of the log–likelihood ratio satisfies
+$\ln\frac{p^{C}(y_{t}\mid d_{t},A_{t},\pi_{t})}{p^{W}(y_{t}\mid d_{t},A_{t},\pi_{t})}=y_{t}\,\ln\frac{S^{C}(A_{t},\pi_{t},d_{t})}{S^{W}(A_{t},\pi_{t},d_{t})}\;+\;(1-y_{t})\,\ln\frac{1-S^{C}(A_{t},\pi_{t},d_{t})}{1-S^{W}(A_{t},\pi_{t},d_{t})}.$
+Proof.
+By the Bernoulli likelihoods above,
+$\frac{p^{C}(y_{t}\mid d_{t},A_{t},\pi_{t})}{p^{W}(y_{t}\mid d_{t},A_{t},\pi_{t})}=\frac{\big[S^{C}\big]^{y_{t}}\big[1-S^{C}\big]^{1-y_{t}}}{\big[S^{W}\big]^{y_{t}}\big[1-S^{W}\big]^{1-y_{t}}},$
+where for brevity we write $S^{\theta}=S^{\theta}(A_{t},\pi_{t},d_{t})$. Taking logs,
+$\ln\frac{p^{C}}{p^{W}}=y_{t}\big(\ln S^{C}-\ln S^{W}\big)+(1-y_{t})\big(\ln(1-S^{C})-\ln(1-S^{W})\big),$
+which is the stated expression. ∎
+Remarks.
+(i) If $S^{C}(A_{t},\pi_{t},d_{t})=S^{W}(A_{t},\pi_{t},d_{t})$ (e.g., at $d_{t}=\mathrm{AGM}$ when $_{\mathrm{AGM}}\equiv 0$), the outcome term is 0 and outcomes do not update beliefs.
+(ii) When $S^{C}>S^{W}$ (e.g., at $d_{t}=\mathrm{early}$ under (4) and (3)), a success $y_{t}=1$ raises the posterior odds of type $C$, while a failure $y_{t}=0$ lowers them.
+(iii) The expected value of the log–likelihood ratio under type $C$ equals the Kullback–Leibler divergence $\mathrm{KL}(\mathrm{Bern}(S^{C})\,\|\,\mathrm{Bern}(S^{W}))\geq 0$, with equality iff $S^{C}=S^{W}$, formalizing informativeness.
+A.2 Belief Consistency
+Iterating (7) gives
+$\operatorname{logit}(\pi_{t})=\operatorname{logit}(\pi_{0})+\sumop\displaylimits_{\tau=0}^{t-1}\Bigg[\ln\frac{p^{C}(d_{\tau}\mid\pi_{\tau},A_{\tau})}{p^{W}(d_{\tau}\mid\pi_{\tau},A_{\tau})}+y_{\tau}\ln\frac{S^{C}}{S^{W}}+(1-y_{\tau})\ln\frac{1-S^{C}}{1-S^{W}}\Bigg].$ (A1)
+Assume (i) the prior is non-degenerate, $\pi_{0}\in(0,1)$; (ii) signals are observed each period; (iii) early campaigns – where $S^{C}\neq S^{W}$ – occur infinitely often with positive probability; and (iv) the expected log-likelihood increments are bounded.
+Then, applying a law of large numbers to the subsequence of informative (early) periods, the cumulative log-likelihood ratio diverges almost surely with sign depending on the true type:
+[…]
+Hence posterior beliefs are consistent: $\pi_{t}\to 1$ almost surely under the credible type and $\pi_{t}\to 0$ under the weak type.
+Even though the timing rule is deterministic, the stochastic nature of proposal outcomes $y_{t}$ guarantees that belief trajectories and the transition from AGM to early campaigning remain random and path-dependent.
+A.3 Existence and Uniqueness of the Cutoff Policy and Equilibrium
+[…]
+Step 1: Existence and uniqueness of the value function
+[…]
+$(\mathcal{T}V)(\pi,A)=\max_{d\in\{\text{early},\text{AGM}\}}\Big\{u(d;\pi,A)+\beta\,\mathbb{E}[V(\pi^{\prime},A^{\prime})\mid\pi,A,d]\Big\}.$ (A2)
+Because $u$ is bounded and continuous, $0<\beta<1$, and the transition kernel $(\pi^{\prime},A^{\prime})$ is well defined and independent of $V$, $\mathcal{T}$ is a contraction on $\mathcal{B}$ with modulus $\beta$.
+Because $u(d;\pi,A)$ is increasing in $(\pi,A)$ by (3), and the transition kernel is monotone in $(\pi,A)$ (awareness increases in $A$, and expected posteriors are nondecreasing in $(\pi,A)$; cf.
+Step 2), $\mathcal{T}$ is monotone, so the fixed point $V^{\star}$ is increasing in each argument.
+This establishes existence and uniqueness of the value function.
+Step 2: Single-crossing in $\pi$ and monotonicity in AA
+Lemma A2 (Single-Crossing and Monotonicity).
+[…]
+Proof.
+Preliminaries.
+By standard contraction arguments, the value function $V(\pi,A)$ solving the Bellman equation is bounded, continuous, and increasing in each argument.
+(The one-period utility is increasing in $(\pi,A)$ and $0<\beta<1$.)
+(i) Current wedge. The current payoff wedge equals
+[…]
+By (3), early is strictly increasing in $\pi$ and weakly increasing in $A$. Hence the current wedge is strictly increasing in $\pi$ and weakly increasing in $A$.
+(ii) Continuation wedge. Write
+$^{\mathrm{cont}}(\pi,A)\;\equiv\;\mathbb{E}\big[V(\pi_{t+1},A_{t+1})\mid\pi,A,\mathrm{early}\big]-\mathbb{E}\big[V(\pi_{t+1},A_{t+1})\mid\pi,A,\mathrm{AGM}\big],$
+with $(\pi_{t+1},A_{t+1})$ evolving by (7) and (1).
+Awareness transition.
+Equation (1) implies that $A_{t+1}^{\mathrm{early}}-A_{t+1}^{\mathrm{AGM}}=(1-\rho)\,\big[v(\mathrm{early})-v(\mathrm{AGM})\big]<0$, a constant gap that is independent of both $\pi$ and $A$.
+As $A$ increases, both $A_{t+1}^{\mathrm{early}}$ and $A_{t+1}^{\mathrm{AGM}}$ rise with the same slope $\rho$, so the awareness gap between timings is unchanged by $A$ or $\pi$.
+Therefore, any dependence of cont on $(\pi,A)$ must come from the belief channel described next.
+Belief transition. In log-odds form,
+[…]
+[…]
+[…]
+$\quad+\;\underbrace{\mathbb{E}[\text{outcome LLR}\mid\mathrm{early}]}_{\geq 0},$
+[…]
+Using the add–subtract device and using superscripts $E$ and $G$ for variables computed for early and AGM timings respectively,
+[…]
+[…]
+[…]
+[…]
+Step 3: Existence and shape of the cutoff curve
+At this point we are ready to state the following proposition.
+Proposition A1 (Existence and Shape of the Cutoff Curve).
+For each level of awareness $A\in[0,1]$, there exists a unique continuous cutoff $\bar{\pi}(A)\in(0,1)$ such that the NGO campaigns early if and only if $\pi>\bar{\pi}(A)$.
+The cutoff is weakly decreasing in $A$.
+[…]
+The cutoff $\bar{\pi}(A)$ therefore separates timing choices: for $\pi<\bar{\pi}(A)$, both types optimally choose AGM campaigns, while for $\pi>\bar{\pi}(A)$ only the credible type finds early campaigning profitable.
+∎
+Step 4: Uniqueness of the Markov-Perfect equilibrium
+Because the Bellman operator (A2) is a contraction, the value function $V^{\star}$ is unique.
+The cutoff $\bar{\pi}(A)$ is also unique for each $A$, and the corresponding optimal policy $d^{\star}(\pi,A)$ is unique except on a measure-zero indifference set $\{\pi=\bar{\pi}(A)\}$.
+Hence the pair $(V^{\star},d^{\star})$ constitutes a unique pure-strategy MPE.
+Proposition A2 (Existence and Uniqueness of the Markov-Perfect Equilibrium).
+Under the maintained assumptions, the model admits a unique bounded, continuous value function $V^{\star}(\pi,A)$ solving the Bellman equation (A2), and a unique cutoff policy $\bar{\pi}(A)$ characterized by Proposition A1.
+The resulting pure-strategy Markov-perfect equilibrium is unique up to a measure-zero indifference set.
+Conclusion of the proof.
+Steps 2–4 jointly establish Proposition 1 in the main text.
+Lemma A2 delivers single crossing of the action–value difference, ensuring that timing incentives are monotone in credibility and awareness.
+Proposition A1 then yields a unique cutoff policy $\bar{\pi}(A)$, under which both types pool on AGM campaigns at low beliefs, while only the credible NGO switches to early campaigning once beliefs exceed the cutoff.
+Finally, Proposition A2 guarantees that this cutoff policy is part of a unique pure-strategy Markov-perfect equilibrium.
+Together, these results formalize the pooling–separating structure described in Proposition 1.
 Appendix B Omitted Figures
-Figure B1: Persistence of the AGM month over time
-Note: This figure plots the percentage of AGMs happening in each month of the year (y-axis), given the month of the previous year’s AGM (x-axis).
-A larger and darker circle indicates a greater proportion of meetings.
-Plot based on 38,471 meetings of 5,745 US companies, from 2010 to 2020.
-Figure B2: Firms’ negative media coverage increases if NGOs campaign during AGMs
-Note: The figure plots the coefficients $\beta_{\tau}$ from the following event study regression where we re-calendarize time around an AGM date so that the AGM month is month 0 in each year for each firm:
-$News_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\!\left(\beta_{\tau}\cdot 1\!\left\{M_{n,w,y}=\tau\!\right\}\cdot NGO\,Campaign_{n,w,y}\!\right)$
-$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
-where $News_{n,w,y}$ is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event.
-The set of dummies $1\{M_{n,w,y}=\tau\}$ indicate the distance in months from the closest AGM.
-The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise.
-We interact these event-study indicators, with the variable $NGO\,Campaign_{n,w,y}$, an indicator variable equal to 1 if firm n experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise.
-The vector $X_{n,w,y}$ includes the direct effect of these variables.
-The reference month is $\tau=-4$ following SEC rules on shareholder proposals.
-Note that, because of the stability of AGM dates across years within firms (Appendix Figure B1), the distribution of months fixed effects is unaffected by the re-calendarization, so that months are not double counted across years.
-Standard errors are clustered by firm.
-Figure B3: NGO campaigns and shareholder proposals
-Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead.
+Figure B1: How does target characteristics affect the decision to campaign on AGM dates?
+Note: The figure reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects.
+Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
+Figure B2: Increased NGO visibility around an AGM campaign – All sample
+(a) Campaign timing and campaigning NGO’s Google Trends scores
+(b) Campaign timing and number of target firms’ negative news (standardized)
+Note: Each panel plots the estimated $\beta^{AGM}$ coefficients from (10).
+In Panel (a), the dependent variable is the monthly Google Trends score for the NGO.
+In Panel (b), the dependent variable is the (normalized) number of monthly negative media coverage received by a target firm.
+Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
+Figure B3: NGO campaigns and shareholder proposals – All Sample
+(a) All NGOs
+Notes: Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead.
 The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects.
-The regression is based on 1,227 NGO campaigns.
-Standard errors are clustered at the firm level.
-Figure B4: NGO campaign timing and consumer backlash
-(a) NGO Campaigns on AGM dates
-(b) NGO Campaigns on Other Days
+Error bars shows 95% confidence intervals using standard errors that are clustered by firm.
+Figure B4: NGO campaign timing and donations – All sample
+(a) Donations
+(b) Operating Margin
 Notes: The figure plots the estimated ${\beta}_{\tau}$ from (12).
-Panel (a) focuses on NGO campaigns on AGM dates, while Panel (b) focuses on campaigns on dates other than AGM dates.
-The reference period is -1 quarter, consistent with the other analyses.
-The number of observations is 239,818 in Panel (a) and 239,925 in Panel (b).
-Standard errors are clustered at the firm level.
-Figure B5: Campaign Timing and E&S Score Improvement
-Note: This figure plots the estimated coefficients of the following regression:
-$_{5}\text{E\&S Score}_{n,q}=\sumop\displaylimits_{l=-6}^{5}\text{Campaign}_{n,q,l}+X_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
-where $_{5}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter q1q-1 and quarter $q+4$, and $\sumop\displaylimits_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM.
-The regression also controls ($X_{n,q}$) for firm size (logged total assets), lagged average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$).
-ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
-The regression has 47,082 observations.
-Standard errors are clustered at the firm level.
-Unlike an event study, equation (B1) is a distributed leads-and-lags regression.
-We omit a reference period, so the coefficients are interpreted relative to firms that did not experience a campaign.
-This choice allows us to directly compare the effects of early campaigns with those launched on the AGM date relative to firms without campaigns.
-Results remain similar if we instead set month –4 as the reference period, as in the other analyses.
+Panel (a) focuses on donations over (lagged) assets, while Panel (b) focuses on donations minus expenditures over (lagged) assets.
+Error bars shows 95% confidence intervals using standard errors that are clustered by NGO.
+Figure B5: Campaign Timing and Changes in E&S Scores – All Sample
+Note: The figure plots the estimated coefficients $\beta_{l}$ from equation (15).
+The dependent variable is the yearly change in a firm’s average E&S scores.
+Shaded areas show 95% confidence intervals based on firm-level clustered standard errors.
+Data combine MSCI ESG scores, NGO campaigns from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS.
 Figure B6: NGO past campaigns and the timing of current NGO campaigns – Alternative set of Fixed Effects
@@
 Standard errors are clustered at the firm level.
-The number of observations is 1,338.
 Figure B9: Past campaigns and target firm’s E&S characteristics
-(a) Target firm’s E score for Environmental campaigns
-(b) Target firm’s S score for Social campaigns
+(a) Target firm’s E score
+(b) Target firm’s S score
 Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating either the number of environmental (top panel) or social (bottom panel) campaigns the focal NGO had launched in previous years as in (16).
 Standard errors are clustered at the firm level.
-The number of observations is 636 and 698, respectively.
-Figure B10: Media coverage of NGO-target firms around their AGM
-Note: To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
-$\text{News}_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\left(\beta_{\tau}\cdot\mathbf{1}_{\{M_{n,w,y}=\tau\}}\cdot\text{NGO Campaign}_{n,w,y}\right)$
-$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
-[…]
 Appendix C Robustness of Google Trends Results
 This section discusses the robustness of the results presented in Figure 3 to alternative specifications.
-In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (Sun and Abraham, 2021), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
+In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (sun2021estimating), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
 C.1 Staggered Difference-in-Differences Specification
-We first replicate Figure 2, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (Sun and Abraham, 2021).
+We first replicate Figure 3, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (sun2021estimating).
 Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1.
-The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%.
+The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for young NGOs (first four campaigns), raising Google search intensity by about 58%.
 By contrast, NGOs see a more modest increase (33%) in search intensity for later campaigns, and only in the month of the campaign.
 Figure C1: Staggered difference-in-differences specification
-Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (1) using a staggered difference-in-differences estimator (Sun and Abraham, 2021).
+Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (10) using a staggered difference-in-differences estimator (sun2021estimating).
 The dependent variable is the monthly Google Trends score for the NGO.